### v1

# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

Abstract.
We explicitly compute the Rankin-Selberg type integral introduced by Piatetski-Shapiro over adeles for vector-valued Siegel cusp forms of square-free levels $\Gamma_{0}(N)$. On the way, for particular test functions in the Bessel models of irreducible admissible representations, exact evaluations of the local zeta-integrals are given.
2020 Mathematics Subject Classification:
primary 11F46, secondary 11F67
1. Introduction
Lately, the Bessel periods on the symplectic group $\mathsf{G}:={\bf GSp}_{2}$ of rank $2$ progressively gain its importance in the arithmetic of Siegel modular forms. The notion of Bessel period and the associated local model of admissible representations of $\mathsf{G}$ were originally introduced by Novodvorski and Piatetski-Shaprio in their work ([26], [27], [30]), where a modern treatment of Andrianov’s integral representation ([1], [2]) of the spinor $L$-function for Siegel modular forms was developed. Related to this, we should also mention independent work by Sugano ([43]), which handles vector-valued Hilbert-Siegel cusp forms on arithmetic groups defined by maximal lattices of indefinite division algebras. The Bessel period also plays an essential role in a formulation of Böcherer’s conjecture posed by Liu [23] in a style of the refined Gan-Gross-Prasad conjecture. This version of Böcherer’s conjecture has been completely solved by Furusawa and Morimoto ([10], [11]).
In this paper, we compute the Rankin-Selberg type integral a la Piatetski-Shapiro for holomorphic vector-valued Siegel cusp forms of square-free levels explicitly, using results by Pitale-Schmidt [32] on local new vectors on $\mathsf{G}({\mathbb{Q}}_{p})\,(p<\infty)$; since we rely on the local-global method, our result is immediately applied to non-holomorphic Siegel cusp forms of discrete series type (cf. [20], [21]). We explain the moral of our investigation for scalar-valued Siegel modular forms of weight $l$. Our focus is an asymptotic formula as $N+l\rightarrow\infty$ of the average
(1.1) $\frac{1}{[{\bf Sp}_{2}({\mathbb{Z}}):\Gamma_{0}(N)]}\sum_{\varphi\in{\mathscr{B}}(l,N)}\langle E(s,\Lambda,\mu)\mid\varphi\rangle\,B^{\Lambda}(\varphi),$
where $\mathsf{G}^{\#}$ is a subgroup of $\mathsf{G}$ defined by a quadratic field $E$ of discriminant $D<0$ (see §2.1), $E(s,\Lambda,\mu)$ is the Eisenstein series on $\mathsf{G}^{\#}({\mathbb{A}})$ attached to an ideal class character $\Lambda$ of $E$, a character $\mu:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}{\mathbb{R}}^{\times}\rightarrow{\mathbb{C}}^{1}$ and $s\in{\mathbb{C}}$ (see §3.1), $B^{\Lambda}(\varphi)$ is the Bessel period of $\varphi$ in an orthonormal basis ${\mathscr{B}}(l,N)$ of Siegel cusp forms of weight $l$ and square free level $\Gamma_{0}(N)$, and $\langle E(s,\Lambda,\mu)\mid\varphi\rangle$ is the $L^{2}$-inner product of $E(s,\Lambda,\mu)$ and $\varphi|\mathsf{G}^{\#}({\mathbb{A}})$. The conductor of $\mu$ is assumed to be relatively prime to $2DN$. Actually, for application ([19]), we consider a more general average than (1.1) by requiring ${\mathscr{B}}(l,N)$ to consist of joint eigenforms by Hecke operators outside $NM$ and by putting the eigenvalue of such a Hecke operator at each $\varphi$ in the formation of the sum (1.1). With such a generality, we obtain a completely explicit expression of (1.1) in terms of
  • The spinor $L$-function $L(s,\pi,\mu)$ (see §4.6) attached to an irreducible cupsidal representation $\pi\cong\otimes_{p}\pi_{p}$ twisted by $\mu$.
  • An average of Fourier coefficient of a newform $\varphi_{\pi}^{0}$ in $\pi$, denoted by $R(\varphi_{\pi}^{0},E,\Lambda)$ (see (4.12)) introduced in [17]).
  • An average of local Bessel periods of $\pi_{p}$ (see §5).
For the precise statement including the vector-valued case, we refer to Proposition 5.1 and Theorem 5.2. Owing to the explicit nature of our result, we are able to show that, asymptotically as the prime level $N$ grows with the weight $l$ being fixed, the contribution to (1.1) of the old forms, the Yoshida lifts and the Saito-Kurokawa lifts are 00 (see Theorem 5.4). These results are necessary in our forthcoming paper [19], where we shall work out the geometric side of a relative trace formula whose spectral side is essentially (1.1). We should remark that an average similar to our computed average (5.9) without the factors ${\widehat{L}}(s+1/2,\pi,\mu)$ and $\widehat{f_{S}}(\pi)$ is studied in [7, §3]. When $E={\mathbb{Q}}(i)$, our average (5.9) is almost the same as the one in [5, §8] except the appearance of the local periods.
Let us explain the structure of this paper. In §2, we introduce basic objects related to algebraic subgroups of the symplectic group $\mathsf{G}$ and Haar measures on their $p$-adic, real and adelic points. In §3.1, the Rankin-Selberg type integral and the global Bessel function for Siegel cusp forms are recalled; after this preparation, the basic identity (Lemma 3.2) is stated, which computes the pairing $\langle E(s,\Lambda,\mu),\varphi\rangle$ in (1.1) in terms of the Mellin transform of the global Bessel function attached to $\varphi$. We review the proof ([30]) to determine a constant depending on Haar measures exactly. In §4, we recall the definition of the quantity $R(\varphi,E,\Lambda)$ and describe its behavior under the Galois conjugation $\Lambda\mapsto\Lambda^{\dagger}$ (Lemma 4.6). The local multiplicity one theorem for Bessel models of $\pi_{p}$’s allows us to split $B^{\Lambda}(\varphi)$ for $\varphi\in\pi$ corresponding to a pure tensor in $\otimes_{p}\pi_{p}$ to an Euler product of local Bessel functions for $\pi_{p}$ up to a constant. To define our global newform $\varphi_{\pi}^{0}$ and identify this constant with $R(\varphi_{\pi}^{0},E,\Lambda)$ (see Lemma 4.7), we fix pairs of a local Bessel functional $\ell_{p}$ on $\pi_{p}$ and a local new vector $\xi_{p}$ of $\pi_{p}$ normalized by $\ell_{p}(\xi_{p})=1$; in §4.4, we explain that the main result of [32] ensures the existence of such $\{(\ell_{p},\xi_{p})\}_{p}$ when $\pi$ corresponds to a new forms on $\Gamma_{0}(N)$. In §6, we compute the local zeta-integrals by using results of [32]. The unramified computations are known as recalled in §6.2. Other cases, as well as the explicit determination of the local periods of prime level (see §6.8), seem to be new, so that some details are given in §6.5, §6.6, and §6.4. In §4.6, the functional equation of the spinor $L$-function is deduced from the basic identity by using Lemma 4.6. Although a “nice” functional equation of the spinor $L$-function itself is known in a broader generality ([40, Lemma 1.2]) owing to Arthur’s classification of automorphic representations of $\mathsf{G}({\mathbb{A}})$, we include this section for the sake of completeness and to provide a proof free from Arthur’s result, confirming that the local $L$-factor and the local $\varepsilon$-factor defined by [30] (in terms of a particular Bessel model) coincides with the ones proposed in [37, Tables 2,3]. In section §5, we prove Proposition 5.1 and Theorem 5.2; by invoking a result by Furusawa-Morimoto ([10, Theorem 8.1]), we obtain an explicit expression of (1.1) solely in terms of $L$-functions (Theorem 5.3).
2. Preliminaries
Let ${\mathsf{V}}$ denotes the space of symmetric matrices of degree $2$; we consider ${\mathsf{V}}$ as a ${\mathbb{Z}}$-scheme by defining
${\mathsf{V}}(R)=\left\{\left[\begin{smallmatrix}a&b\\ b&c\end{smallmatrix}\right]\mid a,b,c\in R\right\}$
for any commutative ring $R$. The group ${\bf GL}_{2}(R)$ acts on this space from the right as
(2.1) ${\mathsf{V}}(R)\times{\bf GL}_{2}(R)\ni(T,h)\quad\longmapsto\quad T{\mathsf{s}}(h):=\tfrac{1}{\det h}{}^{t}hTh\in{\mathsf{V}}(R).$
Let $\mathsf{G}$ denote the symplectic similitude group
$\mathsf{G}:=\{g\in{\bf GL}_{4}\mid{}^{t}g\left[\begin{smallmatrix}O&1_{2}\\ -1_{2}&O\end{smallmatrix}\right]g=\nu(g)\,\left[\begin{smallmatrix}O&1_{2}\\ -1_{2}&O\end{smallmatrix}\right]\quad(\exists\nu(g)\in{\bf GL}_{1}\}$
(usually denoted by ${\bf GSp}_{2}$), where $\nu:\mathsf{G}\rightarrow{\bf GL}_{1}$ is the similitude character. The center of $\mathsf{G}$ is denoted by ${\mathsf{Z}}$, which coincides with the set of all the scalar matrices in $\mathsf{G}$. Let $\mathsf{P}=\mathsf{M}\mathsf{N}$ be the Siegel parabolic subgroup of $\mathsf{G}$, where
$\mathsf{M}$ :=$:=\{\mathsf{m}(A,c):=\left[\begin{smallmatrix}A&O\\ O&{}^{t}A^{-1}c\end{smallmatrix}\right]\mid A\in{\bf GL}_{2}\,,c\in{\bf GL}_{1}\},$
$\mathsf{N}$ :=$:=\{\mathsf{n}(X):=\left[\begin{smallmatrix}1_{2}&X\\ O&1_{2}\end{smallmatrix}\right]\mid X\in{\mathsf{V}}\}.$
Note that $\nu(\mathsf{m}(A,c))=c$ and $\nu(\mathsf{n}(X))=1$. Let
${\mathsf{T}}:=\left\{\mathsf{m}(\left[\begin{smallmatrix}a&0\\ 0&b\end{smallmatrix}\right],c)\mid a,b,c\in{\bf GL}_{1}\right\}.$
The Weyl group $W:=N_{\mathsf{G}}({\mathsf{T}})/{\mathsf{T}}$ of $\mathsf{G}$ is generated by the images of
$s_{1}:=\left[\begin{smallmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{smallmatrix}\right],\quad s_{2}:=\left[\begin{smallmatrix}0&0&1&0\\ 0&1&0&0\\ -1&0&0&0\\ 0&0&0&1\end{smallmatrix}\right].$
2.1. Embedding of groups
Let $E={\mathbb{Q}}(\sqrt{D})$ be an imaginary quadratic field of discriminant $D<0$; we realize $E$ as a subfield of ${\mathbb{C}}$ so that $\sqrt{D}=i|D|^{1/2}$ with $i$ being the imaginary unit. Let $\mathfrak{o}_{E}$ denote the ring of integers of $E$. Choose a ${\mathbb{Z}}$-basis $\{1,\theta\}$ of $\mathfrak{o}_{E}$ so that
(2.2) $\theta-\bar{\theta}=-\sqrt{D},$
where $\tau\mapsto\bar{\tau}$ denotes the non-trivial automorphism of $E/{\mathbb{Q}}$.
The different of $E/{\mathbb{Q}}$, defined as ${\mathfrak{d}}_{E/{\mathbb{Q}}}^{-1}:=\{\tau\in E\mid{\operatorname{tr}}_{E/{\mathbb{Q}}}(\tau\mathfrak{o}_{E})\subset{\mathbb{Z}}\}$, is $\frac{1}{\sqrt{D}}\mathfrak{o}_{E}$; thus a symplectic ${\mathbb{Z}}$-bi-linear form $\langle\cdot,\cdot\rangle:\mathfrak{o}_{E}^{2}\times\mathfrak{o}_{E}^{2}\longrightarrow{\mathbb{Z}}$ is defined by
$\langle x,y\rangle:={\operatorname{tr}}_{E/{\mathbb{Q}}}\left(\tfrac{-1}{\sqrt{D}}\det\left[\begin{smallmatrix}x_{1}&y_{1}\\ x_{2}&y_{2}\end{smallmatrix}\right]\right),\quad x=\left[\begin{smallmatrix}x_{1}\\ x_{2}\end{smallmatrix}\right],\,y=\left[\begin{smallmatrix}x_{1}\\ x_{2}\end{smallmatrix}\right]\in\mathfrak{o}_{E}^{2}.$
Note that the following vectors of $\mathfrak{o}_{E}^{2}$ form a symplectic ${\mathbb{Z}}$-basis, i.e., $\langle v_{i}^{+},v_{j}^{-}\rangle=\delta_{ij}$ and $\langle v_{i}^{+},v_{j}^{+}\rangle=\langle v_{i}^{-},v_{j}^{-}\rangle=0$ for all $i,j$:
$v_{1}^{+}:=\left[\begin{smallmatrix}1\\ 0\end{smallmatrix}\right],\quad v_{2}^{+}:=\left[\begin{smallmatrix}\theta\\ 0\end{smallmatrix}\right],\quad v_{1}^{-}:=\left[\begin{smallmatrix}0\\ -\bar{\theta}\end{smallmatrix}\right],\quad v_{2}^{-}:=\left[\begin{smallmatrix}0\\ 1\end{smallmatrix}\right].$
For any ring $R$, set
$\mathsf{G}^{\#}(R):=\{h\in{\bf GL}_{2}(\mathfrak{o}_{E}\otimes_{\mathbb{Z}}R)\mid\det(h)\in R^{\times}\}.$
For $h\in\mathsf{G}^{\#}(R)$, viewed as an $R$-linear automorphism of $(\mathfrak{o}_{E}\otimes_{\mathbb{Z}}R)^{2}$, let $\iota_{\theta}(h)$ denote the $4\times 4$-matrix representing $h$ by the $R$-basis $\{v_{1}^{+},v_{2}^{+},v_{1}^{-},v_{2}^{-}\}$ of $(\mathfrak{o}_{E}\otimes_{\mathbb{Z}}R)^{2}$, i.e.,
(2.3) $[h(v_{1}^{+}),h(v_{2}^{+}),h(v_{1}^{-}),h(v_{2}^{-})]=[v_{1}^{+},v_{2}^{+},v_{1}^{-},v_{2}^{-}]\,\iota_{\theta}(h).$
Then $\iota_{\theta}(h)\in{\bf GL}_{4}(R)$ satisfies $\langle\iota_{\theta}(h)(x),\iota_{\theta}(h)(y)\rangle=\det(h)\,\langle x,y\rangle$ for all $x,y\in\mathfrak{o}_{E}^{2}$, i.e., $\iota_{\theta}(h)$ belongs to $\mathsf{G}(R)$ and $\nu(\iota_{\theta}(h))=\det h$. Thus, we have an embedding
$\iota_{\theta}:\mathsf{G}^{\#}(R)\longrightarrow\mathsf{G}(R).$
Let ${\mathsf{B}}^{\#}$ be the Borel subgroup of $\mathsf{G}^{\#}$ consisting of all those upper-triangular matrices in $\mathsf{G}^{\#}$; then, for any ${\mathbb{Q}}$-algebra $R$, the set ${\mathsf{B}}^{\#}(R)$ coincides with the set of all points of the form
(2.4) $\left[\begin{smallmatrix}a\tau&\beta\\ 0&\tau^{-1}\end{smallmatrix}\right],\quad\tau\in(E\otimes_{\mathbb{Q}}R)^{\times},\,a\in R^{\times},\,\beta\in R.$
Let $\mathsf{N}^{\#}$ denote the unipotent radical of ${\mathsf{B}}^{\#}$, i.e., $\mathsf{N}^{\#}(R)$ is the set of points with $\tau=a=1$. Set ${\mathsf{Z}}^{\#}:=\{a1_{2}\mid a\in{\bf GL}_{1}\}$; then ${\mathsf{Z}}^{\#}$ is a subgroup of the center of $\mathsf{G}^{\#}$ of index $2$.
Let $I_{\theta}:E\longrightarrow{\bf M}_{2}({\mathbb{Q}})$ denote the matrix representation of the regular representation of the ${\mathbb{Q}}$-algebra $E$ with respect to the basis $\{1,\theta\}$, i.e.,
(2.5) $[\tau,\tau\theta]=[1,\theta]\,I_{\theta}(\tau),\quad\tau\in E,$
or explicitly,
(2.6) $I_{\theta}(a+b\theta)=\left[\begin{matrix}a&-b{\operatorname{N}}_{E/{\mathbb{Q}}}(\theta)\\ b&a+b{\operatorname{tr}}_{E/{\mathbb{Q}}}(\theta)\end{matrix}\right],\quad a,b\in{\mathbb{Q}}.$
For $\beta=b_{2}+b_{3}\theta\in E_{R}$ with $b_{2},b_{3}\in R$, define an element $X_{\beta}$ of ${\mathsf{V}}(R)$ as
(2.7) $X_{\beta}:=\left[\begin{smallmatrix}b_{1}&b_{2}\\ b_{2}&b_{3}\end{smallmatrix}\right]\quad\text{with $b_{1}:=-b_{2}{\operatorname{tr}}_{E/{\mathbb{Q}}}(\theta)-b_{3}{\operatorname{N}}_{E/{\mathbb{Q}}}(\theta)$}.$
We have
(2.8) $\iota_{\theta}\left(\left[\begin{smallmatrix}\tau&0\\ 0&a\tau^{-1}\end{smallmatrix}\right]\right)$ $=\mathsf{m}(I_{\theta}(\tau),a),\quad a\in R^{\times},\,\tau\in E_{R}^{\times},$
(2.9) $\iota_{\theta}\left(\left[\begin{smallmatrix}1&\beta\\ 0&1\end{smallmatrix}\right]\right)$ $=\mathsf{n}(X_{\beta}),\quad\beta\in E_{R}.$
By formulas (2.6) and (2.8), we have $\iota_{\theta}(a1_{2})=a1_{4}$ for $(a\in R^{\times})$, which implies
(2.10) $\iota_{\theta}({\mathsf{Z}}^{\#})={\mathsf{Z}}.$
We then have ${\operatorname{N}}_{E/{\mathbb{Q}}}(x+\theta y)=\left[x,y\right]T\left[\begin{smallmatrix}x\\ y\end{smallmatrix}\right]$ with
(2.11) $T_{\theta}$ $:=\left[\begin{matrix}1&2^{-1}{\operatorname{tr}}_{E/{\mathbb{Q}}}(\theta)\\ 2^{-1}{\operatorname{tr}}_{E/{\mathbb{Q}}}(\theta)&{\operatorname{N}}_{E/{\mathbb{Q}}}(\theta)\end{matrix}\right].$
Let ${\mathsf{V}}^{T_{\theta}}({\mathbb{Q}})$ be the orthogonal of $T_{\theta}$ in ${\mathsf{V}}({\mathbb{Q}})$ with represent to the non-degenerate quadratic form ${\operatorname{tr}}(XY)$ on ${\mathsf{V}}({\mathbb{Q}})$, i.e.,
${\mathsf{V}}^{T_{\theta}}({\mathbb{Q}})$ :=$:=\{X\in{\mathsf{V}}({\mathbb{Q}})\mid{\operatorname{tr}}(T_{\theta}X)=0\}.$
Note that ${\operatorname{tr}}(T_{\theta}^{2})=1+({\operatorname{tr}}_{E/{\mathbb{Q}}}(\theta)/2)^{2}+{\operatorname{N}}_{E/{\mathbb{Q}}}(\theta)^{2}>0$, $\det(T_{\theta})=-D/4>0$ by (2.2), and $T_{\theta}\in{\mathsf{V}}({\mathbb{Z}})$. We have ${\mathsf{V}}^{T_{\theta}}({\mathbb{Q}})=\{X_{\beta}\mid\beta\in E\}$, and
(2.12) ${\mathsf{V}}({\mathbb{Q}})={\mathbb{Q}}T_{\theta}\oplus{\mathsf{V}}^{T_{\theta}}({\mathbb{Q}}),\qquad\iota_{\theta}(\mathsf{N}^{\#})=\mathsf{n}({\mathsf{V}}^{T_{\theta}})$
by (2.9). The group $\mathsf{M}$ acts on the space of rational homomomorhisms ${\rm{Hom}}(\mathsf{N},{\mathbb{G}}_{a})$ by the rule ${\rm Ad}^{*}(m)\chi(n)=\chi(m^{-1}nm)$ for $m\in\mathsf{M}$, $n\in\mathsf{N}$ and $\chi\in{{\rm{Hom}}}(\mathsf{N},{\mathbb{G}}_{a})$. For $T\in{\mathsf{V}}$, let $\mathsf{M}_{T}$ denote the stabilizer of $\chi_{T}:\mathsf{n}(X)\mapsto{\operatorname{tr}}(TX)$, and $\mathsf{M}_{T}^{\circ}$ the identity component with respect to the Zariski topology. Then, for any ${\mathbb{Q}}$-algebra $R$,
$\mathsf{M}_{T}^{\circ}(R)=\{\mathsf{m}(I_{\theta}(\tau),{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau))\mid\tau\in E_{R}\}.$
2.2. Open compact subgroups at finite places
Let $p$ be a prime number. Set $E_{p}:=E\otimes_{\mathbb{Q}}{\mathbb{Q}}_{p}$ and $\mathfrak{o}_{E,p}:=\mathfrak{o}_{E}\otimes_{\mathbb{Z}}{\mathbb{Z}}_{p}$. Define
${\bf K}^{\#}_{p}:=\mathsf{G}^{\#}({\mathbb{Z}}_{p}),\quad{\bf K}_{p}:=\mathsf{G}({\mathbb{Z}}_{p}).$
Then, ${\bf K}^{\#}_{p}$, which coincides with the stabilizer in $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ of the $\mathfrak{o}_{E,p}$-lattice $\mathfrak{o}_{E,p}^{2}\subset E_{p}^{2}$, is a maximal compact subgroup of $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$, and ${\bf K}_{p}$ is the standard maximal compact subgroup of $\mathsf{G}({\mathbb{Q}}_{p})$. For a non-zero ideal ${\mathfrak{n}}\subset{\mathbb{Z}}_{p}$, set
${\bf K}_{0}({\mathfrak{n}}):=\{\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in\mathsf{G}({\mathbb{Z}}_{p})\mid C\in{\mathfrak{n}}\}.$
For a non-zero ideal ${\mathfrak{a}}\subset\mathfrak{o}_{E,p}$, set
${\bf K}^{\#}_{0}({\mathfrak{a}}):=\{\left[\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right]\in\mathsf{G}^{\#}({\mathbb{Z}}_{p})\mid c\in{\mathfrak{a}}\}.$
Thus, ${\bf K}_{0}({\mathbb{Z}}_{p})={\bf K}_{p}$ and ${\bf K}_{0}^{\#}(\mathfrak{o}_{E,p})={\bf K}_{p}^{\#}$.
Lemma 2.1.
For a non-zero ideal ${\mathfrak{n}}\subset{\mathbb{Z}}_{p}$, we have
${\bf K}_{0}^{\#}({\mathfrak{n}}\mathfrak{o}_{E,p})=\iota_{\theta}^{-1}({\bf K}_{0}({\mathfrak{n}})).$
Proof.
Indeed, both $\iota_{\theta}^{-1}({\bf K}_{p})$ and ${\bf K}^{\#}_{p}$ coincides with the stabilizer of $\mathfrak{o}_{E,p}^{2}=\langle v_{1}^{+},v_{2}^{+},v_{1}^{-},v_{2}^{-}\rangle_{{\mathbb{Z}}_{p}}$ in $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$. Hence $\iota_{\theta}^{-1}({\bf K}_{p})={\bf K}^{\#}_{p}$. In the remaining part of the proof, we suppose ${\mathfrak{n}}\subset p{\mathbb{Z}}_{p}$. Then for $g=\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in{\bf K}_{p}$, we have $g\in{\bf K}_{0}({\mathfrak{n}})$ if and only if $g(v_{1}^{+}),g(v_{2}^{+})\in\langle v_{1}^{+},v_{2}^{+}\rangle_{{\mathbb{Z}}_{p}}+{\mathfrak{n}}\langle v_{1}^{-},v_{2}^{-}\rangle_{{\mathbb{Z}}_{p}}$. For $g=\iota_{\theta}(h)$ with $h=\left[\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right]\in{\bf K}_{p}^{\#}$, this last condition becomes $\left[\begin{smallmatrix}a\\ c\end{smallmatrix}\right],\,\left[\begin{smallmatrix}a\theta\\ c\theta\end{smallmatrix}\right]\in\left[\begin{smallmatrix}\mathfrak{o}_{E,p}\\ {\mathfrak{n}}\mathfrak{o}_{E,p}\end{smallmatrix}\right],$ or equivalently, $c\in{\mathfrak{n}}\mathfrak{o}_{E,p}$. ∎
For any $N\in{\mathbb{Z}}_{>0}$, define open compact subgroups ${\bf K}_{0}(N)\subset\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}})$ and ${\bf K}^{\#}(N\mathfrak{o}_{E})\subset\mathsf{G}^{\#}({\mathbb{A}}_{\rm{\bf{f}}})$ as
${\bf K}_{0}(N):=\prod_{p<\infty}{\bf K}_{0}(N{\mathbb{Z}}_{p}),\quad{\bf K}^{\#}_{0}(N\mathfrak{o}_{E}):=\prod_{p<\infty}{\bf K}_{0}^{\#}(N\mathfrak{o}_{E,p}).$
2.3. Maximal compact subgroup at the archimedean place
The identity connected component of $\mathsf{G}({\mathbb{R}})$ is $\mathsf{G}({\mathbb{R}})^{0}=\{g\in\mathsf{G}({\mathbb{R}})\mid\nu(g)>0\}$. Set ${\bf K}_{\infty}:=\mathsf{G}({\mathbb{R}})^{0}\cap{\bf O}(4)$, which is a maximal compact subgroup of $\mathsf{G}({\mathbb{R}})^{0}$ given as
${\bf K}_{\infty}=\{\left[\begin{smallmatrix}A&B\\ -B&A\end{smallmatrix}\right]\mid A,B\in{\bf M}_{2}({\mathbb{R}}),\,A+iB\in{\bf U}(2)\}.$
Note that ${\bf K}_{\infty}\subset{\bf Sp}_{2}({\mathbb{R}})$. The action of $\mathsf{G}({\mathbb{R}})^{0}$ on the Siegel upper-half space
${\mathfrak{h}}_{2}:=\{Z\in{\bf M}_{2}({\mathbb{C}})\mid{}^{t}Z=Z,\,{\operatorname{Im}}(Z)\gg 0\},$
denoted by $(g,Z)\mapsto g\langle Z\rangle$ is defined by the usual formula, i.e.,
$g\langle Z\rangle:=(AZ+B)(CZ+D)^{-1},\quad g=\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in\mathsf{G}({\mathbb{R}})^{0},\quad Z\in{\mathfrak{h}}_{2}.$
The stabilizer of $i1_{2}\in{\mathfrak{h}}_{2}$ coincides with ${\bf K}_{\infty}$. Set $\gamma_{\infty}:=\mathsf{m}(-1_{2},-1)=\left[\begin{smallmatrix}-1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\in\mathsf{G}({\mathbb{R}})$; then ${\bf K}_{\infty}\cup{\bf K}_{\infty}\gamma_{\infty}$ is a maximal compact subgroup of $\mathsf{G}({\mathbb{R}})$. The four vectors
$u_{1}:=\left[\begin{smallmatrix}1\\ 0\end{smallmatrix}\right],\quad u_{2}:=\left[\begin{smallmatrix}i\\ 0\end{smallmatrix}\right],\quad u_{3}:=\left[\begin{smallmatrix}0\\ -i\end{smallmatrix}\right],\quad u_{4}:=\left[\begin{smallmatrix}0\\ -1\end{smallmatrix}\right].$
in $E_{\mathbb{R}}^{2}$ form an ${\mathbb{R}}$-basis such that
(2.13) $[v_{1}^{+},v_{2}^{+},v_{1}^{-},v_{2}^{-}]=[u_{1},u_{2},u_{3},u_{4}]\,(b_{\mathbb{R}}^{\theta})^{-1}$
with
(2.14) $b_{\mathbb{R}}^{\theta}:=\mathsf{m}(A_{\theta},2^{-1}\sqrt{|D|})^{-1},\quad A_{\theta}:=\left[\begin{matrix}1&2^{-1}{\operatorname{tr}}_{E/{\mathbb{Q}}}(\theta)\\ 0&-2^{-1}\sqrt{|D|}\end{matrix}\right]\,\in{\bf GL}_{2}({\mathbb{R}}).$
Note that $b_{{\mathbb{R}}}^{\theta}\in\mathsf{G}({\mathbb{R}})^{0}$ because $\nu(b_{{\mathbb{R}}}^{\theta})^{-1}=2^{-1}\sqrt{|D|}>0$. Set
${\bf K}^{\#}_{\infty}:={\bf U}(2)\cap\mathsf{G}^{\#}({\mathbb{R}})=\{k^{\#}\in{\bf U}(2)\mid\det(k^{\#})=\pm 1\},$
which is a maximal compact subgroup of $\mathsf{G}^{\#}({\mathbb{R}})$. The (topological identity connected component $({\bf K}^{\#}_{\infty})^{0}$ of ${\bf K}^{\#}_{\infty}$ is ${\bf SU}(2)$, and ${\bf K}^{\#}_{\infty}=({\bf K}^{\#}_{\infty})^{0}\cup({\bf K}_{\infty}^{\#})^{0}\delta_{\infty}$ with $\delta_{\infty}:=\left[\begin{smallmatrix}1&0\\ 0&-1\end{smallmatrix}\right]$. A general element of ${\bf SU}(2)$ is written in the form
(2.15) $h=\left[\begin{smallmatrix}a&-\bar{b}\\ b&\bar{a}\end{smallmatrix}\right]\quad\text{with $a=a^{\prime}+ia^{\prime\prime},b=b^{\prime}+ib^{\prime\prime}\in{\mathbb{C}}$, $|a|^{2}+|b|^{2}=1$}.$
For such an $h$, a computation reveals the relation
(2.16) $[hu_{1},hu_{2},hu_{3},hu_{4}]=[u_{1},u_{2},u_{3},u_{4}]\left[\begin{smallmatrix}A&B\\ -B&A\end{smallmatrix}\right]\quad\text{with $A:=\left[\begin{smallmatrix}a^{\prime}&-a^{\prime\prime}\\ a^{\prime\prime}&a^{\prime}\end{smallmatrix}\right],\,B:=\left[\begin{smallmatrix}b^{\prime\prime}&b^{\prime}\\ b^{\prime}&-b^{\prime\prime}\end{smallmatrix}\right]$}.$
Lemma 2.2.
We have $\iota_{\theta}^{-1}({\bf K}_{\infty})=({\bf K}_{\infty}^{\#})^{0}$. For $k_{\infty}^{\#}\in({\bf K}^{\#}_{\infty})^{0}$ as in (2.15), and define $A,B\in{\bf M}_{2}({\mathbb{R}})$ as in (2.16). Then,
(2.17) $\iota_{\theta}(k_{\infty}^{\#})=b_{{\mathbb{R}}}^{\theta}\left[\begin{smallmatrix}A&B\\ -B&A\end{smallmatrix}\right]\,(b_{{\mathbb{R}}}^{\theta})^{-1}.$
Proof.
The formula in (2.17) follows from (2.4), (2.16) and (2.13) immediately. From
$[hu_{1},hu_{2},hu_{3},hu_{4}]=[u_{1},u_{2},u_{3},u_{4}](-\gamma)$
and $-\gamma\not\in{\bf K}_{\infty}$, the assertion follows. ∎
2.4. Haar measures
For locally compact unimodular topological groups $H$ relevant to us, we fix Haar measures $\eta_{H}$ on $H$ in the following manner.
2.4.1.
On the additive group ${\mathbb{C}}$, define ${\rm{d}}\eta_{{\mathbb{C}}}(\tau):={\rm{d}}x{\rm{d}}y=2^{-1}|{\rm{d}}\tau\wedge{\rm{d}}\bar{\tau}|$ for $\tau=x+iy\,(x,y\in{\mathbb{R}})$ with ${\rm{d}}x\,{\rm{d}}y$ being the Lebesgue measure on ${\mathbb{R}}^{2}$.
The Haar measure $\eta_{{\mathbb{C}}^{\times}}$ on ${\mathbb{C}}^{\times}$ (resp. ${\mathbb{R}}^{\times})$ is defined by ${\rm{d}}\eta_{{\mathbb{C}}^{\times}}(\tau)=|\tau|_{{\mathbb{C}}}^{-1}{\rm{d}}\tau$ for $\tau=x+iy\in{\mathbb{C}}$ (resp .$|x|_{\mathbb{R}}^{-1}{\rm{d}}x$). For each $p<\infty$, $E_{p}^{\times}$ (resp. ${\mathbb{Q}}_{p}^{\times}$) is endowed with the Haar measure $\eta_{E_{p}^{\times}}$ such that $\eta_{E_{p}^{\times}}(\mathfrak{o}_{E,p}^{\times})=1$ (resp. $\eta_{{\mathbb{Q}}_{p}^{\times}}({\mathbb{Z}}_{p}^{\times})=1$). Then, viewing ${\mathbb{A}}_{E}^{\times}$ as the restricted product of $E_{p}^{\times}$, we define $\eta_{{\mathbb{A}}_{E}^{\times}}=\prod_{p\leqslant\infty}\eta_{E_{p}^{\times}}$. Similarly, we set $\eta_{{\mathbb{A}}^{\times}}=\prod_{p\leqslant\infty}\eta_{{\mathbb{Q}}_{p}^{\times}}$.
2.4.2.
[…]
$y=\tfrac{-D}{2}\,x,\qquad\alpha=\beta+2^{-1}{\operatorname{tr}}_{E/{\mathbb{Q}}}(\theta)(x+y)+({\operatorname{N}}_{E/{\mathbb{Q}}}(\theta)x-y)\,\theta$
and since $|\tfrac{-D}{2}|_{\mathbb{A}}=1$, we get ${\rm{d}}\eta_{{\mathsf{V}}({\mathbb{A}})}(X)={\rm{d}}\eta_{{\mathbb{A}}}(y)\otimes{\rm{d}}\eta_{{\mathbb{A}}_{E}}(\alpha)$. Through the identification ${\mathsf{V}}({\mathbb{A}})\ni X\mapsto\mathsf{n}(X)\in\mathsf{N}({\mathbb{A}})$ and $\beta\ni{\mathbb{A}}_{E}\mapsto\left[\begin{smallmatrix}1&\beta\\ 0&1\end{smallmatrix}\right]\in\mathsf{N}^{\#}({\mathbb{A}})$, the groups $\mathsf{N}({\mathbb{A}})$ and $\mathsf{N}^{\#}({\mathbb{A}})$ acquire Haar measures. Thus, we have the integration formula:
(2.18) $\int_{\mathsf{N}({\mathbb{A}})}f(n)\,{\rm{d}}n=\int_{{\mathbb{A}}}\int_{\mathsf{N}^{\#}({\mathbb{A}})}f(\mathsf{n}(xT_{\theta})\,\iota_{\theta}(n^{\#}))\,{\rm{d}}x\,{\rm{d}}n^{\#}$
for any $f\in L^{2}(\mathsf{N}({\mathbb{A}}))$. Note that ${\operatorname{vol}}(\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}}))={\operatorname{vol}}(\mathsf{N}^{\#}({\mathbb{Q}})\backslash\mathsf{N}^{\#}({\mathbb{A}}))=1$.
2.4.3.
On the Lie group $\mathsf{G}^{\#}({\mathbb{R}})=\{g^{\#}\in{\bf GL}_{2}({\mathbb{C}})\mid\det(g^{\#})\in{\mathbb{R}}^{\times}\}$, we fix a Haar measure as follows. By the Iwasawa decomposition, any element $g^{\#}\in\mathsf{G}^{\#}({\mathbb{R}})$ is uniquely written as
(2.19) $g^{\#}=\left[\begin{smallmatrix}1&\beta\\ 0&1\end{smallmatrix}\right]\left[\begin{smallmatrix}a\tau&0\\ 0&\tau^{-1}\end{smallmatrix}\right]\,k^{\#}$
with $a\in{\mathbb{R}}^{\times}$, $\tau\in{\mathbb{C}}^{\times}$, $\beta\in{\mathbb{C}}$ and $k^{\#}\in{\bf K}_{\infty}^{\#}$. Then, our Haar measure on $\mathsf{G}^{\#}({\mathbb{R}})$ is symbolically defined as
(2.20) ${\rm{d}}g^{\#}=|a|_{\mathbb{R}}^{-2}|\tau|_{\mathbb{C}}^{-2}\times{\rm{d}}\eta_{{\mathbb{C}}}(\beta)\,\tfrac{{\rm{d}}\eta_{{\mathbb{R}}}(a)}{|a|_{\mathbb{R}}}\,\tfrac{{\rm{d}}\eta_{{\mathbb{C}}}(\tau)}{|\tau|_{\mathbb{C}}}\,\,{\rm{d}}k^{\#},$
where ${\rm{d}}k^{\#}$ is the Haar measure on ${\bf K}_{\infty}^{\#}$ with volume $1$, and $|\tau|_{\mathbb{C}}:=\tau\bar{\tau}$. For $p<\infty$, the group $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ is endowed with the Haar measure ${\rm{d}}g_{p}^{\#}$ uniquely determined by the condition ${\operatorname{vol}}({\bf K}_{p}^{\#})=1$. As in the archimedean case, if we write a general element $g^{\#}$ of $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ as in (2.19) with $a\in{\mathbb{Q}}_{p}^{\times}$, $\tau\in E_{p}^{\times}$, $\beta\in E_{p}$ and $k^{\#}\in{\bf K}_{p}^{\#}$, then
(2.21) ${\rm{d}}g^{\#}_{p}=|a|_{p}^{-2}|{\operatorname{N}}_{E_{p}/{\mathbb{Q}}_{p}}\tau|_{p}^{-2}\times{\rm{d}}\eta_{E_{p}}(\beta)\,{{\rm{d}}\eta_{{\mathbb{Q}}_{p}^{\times}}(a)}\,{{\rm{d}}\eta_{E_{p}^{\times}}(\tau)}\,{\rm{d}}k^{\#},$
where ${\rm{d}}k^{\#}$ is the normalized Haar measure of ${\bf K}^{\#}_{p}$. On the adele group $\mathsf{G}^{\#}({\mathbb{A}})$, we use the product measure of Haar measures on $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ for $p\leqslant\infty$. Then, from (2.20) and (2.21), we get
(2.22) $\int_{\mathsf{G}^{\#}({\mathbb{A}})}f(g^{\#})\,{\rm{d}}g^{\#}=\tfrac{\sqrt{|D|}}{2}\int_{\mathsf{N}^{\#}({\mathbb{A}})}\int_{{\mathbb{A}}^{\times}}\int_{{\mathbb{A}}^{\times}_{E}}\int_{{\bf K}^{\#}}f$ $\left(n^{\#}\left[\begin{smallmatrix}a\tau&0\\ 0&\tau^{-1}\end{smallmatrix}\right]\,k^{\#}\right)|a|_{\mathbb{A}}^{-2}|{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau)|_{\mathbb{A}}^{-2}$
×$\qquad\times{\rm{d}}n^{\#}\,{\rm{d}}\eta_{{\mathbb{A}}^{\times}}(a)\,{\rm{d}}\eta_{{\mathbb{A}}_{E}^{\times}}(\tau)\,{\rm{d}}k^{\#},$
where ${\rm{d}}k^{\#}$ is the normalized Haar measure on ${\bf K}^{\#}=\prod_{p\leqslant\infty}{\bf K}_{p}^{\#}$. Note that the measure on $\mathsf{N}^{\#}({\mathbb{A}})\cong{\mathbb{A}}_{E}$ is $\eta_{{\mathbb{A}}_{E}}$ defined above, which is $\frac{\sqrt{|D|}}{2}\times\prod_{p}\eta_{E_{p}}\,(p\leqslant\infty)$ due to the formulas $(\prod_{p}\eta_{E_{p}})({\mathbb{A}}_{E}/E)=|D|^{1/2}/2$ ([49, Chap V §4 Proposition 7]) and $\eta_{{\mathbb{A}}_{E}}({\mathbb{A}}_{E}/E)=1$.
2.4.4.
We fix $\eta_{{\bf K}_{\infty}}$ so that the measure of ${\bf K}_{\infty}$ is $1$. Then we normalize $\eta_{{\bf Sp}_{2}({\mathbb{R}}))}$ in such a way that the quotient $\eta_{{\bf Sp}_{2}({\mathbb{R}})}/\eta_{{\bf K}_{\infty}}$ corresponds to the measure $(\det Z)^{-3}{\rm{d}}X{\rm{d}}Y$ on the Siegel upper-half space ${\mathfrak{h}}_{2}\cong{\bf Sp}_{2}({\mathbb{R}})/{\bf K}_{\infty}$. By $\mathsf{G}({\mathbb{R}})^{0}={\mathsf{Z}}({\mathbb{R}})^{0}\,{\bf Sp}_{2}({\mathbb{R}})\cong{\mathbb{R}}_{>0}\times{\bf Sp}_{2}({\mathbb{R}})$, $\eta_{\mathsf{G}({\mathbb{R}})}$ is defined by demanding that its restriction to $\mathsf{G}({\mathbb{R}})^{0}$ is $\eta_{{\mathbb{R}}^{\times}}\otimes\eta_{{\bf Sp}_{2}({\mathbb{R}})}$. For $p<\infty$, we fix $\eta_{\mathsf{G}({\mathbb{Q}}_{p})}$ by ${\operatorname{vol}}({\bf K}_{p})=1$. Then, $\eta_{\mathsf{G}({\mathbb{A}})}$ is defined to be the restricted product of $\eta_{\mathsf{G}({\mathbb{Q}}_{v})}\,(v\leqslant\infty)$.
2.5. Idele class characters
Let $\Lambda$ denote a character of the finite group
${\rm Cl}(E):={\mathbb{A}}_{E}^{\times}/E^{\times}E_{\mathbb{R}}^{\times}\widehat{\mathfrak{o}}_{E}^{\times}.$
As is well known, this group is isomorphic to the ideal class group of $E$. We regard $\Lambda$ as an idele class character of $E^{\times}$ of finite order. Since $^{\mathbb{A}}^{\times}={\mathbb{Q}}^{\times}({\mathbb{R}}_{>0})\widehat{\mathbb{Z}}^{\times}$ is contained in the subgroup $E^{\times}E_{\mathbb{R}}^{\times}\widehat{\mathfrak{o}}_{E}^{\times}$, we have
(2.23) $\Lambda|{\mathbb{A}}^{\times}={\bf 1}.$
Let $v$ be a place of $E$ and $\Lambda_{v}$ the $v$-component of $\Lambda$, i.e., $\Lambda=\otimes_{v}\Lambda_{v}$. We say that $v$ is inert in $E/{\mathbb{Q}}$, splits in $E/{\mathbb{Q}}$ or ramifies in $E/{\mathbb{Q}}$ if $E_{v}:=E\otimes_{\mathbb{Q}}{\mathbb{Q}}_{p}$ is an unramified field extension of ${\mathbb{Q}}_{p}$, is isomorphic to ${\mathbb{Q}}_{p}\oplus{\mathbb{Q}}_{p}$, or is a ramified field extension of ${\mathbb{Q}}_{p}$, respectively. Since $\Lambda$ is supposed to be trivial in $E_{\mathbb{R}}^{\times}$, we have $\Lambda_{\infty}={\bf 1}$; moreover,
(2.24) If v is inert in E/, then $\text{If $v$ is inert in $E/{\mathbb{Q}}$, then $\Lambda_{v}={\bf 1}$}.$
Indeed, since $\Lambda$ is supposed to be trivial on $\widehat{\mathfrak{o}}_{E}^{\times}$, the restriction $\Lambda_{v}|{\mathfrak{o}_{E,v}^{\times}}$ is trivial. If $v$ is inert and $p$ denote the residue characteristic of $E_{v}$, then $p\in{\mathbb{Q}}_{p}$ is a prime element of the local field $E_{v}$. We have $\Lambda_{v}(p)=1$ due to (2.23). Hence, $\Lambda_{v}$ is trivial on $p^{{\mathbb{Z}}}\mathfrak{o}_{E,v}^{\times}=E_{p}^{\times}$.
Define the Galois conjugate $\Lambda^{\dagger}$ of $\Lambda$ by setting
(2.25) $\Lambda^{\dagger}(\tau):=\Lambda(\bar{\tau}),\quad\tau\in{\mathbb{A}}_{E}^{\times}.$
We have $\Lambda(\tau)\Lambda^{\dagger}(\tau)=\Lambda(\tau\bar{\tau})=1$ for $\tau\in{\mathbb{A}}_{E}^{\times}$ by (2.23). Hence
(2.26) $\Lambda^{\dagger}=\Lambda^{-1}=\bar{\Lambda},$
where $\bar{\Lambda}$ is the complex conjugate of $\Lambda$.
For an idele class character $\xi$ of $E^{\times}$, let $\widehat{L}(s,\xi)$ be the completed Hecke $L$-function of $\xi$, and $L_{p}(s,\xi)$ its local $p$-factor for $p\leqslant\infty$. Then, $L_{p}(s,\xi^{\dagger})=L_{p}(s,\xi)$ for any $p<\infty$. Indeed, if $p$ is not ramified in $E/{\mathbb{Q}}$, the equality is trivial. Suppose that $E/{\mathbb{Q}}$ is ramified; if $\xi_{p}$ is a ramified character of $E_{p}^{\times}$, then both $L$-factors are $1$. If $\xi_{p}$ is unramified, then for any prime element $\varpi$ of $E_{p}$, we have $(\xi^{\dagger})_{p}(\varpi)=\xi_{p}(\bar{\varpi})=\xi_{p}(\varpi)$, which implies the equality between local $p$-factors.
For any character $\mu:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}{\mathbb{R}}^{\times}\rightarrow{\mathbb{C}}^{1}$, we define $\mu_{E}:=\mu\circ{\operatorname{N}}_{E/{\mathbb{Q}}}:{\mathbb{A}}^{\times}_{E}/E^{\times}\rightarrow{\mathbb{C}}^{1}$. Then, $\mu_{E}$ is Galois invariant, and $\mu_{E}|E_{\infty}^{\times}={\bf 1}$. Hence, $(\Lambda\mu_{E})^{\dagger}=\Lambda^{\dagger}\mu_{E}^{\dagger}=\Lambda^{-1}\mu_{E}$, and
(2.27) $\widehat{L}(s,\Lambda^{-1}\mu_{E})=\widehat{L}(s,\Lambda^{\dagger}\mu_{E})=\widehat{L}(s,\Lambda\mu_{E}),\quad({\operatorname{Re}}(s)>1).$
Note that $\widehat{L}(s,\Lambda\mu_{E})$ is holomorphic except for possible simple poles at $s=1,0$, which occurs if and only if both $\Lambda$ and $\mu$ are trivial.
3. Eisenstein series and Rankin-Selberg integral
3.1. Eisenstein series
For details, we refer to [15, §19]; the theory on ${\bf GL}_{2}({\mathbb{A}}_{E})$ developed there carries over into the group $\mathsf{G}^{\#}({\mathbb{A}})$ by a minor modification. For a finite-dimensional vector space $V$ over a local field, let $\mathcal{S}(V)$ be the space of all Schwartz-Bruhat functions on $V$. For $p{\leqslant}\infty$, $\phi\in\mathcal{S}(E_{p}^{2})$ and $s\in\mathbb{C}$ with ${\operatorname{Re}}(s)>0$, we define a function $f_{\phi}^{(s,\Lambda_{p},\mu_{p})}:\mathsf{G}^{\#}({\mathbb{Q}}_{p})\rightarrow{\mathbb{C}}$ by
[…]
When $p=\infty$, we assume that $\phi$ is ${\bf K}_{\infty}^{\#}$-finite. By the local Tate’s theory, the function $s\mapsto f_{\phi}^{(s,\Lambda_{p},\mu_{p})}(g^{\#})$ is continued meromorphically to ${\mathbb{C}}$ in such a way that
(3.2) $\mathsf{f}_{\phi}^{(s,\Lambda_{p},\mu_{p})}(g^{\#}):=L_{p}(s+1,\Lambda_{p}\mu_{E,p})^{-1}\times f_{\phi}^{(s,\Lambda_{p},\mu_{p})}(g^{\#})$
is holomorphic on ${\mathbb{C}}$. Then $\mathsf{f}_{\phi}^{(s,\Lambda_{p},\mu_{p})}$ belongs to the space $\mathscr{V}^{\#}(s,\Lambda_{p},\mu_{p})$ consisting of all smooth functions $f$ on $\mathsf{G}^{\#}({\mathbb{Q}}_{p})$ satisfying
$f([\begin{smallmatrix}\tau&\beta\\ 0&a\bar{\tau}\end{smallmatrix}]g^{\#})=\Lambda_{p}(\tau)^{-1}\mu_{p}(a)^{-1}|a|_{p}^{-(s+1)}f(g^{\#})$
for any $[\begin{smallmatrix}\tau&\beta\\ 0&a\bar{\tau}\end{smallmatrix}]\in{\mathsf{B}}^{\#}({\mathbb{Q}}_{p})$ and $g^{\#}\in\mathsf{G}^{\#}({\mathbb{Q}}_{p})$.
Let $\mathcal{S}({\mathbb{A}}_{E}^{2})$ be the space of all Schwartz-Buruhat functions on ${\mathbb{A}}_{E}^{2}$. The Fourier transform $\widehat{\phi}$ of $\phi\in\mathcal{S}({\mathbb{A}}_{E}^{2})$ is defined by
[…]
Set $w_{0}:=\left[\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right]\in\mathsf{G}^{\#}({\mathbb{Q}}_{p})$. The standard intertwining operator
$M(s):{\mathscr{V}}^{\#}(s,\Lambda_{p},\mu_{p})\longrightarrow{\mathscr{V}}^{\#}(-s,\Lambda_{p}^{-1},\mu_{p}^{-1})$
is defined on ${\operatorname{Re}}(s)>0$ as the absolutely convergent integral
$M(s)f(g^{\#})=\int_{E_{p}}f\left(w_{0}\left[\begin{smallmatrix}1&\beta\\ 0&1\end{smallmatrix}\right]g^{\#}\right)\,{\rm{d}}\eta_{E_{p}}(\beta),\quad g^{\#}\in\mathsf{G}^{\#}({\mathbb{Q}}_{p}).$
The effect of $M(s)$ on the section $f_{\phi}^{(s,\Lambda_{p},\mu_{p})}$ is described by the Fourier transform $\widehat{\phi}$ as
(3.3) $f_{\widehat{\phi}}^{(-s,\Lambda_{p}^{-1},\mu_{p}^{-1})}(g^{\#})=c_{p}\Lambda_{p}(\sqrt{D})^{-1}|D|_{p}^{-s+1/2}\,\varepsilon_{p}(s,\Lambda\mu_{E},\psi_{E_{p}})\frac{L_{p}(1-s,\Lambda^{-1}\mu_{E}^{-1})}{L_{p}(s,\Lambda\mu_{E})}\times M(s)f_{\phi}^{(s,\Lambda_{p},\mu_{p})}(g^{\#})$
with $c_{p}=1$ if $p<\infty$ and $c_{\infty}=\frac{\sqrt{|D|}}{2}$ if $p=\infty$, where $\varepsilon(s,\bullet,\psi_{E_{p}})$ denote Tate’s local epsilon factor defined by the character $\psi_{E_{p}}$ of $E_{p}$ and the associated self-dual measure on $E_{p}$ as usual. For a decomposable element $\phi=\otimes_{p{\leqslant}\infty}\phi_{p}\in\mathcal{S}({\mathbb{A}}_{E}^{2})$, we define
$\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(g^{\#})=\prod_{p{\leqslant}\infty}\mathsf{f}_{\phi_{p}}^{(s,\Lambda_{p},\mu_{p})}(g_{p}^{\#}),\quad g^{\#}=(g_{p}^{\#})_{p}\in\mathsf{G}^{\#}({\mathbb{A}}).$
Note that $\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}$ is left ${\mathsf{B}}^{\#}({\mathbb{Q}})$- invariant and right ${\bf K}^{\#}$-finite.
The Eisenstein series attached to $\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}$ is defined by the absolutely convergent series
(3.4) $E(\phi,s,\Lambda,\mu;g^{\#}):=\widehat{L}(s+1,\Lambda\mu_{E})\times\sum_{\delta\in{\mathsf{B}}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{Q}})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(\delta g^{\#}),\quad g^{\#}\in\mathsf{G}^{\#}({\mathbb{A}})$
for ${\operatorname{Re}}(s)>1$. The following property of the Eisenstein series is standard.
Proposition 3.1.
Let $\phi\in\mathcal{S}({\mathbb{A}}_{E}^{2})$, $s\in{\mathbb{C}}$, $\Lambda\in\widehat{{\rm Cl}({E})}$, $\mu\in\widehat{{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}{\mathbb{R}}^{\times}}$ and $g^{\#}\in\mathsf{G}^{\#}({\mathbb{A}})$.
  • (i)
    The map $s\mapsto E(\phi,s,\Lambda,\mu;g^{\#})$ $({\operatorname{Re}}(s)>1)$ has a meromorphic continuation to ${\mathbb{C}}$, holomorphic in $s$ unless $\Lambda\mu_{E}\not={\bf 1}$ in which case it has possible simple poles only at $s=1,-1$. The function $g^{\#}\mapsto E(\phi,s,\Lambda,\mu;g^{\#})$ is an automorphic form on $\mathsf{G}^{\#}({\mathbb{Q}}){\mathsf{Z}}^{\#}({\mathbb{A}})\backslash\mathsf{G}^{\#}({\mathbb{A}})$.
  • (ii)
    We have the functional equation $E(\widehat{\phi},-s,\Lambda^{-1},\mu^{-1};g^{\#})=E(\phi,s,\Lambda,\mu;g^{\#})$.
  • (iii)
    For a relatively compact subset ${\mathcal{N}}\subset{\mathbb{C}}$ on which $s\mapsto E(\phi,s,\Lambda,\mu;g^{\#})$ is regular and for a compact set ${\mathcal{U}}\subset\mathsf{G}^{\#}({\mathbb{A}})$, there exists constants $C>0$ and $N>0$ such that
    $|E(\phi,s,\Lambda,\mu;\left[\begin{smallmatrix}a&0\\ 0&1\end{smallmatrix}\right]\,g^{\#})|\leqslant C\,|a|^{N}\quad(a\in{\mathbb{R}}_{>1},\,g^{\#}\in{\mathcal{U}},\,s\in{\mathcal{N}}).$
3.2. Rankin-Selberg integral and the basic identity
For any cusp form $\varphi$ on $\mathsf{G}({\mathbb{A}})$, $\Lambda\in\widehat{{\rm Cl}(E)}$ and $s\in{\mathbb{C}}$, the Rankin-Selberg integral is defined by
(3.5) $\langle E(\phi,s,\Lambda,\mu),\varphi\rangle:=\int_{{\mathsf{Z}}^{\#}({\mathbb{A}})\mathsf{G}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{A}})}E(\phi,s,\Lambda,\mu;g^{\#})\,\varphi(\iota_{\theta}(g^{\#}))\,{\rm{d}}g^{\#}.$
By Lemma 3.1(iii) and by the Fourier expansion (4.5), it is straightforward to show the absolute convergence of the integral for $s\in{\mathbb{C}}$ where the Eisenstein series is regular. Let $\psi:{\mathbb{A}}/{\mathbb{Q}}\rightarrow{\mathbb{C}}^{1}$ be the basic character. For $T\in{\mathsf{V}}({\mathbb{Q}})$, define a character of $\mathsf{N}({\mathbb{A}})$ by
$\psi_{T}(\mathsf{n}(X))=\psi({\operatorname{tr}}(TX)),\quad X\in{\mathsf{V}}({\mathbb{A}}).$
The $(T_{\theta},\Lambda)$-Bessel period of a cusp form $\varphi$ on ${\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})$ is defined by the integral111 Note that $\mathsf{m}(I_{\theta}(a),{\operatorname{N}}_{E/{\mathbb{Q}}}(a))=a1_{4}$ for $a\in{\mathbb{A}}^{\times}$. Then, due to $\varphi$ being ${\mathsf{Z}}({\mathbb{A}})$-invariant, the integral $B^{T_{\theta},\Lambda}(\varphi;g)$ is 00 if $\Lambda|{\mathbb{A}}^{\times}\not=1$. This trivial vanishing is not the case by (2.23).
(3.6) $B^{T_{\theta},\Lambda}(\varphi;g):=\int_{{\mathbb{A}}_{E}^{\times}/E^{\times}{\mathbb{A}}^{\times}}\Lambda(\tau)^{-1}\,\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\psi_{T_{\theta}}(n)^{-1}\varphi(\mathsf{m}(I_{\theta}(\tau),{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau))\,n\,g)\,{\rm{d}}n\,{\rm{d}}^{\times}\tau$
for $g\in\mathsf{G}({\mathbb{A}})$, where ${\rm{d}}^{\times}\tau$ denote the quotient measure on ${\mathbb{A}}_{E}^{\times}/{\mathbb{A}}^{\times}$ (see §2.4.1). Here is the basic identity, which is originally due to Andrianov in the classical setting ([1], [2]) and is adelically reformulated by Piatetski-Shapiro ([30]). To determine constants exactly under our normalization of Haar measures, we reproduce the proof briefly.
Lemma 3.2.
Let $\varphi:{\mathsf{Z}}({\mathbb{A}})\backslash\mathsf{G}({\mathbb{Q}})\mathsf{G}({\mathbb{A}})\longrightarrow{\mathbb{C}}$ be a cusp form. For ${\operatorname{Re}}(s)>1$, we have
(3.7) $\langle E(\phi,s,\Lambda,\mu),\varphi\rangle$ $=\tfrac{\sqrt{|D|}}{2}L(s+1,\Lambda\mu_{E})$
×$\times\int_{{\mathbb{A}}^{\times}}\int_{{\bf K}_{\infty}^{\#}\,{\bf K}_{0}^{\#}(\mathfrak{o}_{E})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(k^{\#})\mu(a)|a|_{\mathbb{A}}^{s-1}\,B^{T_{\theta},\Lambda}(\varphi;\mathsf{m}(a1_{2},a)\iota_{\theta}(k^{\#})b_{\mathbb{R}}^{\theta})\,{\rm{d}}^{\times}a\,{\rm{d}}k^{\#}.$
Proof.
By substituting (3.4) to (3.5),
$L(s+1,\Lambda\mu_{E})^{-1}\times\langle E(\phi,s,\Lambda,\mu),R(b_{\mathbb{R}}^{\theta})\varphi\rangle$
$=\int_{{\mathsf{Z}}^{\#}({\mathbb{A}})\mathsf{G}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{A}})}\sum_{\gamma\in{\mathsf{B}}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{Q}})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(\gamma g^{\#})\,\varphi(\iota_{\theta}(g^{\#})b_{\mathbb{R}}^{\theta})\,{\rm{d}}g^{\#}$
$=\int_{{\mathsf{Z}}^{\#}({\mathbb{A}}){\mathsf{B}}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{A}})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(g^{\#})\,\varphi(\iota_{\theta}(g^{\#})b_{\mathbb{R}}^{\theta})\,{\rm{d}}g^{\#}$
$=\tfrac{\sqrt{|D|}}{2}\int_{{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}}\int_{{\mathbb{A}}_{E}^{\times}/{\mathbb{A}}^{\times}E^{\times}}\int_{\mathsf{N}^{\#}({\mathbb{Q}})\backslash\mathsf{N}^{\#}({\mathbb{A}})}\int_{{\bf K}_{0}^{\#}(\mathfrak{o}_{E}){\bf K}_{\infty}^{\#}}|a|_{\mathbb{A}}^{-(s+1)}\Lambda(\tau)^{-1}\mu(a)^{-1}$
×$\quad\times\varphi(\iota_{\theta}(n^{\#})\,\mathsf{m}(I_{\theta}(\tau),{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau))\,\left[\begin{smallmatrix}a^{-1}1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\iota_{\theta}(k_{\rm{\bf{f}}}^{\#}))\,|a|_{\mathbb{A}}^{2}\,{\rm{d}}^{\times}a\,{\rm{d}}^{\times}\tau\,{\rm{d}}n^{\#}\,{\rm{d}}k_{\rm{\bf{f}}}^{\#}\,{\rm{d}}k_{\infty}^{\#},$
where the last equality is proved by (2.22). By Lemma 3.3, the last expression becomes
$\tfrac{\sqrt{|D|}}{2}\int_{{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}}\int_{{\mathbb{A}}_{E}^{\times}/{\mathbb{A}}^{\times}E^{\times}}\int_{\mathsf{N}^{\#}({\mathbb{Q}})\backslash\mathsf{N}^{\#}({\mathbb{A}})}\int_{{\bf K}_{0}^{\#}(\mathfrak{o}_{E}){\bf K}_{\infty}^{\#}}|a|_{\mathbb{A}}^{-(s+1)}\Lambda(\tau)^{-1}\mu(a)^{-1}$
×$\quad\times\sum_{\alpha\in{\mathbb{Q}}^{\times}}\varphi(n\,\mathsf{m}(I_{\theta}(\tau),{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau))\,\left[\begin{smallmatrix}\alpha a^{-1}1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\iota_{\theta}(k_{\rm{\bf{f}}}^{\#}))\,\psi_{T_{\theta}}(n)^{-1}\,|a|_{\mathbb{A}}^{2}\,{\rm{d}}^{\times}a\,{\rm{d}}^{\times}\tau\,{\rm{d}}n\,{\rm{d}}k_{\rm{\bf{f}}}^{\#}.$
The $\alpha$-summation and the $a$-integral over ${\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}$ are combined to yields an integral over ${\mathbb{A}}^{\times}$. Then, by the change of variables $a\mapsto a^{-1}$, we get the desired formula by (3.6). ∎
Lemma 3.3.
$\int_{\mathsf{N}^{\#}({\mathbb{Q}})\backslash\mathsf{N}^{\#}({\mathbb{A}})}\varphi(\iota_{\theta}(n^{\#})g)\,{\rm{d}}n^{\#}=\sum_{\alpha\in{\mathbb{Q}}^{\times}}\int_{{\mathsf{V}}({\mathbb{Q}})\backslash{\mathsf{V}}({\mathbb{A}})}\varphi(\mathsf{n}(X)\left[\begin{smallmatrix}\alpha 1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]g)\,\psi_{T_{\theta}}(\mathsf{n}(X))^{-1}\,{\rm{d}}X,\quad g\in\mathsf{G}({\mathbb{A}}).$
Proof.
The proof is straightforward. ∎
4. Automorphic forms and Fourier coefficients
Let $({\mathbb{Z}}^{2})_{\rm dom}:=\{\lambda=(l_{1},l_{2})\in{\mathbb{Z}}^{2}\mid l_{1}\geqslant l_{2}\,\}$. Let $\lambda=(l_{1},l_{2})\in({\mathbb{Z}}^{2})_{\rm dom}$ and $\varrho$ the representation of ${\bf GL}_{2}({\mathbb{C}})$ on the space $V_{\varrho}$ of homogeneous polynomials in $X,Y$ of degree $l_{1}-l_{2}$ defined as $\varrho(h)f(X,Y)=(\det h)^{l_{2}}f(aX+bY,cX+dY)$ for $h=\left[\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right]\in{\bf GL}_{2}({\mathbb{C}})$ and $f\in V_{\varrho}$. As is well-known, any irreducible rational representation of ${\bf GL}_{2}({\mathbb{C}})$, up to equivalence, is obtained this way. The space $V_{\varrho}$ carries an ${\bf SU}(2)$-invariant hermitian inner product product $(\cdot\mid\cdot)_{\varrho}$ given by [10, ((8.2.6)]. Recall ${\bf K}_{\infty}=\{k_{\infty}(u):=\left[\begin{smallmatrix}A&B\\ -B&A\end{smallmatrix}\right]\mid u=A+iB\in{\bf U}(2)\}$. For $N\in{\mathbb{Z}}_{>0}$, let $S_{\varrho}({\bf K}_{0}(N))$ denotes the space of smooth functions $\varphi:\mathsf{G}({\mathbb{A}})\rightarrow V_{\varrho}$ that satisfies the following conditions:
  • (i)
    $\varphi(z\gamma g)=\varphi(g)$ for $(z,\gamma,g)\in{\mathsf{Z}}({\mathbb{A}})\times\mathsf{G}({\mathbb{Q}})\times\mathsf{G}({\mathbb{A}})$.
  • (ii)
    $\varphi(gk_{\rm{\bf{f}}}k_{\infty}(u))=\varrho(\bar{u})^{-1}\varphi(g)$ for $k_{\rm{\bf{f}}}\in{\bf K}_{0}(N)$ and $k_{\infty}(u)\in{\bf K}_{\infty}$.
  • (iii)
    $R(X)\varphi=0$ for all $X\in{\mathfrak{p}}^{-}$.
  • (iv)
    $\varphi$ is bounded on $\mathsf{G}({\mathbb{A}})$.
For $T\in{\mathsf{V}}({\mathbb{R}})$ with $\det(T)\not=0$, let ${\bf B}_{\varrho}^{T}:\mathsf{G}({\mathbb{R}})^{0}\longrightarrow{\rm End}_{{\mathbb{C}}}(V_{\varrho})$ be a function defined by
(4.1) ${\bf B}_{\lambda}^{T}(g):=\nu(g)^{(l_{1}+l_{2})/2}\varrho(Ci+D)^{-1}\exp(2\pi i{\operatorname{tr}}(T\,g\langle i1_{2}\rangle),\quad g=\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in\mathsf{G}({\mathbb{R}})^{0}.$
This function satisfies the conditions:
(4.2) ${\bf B}^{T}_{\varrho}(\mathsf{m}(A,c)g)$ $={\bf B}^{T}_{\varrho}(g)\circ\varrho({}^{t}A^{-1}\,c)^{-1},\quad\mathsf{m}(A,c)\in{\mathsf{M}}_{T}^{\circ}({\mathbb{R}}),\,g\in\mathsf{G}({\mathbb{R}})^{0}$
(4.3) ${\bf B}^{T}_{\varrho}(ngk_{\infty}(u))$ $=\psi_{T}(n)\varrho(\bar{u})^{-l}\circ{\bf B}(g),\quad(n,g,u)\in\mathsf{N}({\mathbb{R}})\times\mathsf{G}({\mathbb{R}})^{0}\times{\bf U}(2),$
(4.4) $R({\mathfrak{p}}^{-}){\bf B}^{T}_{\varrho}$ $=0.$
The function $B^{T}_{\varrho,v}:g\longmapsto{\bf B}_{\varrho}^{T}(g)(v)$ with $v\in V_{\varrho}-(0)$ is bounded on $\mathsf{G}({\mathbb{R}})^{0}$ if and only if $T\in{\mathsf{V}}({\mathbb{R}})^{+}$, where ${\mathsf{V}}({\mathbb{R}})^{+}$ denote the set of positive definite elements in ${\mathsf{V}}({\mathbb{R}})$. Set
$\gamma$ $:=\mathsf{m}(-1_{2},-1)\in\mathsf{M}({\mathbb{Q}}).$
Note that $\nu(\gamma)=-1$.
Lemma 4.1.
Let $\varphi\in S_{\varrho}({\bf K}_{0}(N))$. There exists a unique family of vectors $a_{\varphi}(T;g_{\rm{\bf{f}}})\in V_{\varrho}$ $(T\in{\mathsf{V}}({\mathbb{Q}}))$, $g_{\rm{\bf{f}}}\in\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}})$ such that
(4.5) $\varphi(g_{\infty}g_{\rm{\bf{f}}})=\sum_{T\in{\mathsf{V}}({\mathbb{Q}})}{\bf B}_{\varrho}^{T}(g_{\infty})(a_{\varphi}(T;g_{\rm{\bf{f}}}))\,,\quad g_{\infty}\in\mathsf{G}({\mathbb{R}})^{\circ},\,g_{\rm{\bf{f}}}\in\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}}).$
If $T\not\in{\mathsf{V}}({\mathbb{Q}})^{+}$, then $a_{\varphi}(T;g_{\rm{\bf{f}}})=0$ for all $g_{\rm{\bf{f}}}\in\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}})$. For $g_{\rm{\bf{f}}}\in\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}})$,
(4.6) $\int_{\mathsf{N}({\mathbb{A}})/\mathsf{N}({\mathbb{Q}})}\varphi(ng_{\infty}g_{\rm{\bf{f}}})\,\psi_{T}(n)^{-1}{\rm{d}}n=\begin{cases}{\bf B}_{\varrho}^{T}(g_{\infty})(a_{\varphi}(T;g_{\rm{\bf{f}}})),\quad&(g_{\infty}\in\mathsf{G}({\mathbb{R}})^{\circ}),\\ 0,\quad&(g_{\infty}\in\mathsf{G}({\mathbb{R}})-\mathsf{G}({\mathbb{R}})^{\circ}).\end{cases}$
Proof.
Fix $g_{\rm{\bf{f}}}\in\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}})$ and examine the integral, say $W(g_{\infty})$, on the left-hand side of (4.6).The function $W$ on $\mathsf{G}({\mathbb{R}})^{0}$ satisfies conditions (4.3) and (4.4). As such, there corresponds a function $F:{\mathfrak{h}}_{2}\longrightarrow V_{\varrho}$ determined by the relation $F(g_{\infty}\langle i1_{2}\rangle)=\nu(g)^{-(l_{1}+l_{2})/2}\varrho(Ci+D)\,W(g_{\infty})$ with $g_{\infty}=\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in\mathsf{G}({\mathbb{R}})^{0}$ such that $Z=(Ai+B)(Ci+D)^{-1}$. Since $F(Z)$ satisfies $F(Z+X)=\psi_{T}(X)\,F(Z)$ ($X\in{\mathsf{V}}({\mathbb{R}}))$, it is of the form $F(Z)=F_{1}(Y)\exp(2\pi i{\operatorname{tr}}(TX))$ with a $V_{\varrho}$-valued $C^{\infty}$-function $F_{1}(Y)$. Since $F(Z)$ is holomorphic, the Cauchy-Riemann equations yield $\frac{{\rm{d}}}{{\rm{d}}y_{ij}}F_{1}(Y)=-2\pi t_{ij}F_{1}(T)$, which is uniquely solved as
$F_{1}(Y)=a_{\varphi}(T;g_{\rm{\bf{f}}})\exp(-2\pi{\operatorname{tr}}(TY))$
with a vector $a_{\varphi}(T;g_{\rm{\bf{f}}})\in V_{\varrho}$, or equivalently $W(g_{\infty})={\bf B}_{\varrho}^{T}(g_{\infty})(a_{\varphi}(T;g_{\rm{\bf{f}}}))$. Then, the formula in (4.5) is a consequence of the Fourier expansion of a $\mathsf{N}({\mathbb{Q}})$-periodic function on $\mathsf{N}({\mathbb{A}})$. Since $\varphi$ is bounded on $\mathsf{G}({\mathbb{A}})$, the function $W(g_{\infty})$ should be bounded on $\mathsf{G}({\mathbb{R}})^{0}$; since $g_{\infty}\mapsto{\bf B}_{\varrho}^{T}(g_{\infty})(v)$ with $v\in V_{\varrho}-\{0\}$ is unbounded if $T\not\in{\mathsf{V}}({\mathbb{R}})^{+}$, we have $a_{\varphi}(T;g_{\rm{\bf{f}}})=0$ for all $T\in{\mathsf{V}}({\mathbb{Q}})-{\mathsf{V}}({\mathbb{R}})^{+}$. It remains to show the second case in (4.6). Let $g_{\infty}\not\in\mathsf{G}({\mathbb{R}})^{\circ}$. Decompose $\gamma\in\mathsf{M}({\mathbb{Q}})$ as $\gamma_{\infty}\gamma_{{\rm{\bf{f}}}}$ $(\gamma_{\infty}\in\mathsf{M}({\mathbb{R}}),\,\gamma_{\rm{\bf{f}}}\in\mathsf{M}({\mathbb{A}}_{\rm{\bf{f}}}))$ in $\mathsf{M}({\mathbb{A}})$. By the left $\mathsf{G}({\mathbb{Q}})$-invariance of $\varphi$,
$\varphi(\mathsf{n}(X)g_{\infty}g_{\rm{\bf{f}}})=\varphi(\gamma\mathsf{n}(X)g_{\infty}g_{\rm{\bf{f}}})=\varphi(\mathsf{n}(-X)\gamma_{\infty}g_{\infty}\gamma_{\rm{\bf{f}}}g_{\rm{\bf{f}}}).$
Integrate this in $X$ over ${\mathsf{V}}({\mathbb{Q}})\backslash{\mathsf{V}}({\mathbb{A}})$ and make a change of variables $n\rightarrow n^{-1}$ on the way; we have
$\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(ng_{\infty}g_{\rm{\bf{f}}})\,\psi_{T}(n)^{-1}{\rm{d}}n$ $=\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(n^{-1}\gamma_{\infty}g_{\infty}\gamma_{\rm{\bf{f}}}g_{\rm{\bf{f}}})\,\psi_{T}(n)\,{\rm{d}}n$
$=\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(n\gamma_{\infty}g_{\infty}\gamma_{\rm{\bf{f}}}g_{\rm{\bf{f}}})\,\psi_{-T}(n)^{-1}\,{\rm{d}}n.$
Since $\gamma_{\infty}g_{\infty}\in\mathsf{G}({\mathbb{R}})^{\circ}$, the last integral becomes ${\bf B}_{\varrho}^{-T}(\gamma_{\infty}g_{\infty})(a_{\varphi}(-T;\gamma_{\rm{\bf{f}}}g_{\rm{\bf{f}}}))$ by the first case. Since $-T\not\in{\mathsf{V}}({\mathbb{R}})^{+}$, we have $a(-T;\gamma_{\rm{\bf{f}}}g_{\rm{\bf{f}}})=0$. ∎
The vectors $a_{\varphi}(T;g_{\rm{\bf{f}}})$ in $V_{\varrho}$ are referred to as the adelic Fourier coefficients of $\varphi$. As is well-known, there is a linear bijective correspondence between $\varphi\in S_{\varrho}({\bf K}_{0}(N))$ and the classical $V_{\varrho}$-valued Siegel cusp forms $f(Z)$ on ${\bf\Gamma}_{0}(N):=\{\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in{\bf Sp}_{2}({\mathbb{Z}})\mid C\equiv 0\pmod{N}\}$ determined by the relation
$f(g_{\infty}\langle i1_{2}\rangle)=\nu(g_{\infty})^{-(l_{1}+l_{2})/2}\varrho(Ci+D)\varphi(g_{\infty}),\quad g_{\infty}=\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in\mathsf{G}({\mathbb{R}})^{0}.$
By the modular transformation law $f((AZ+B)(CZ+D)^{-1})=\varrho(CZ+D)^{-1}f(Z)$ for $\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in\Gamma_{0}(N)$, since $-1_{4}\in\Gamma_{0}(N)$, we have that $f(Z)$ is identically 00, hence $S_{\varrho}({\bf K}_{0}(N))=(0)$, unless $l_{1}\equiv l_{2}\pmod{2}$. (For more details, see [37, §3.2] and [4, §4].)
Since $\mathsf{N}({\mathbb{A}})\cap{\bf K}_{0}(N)=\{\mathsf{n}(X)\mid X\in{\mathsf{V}}(\widehat{\mathbb{Z}})\}$, by (4.6), we have $a_{\varphi}(T;1)=0$ unless $X\in{\mathcal{Q}}$, where
${\mathcal{Q}}:=\{\left[\begin{smallmatrix}a&b/2\\ b/2&c\end{smallmatrix}\right]\mid a,b,c\in{\mathbb{Z}}\}$
is the dual lattice of ${\mathsf{V}}({\mathbb{Z}})$. Then, (4.5) with $g_{\rm{\bf{f}}}=1_{4}$ reduces to
$f(Z)=\sum_{T\in{\mathcal{Q}}^{+}}a_{\varphi}(T;1_{4})\,\exp(2\pi i{\operatorname{tr}}(TZ)),\quad Z\in{\mathfrak{h}}_{2},$
where ${\mathcal{Q}}^{+}:={\mathcal{Q}}\cap{\mathsf{V}}({\mathbb{Q}})^{+}$. This means $a_{\varphi}(T):=a_{\varphi}(T;1_{4})$ $(T\in{\mathcal{Q}}^{+})$ is the Fourier coefficient of $f(Z)$ in the classical sense. Set
$\delta:=\left[\begin{smallmatrix}1&0\\ 0&-1\end{smallmatrix}\right]\in{\bf GL}_{2}({\mathbb{Q}}).$
Note that the map $T\mapsto-T{\mathsf{s}}(\delta)$ preserves the set ${\mathcal{Q}}^{+}$. If we identify ${\mathcal{Q}}^{+}$ with the set of positive definite integral binary forms $aX^{2}+bXY+cY^{2}$, this operation corresponds to the sign change of $b$.
Lemma 4.2.
The set of adelic Fourier coefficients of $\varphi\in S_{\varrho}({\bf K}_{0}(N))$ has the following properties:
(4.7) $a_{\varphi}(T;\mathsf{m}(h_{\rm{\bf{f}}},\det h_{\rm{\bf{f}}})g_{\rm{\bf{f}}}\kappa)=a_{\varphi}(T{\mathsf{s}}(h);g_{\rm{\bf{f}}}),\quad h\in{\bf GL}_{2}({\mathbb{Q}}),\,g_{\rm{\bf{f}}}\in\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}}),\,\kappa\in{\bf K}_{0}(N).$
(4.8) $a_{\varphi}(-T{\mathsf{s}}\left(\delta\right);1)=\varrho\left(\delta\right)(a_{\varphi}(T;1)),\quad T\in{\mathsf{V}}({\mathbb{Q}}).$
Proof.
The relation in (4.6) follows from (4.7) by a simple change of variables. Let us show (4.8). We have
(4.9) $\mathsf{n}\left(-X{\mathsf{s}}\left(\delta\right)\right)=\kappa\mathsf{n}(X)\kappa^{-1},\quad X\in{\mathsf{V}}({\mathbb{A}})$
with $\kappa:=\mathsf{m}(\delta,1)\in\mathsf{G}({\mathbb{Q}})$. Write $\kappa=\kappa_{\rm{\bf{f}}}\kappa_{\infty}\in\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}})\,\mathsf{G}({\mathbb{R}})$; then, $\kappa_{\rm{\bf{f}}}\in{\bf K}_{0}(N)$ and $\kappa_{\infty}\in{\bf K}_{\infty}$. We have
$\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(n)\,\psi_{-T{\mathsf{s}}\left(\delta\right)}(n)^{-1}\,{\rm{d}}n$ $=\int_{{\mathsf{V}}({\mathbb{Q}})\backslash{\mathsf{V}}({\mathbb{A}})}\varphi(\mathsf{n}(X))\,\psi({\operatorname{tr}}(-(T{\mathsf{s}}\left(\delta\right))\,X))^{-1}\,{\rm{d}}X$
$=\int_{{\mathsf{V}}({\mathbb{Q}})\backslash{\mathsf{V}}({\mathbb{A}})}\varphi(\mathsf{n}(-{\mathsf{s}}\left(\delta\right)X))\,\psi({\operatorname{tr}}(TX))^{-1}\,{\rm{d}}X$
$=\int_{{\mathsf{V}}({\mathbb{Q}})\backslash{\mathsf{V}}({\mathbb{A}})}\varphi(\kappa\mathsf{n}(X)\kappa^{-1})\,\psi({\operatorname{tr}}(TX))^{-1}\,{\rm{d}}X$
$=\varrho(\delta)\,\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(n)\psi_{T}(n)^{-1}\,{\rm{d}}n.$
From this and (4.6),
${\bf B}_{\varrho}^{-T{\mathsf{s}}\left(\delta\right)}(1_{4})(a_{\varphi}(-T{\mathsf{s}}\left(\delta\right);1))=\varrho(\delta)\,{\bf B}_{\varrho}^{T}(1_{4})(a_{\varphi}(T;1)).$
By (4.1), ${\bf B}_{\varrho}^{-T{\mathsf{s}}\left(\delta\right)}(1_{4})={\bf B}_{\varrho}^{T}(1_{4})=e^{-2\pi{\operatorname{tr}}(T)}{\rm Id}_{V_{\varrho}}$. ∎
Define
(4.10) ${\mathcal{Q}}^{+}_{\rm prim}(D):=\{T=\left[\begin{smallmatrix}a&2^{-1}b\\ 2^{-1}b&c\end{smallmatrix}\right]\in{\mathcal{Q}}\mid a>0,\,-\det T=D/4\}.$
Since $D$ is a fundamental discriminant, $T\in{\mathcal{Q}}^{+}_{\rm prim}(D)$ is primitive, i.e., ${\gcd}(a,b,c)=1$; ${\mathcal{Q}}^{+}_{\rm prim}(D)$ is a subset of ${\mathsf{V}}({\mathbb{Q}})$, which is stable under the action of the unimodular group ${\bf SL}_{2}({\mathbb{Z}})$ induced by (2.1). The matrix $T_{\theta}$ defined by (2.11) belongs to ${\mathcal{Q}}_{\rm prim}^{+}(D)$. The ${\bf SL}_{2}({\mathbb{Z}})$-orbit of $T\in{\mathcal{Q}}^{+}_{\rm prim}(D)$ is denoted by $[T]$.
Lemma 4.3.
For each $u\in{\mathbb{A}}_{E}^{\times}$, let $I_{\theta}(u)$ decompose in ${\bf GL}_{2}({\mathbb{A}})$ as
(4.11) $I_{\theta}(u)=\gamma_{u}h_{u}\kappa_{u},\quad\gamma_{u}\in{\bf GL}_{2}({\mathbb{Q}}),\,h_{u}\in{\bf GL}_{2}({\mathbb{R}})^{\circ},\,\kappa_{u}=(\kappa_{u,p})_{p<\infty}\in{\bf GL}_{2}(\widehat{\mathbb{Z}})$
and set $T_{\theta}(u):=T_{\theta}\,{\mathsf{s}}(\gamma_{u})$. Then ,
  • (i)
    $T_{\theta}(u)\in{\mathcal{Q}}_{\rm prim}^{+}(D)$.
  • (ii)
    The ${\bf SL}_{2}({\mathbb{Z}})$-equivalence class of $T_{\theta}(u)$ does not depend on the decomposition (4.11). If $u$, $u^{\prime}$ belongs to the same $E^{\times}E^{\times}_{\infty}\widehat{\mathfrak{o}}_{E}^{\times}$-coset, then $T_{\theta}(u)$ and $T_{\theta}(u^{\prime})$ are ${\bf SL}_{2}({\mathbb{Z}})$-equivalent.
  • (iii)
    The map $[u]\mapsto[T_{\theta}(u)]$ from ${\rm Cl}(E)={\mathbb{A}}_{E}^{\times}/E^{\times}E_{\infty}^{\times}\widehat{\mathfrak{o}}_{E}^{\times}$ to ${\mathcal{Q}}_{\rm prim}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ is a bijection.
Proof.
This is an adelic reformulation of the classically well-known correspondence between the ideal classes and the ${\bf SL}_{2}({\mathbb{Z}})$-equivalence classes of integral binary quadratic forms. The proof is straight-forward. ∎
Since $\mathsf{m}({\bf GL}_{2}(\widehat{\mathbb{Z}}),1)\subset{\bf K}_{0}(N)$, the relation in (4.7) shows that the function $T\mapsto a_{\varphi}(T):=a_{\varphi}(T;1_{4})$ on ${\mathsf{V}}({\mathbb{Z}})^{+}$ is ${\bf SL}_{2}({\mathbb{Z}})$-invariant. Hence the following sum of vectors in $V_{\varrho}$ is well-defined:
(4.12) ${\bf R}(\varphi,E,\Lambda):=\sum_{[u]\in{\rm Cl}(E)}a_{\varphi}([T_{\theta}(u)])\,\Lambda(u)^{-1}.$
Recall the integral in (3.6) and the function in (4.1).
Lemma 4.4.
Let $\varphi\in S_{\varrho}({\bf K}_{0}(N))$.
  • (i)
    For $g_{\infty}\in\mathsf{G}({\mathbb{R}})^{0}$,
    $B^{T_{\theta},\Lambda}(\varphi;g_{\infty})=w_{D}^{-1}\,{\bf B}_{\varrho}^{T_{\theta}}(g_{\infty})({\bf R}(\varphi,E,\Lambda)).$
  • (ii)
    We have
    (4.13) $\varrho(r)(B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta}))$ $=B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta}),\quad r\in{\bf SO}(2),$
    (4.14) $\varrho(\delta)(B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta}))$ $=B^{-T_{\theta}{\mathsf{s}}(\sigma),\Lambda}(\varphi;b_{\mathbb{R}}^{\theta}),$
    where $\sigma:=\left[\begin{smallmatrix}1&t\\ 0&-1\end{smallmatrix}\right]$.
Proof.
(i) The set ${\mathbb{A}}_{E}^{\times}/E^{\times}E_{\mathbb{R}}^{\times}$ is decomposed to a disjoint union of $\widehat{\mathfrak{o}}_{E}^{\times}$-orbits $[u]\widehat{\mathfrak{o}}_{E}^{\times}$ for different classes $[u]\in{\rm Cl}(E)$. Let $\mu$ denote the quotient measure on ${\mathbb{A}}_{E}^{\times}/E^{\times}E_{\mathbb{R}}^{\times}$. For $\tau\in{\mathbb{A}}_{E}^{\times}$, set $\mathsf{m}_{\theta}(\tau):=\mathsf{m}(I_{\theta}(\tau),{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau))$. Then by (3.6) and (4.6),
$B^{T_{\theta},\Lambda}(\varphi;g_{\infty})$ $={\bf B}_{\varrho}^{T_{\theta}}(g_{\infty})\Bigl{(}\int_{{\mathbb{A}}_{E}^{\times}/E^{\times}E_{\infty}^{\times}}a_{\varphi}(T_{\theta};\mathsf{m}_{\theta}(u))\,\Lambda(u)^{-1}\,{\rm{d}}\mu(u)\Bigr{)}$
$={\bf B}_{\varrho}^{T_{\theta}}(g_{\infty})\Bigl{(}\sum_{[u]\in{\rm Cl}(E)}\int_{[u]\widehat{\mathfrak{o}}_{E}^{\times}}a_{\varphi}(T_{\theta};\mathsf{m}_{\theta}(u))\,\Lambda(u)^{-1}\,{\rm{d}}\mu(u)\Bigr{)}.$
Since $\mathsf{m}_{\theta}(\widehat{\mathfrak{o}}_{E}^{\times})\subset{\bf K}_{0}(N)$, by (4.6), the formula in the parenthesis becomes
$\sum_{[u]\in{\rm Cl}(E)}\mu([u]\widehat{\mathfrak{o}}_{E}^{\times})\,a_{\varphi}(T_{\theta};\mathsf{m}_{\theta}(u))\,\Lambda(u)^{-1}.$
Let $I_{\theta}(u)$ decompose as in (4.11). Then, by (4.7) and Lemma 4.3, we have
$a_{\varphi}(T_{\theta};\mathsf{m}_{\theta}(u))=a_{\varphi}(T_{\theta}{\mathsf{s}}(\gamma_{u});1)=a_{\varphi}(T_{\theta}(u)).$
It remains to compute $\mu([u]\widehat{\mathfrak{o}}_{E}^{\times})$. Since any fiber of the natural surjection from $\widehat{\mathfrak{o}}_{E}^{\times}$ onto $[u]\widehat{\mathfrak{o}}_{E}^{\times}$ has a simply transitive action of the global unit group $\mathfrak{o}_{E}^{\times}$, which coincides with the set of roots of unity in $E^{\times}$, we have $\mu([u]\widehat{\mathfrak{o}}_{E}^{\times})=(1/\#\mathfrak{o}_{E}^{\times})=1/w_{D}$.
(ii) We have $I_{\theta}(e^{i\alpha})=A_{\theta}^{-1}\left[\begin{smallmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\end{smallmatrix}\right]A_{\theta}$, or equivalently
$\mathsf{m}(I_{\theta}(e^{i\alpha}),1)=b_{\mathbb{R}}^{\theta}{\operatorname{diag}}(\left[\begin{smallmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\end{smallmatrix}\right],\left[\begin{smallmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\end{smallmatrix}\right])(b_{\mathbb{R}}^{\theta})^{-1}$
for $\alpha\in{\mathbb{R}}$. Due to this and $\Lambda_{\infty}=1$, formula (4.13) is proved by a change of variables $\tau\rightarrow\tau u^{-1}$ in the integral (3.6). Since $b_{\mathbb{R}}^{\theta}\mathsf{m}(\delta,1)(b_{\mathbb{R}}^{\theta})^{-1}=\mathsf{m}(\sigma,1)$, formula (4.14) is proved in the same way as the proof of Lemma 4.2. ∎
Let $\varphi\in S_{\varrho}({\bf K}_{0}(N))-(0)$; then $l_{1}\equiv l_{2}\pmod{2}$ as noted above. As such, $\dim_{\mathbb{C}}(V_{\varrho}^{{\bf SO}(2)})=1$; indeed, $V_{\varrho}^{{\bf SO}(2)}={\mathbb{C}}\,\mathsf{v}_{\varrho}^{0}$ with $\mathsf{v}_{\varrho}^{0}:=(X^{2}+Y^{2})^{(l_{1}-l_{2})/2}$. Set
$\mathsf{v}_{\varrho}^{\theta}:=\bigl{(}\tfrac{-2}{\sqrt{|D|}}\bigr{)}^{\frac{l_{1}+l_{2}}{2}}\,\varrho({}^{t}A_{\theta})\mathsf{v}_{\varrho}^{0}=\bigl{(}\tfrac{-2}{\sqrt{|D|}}\bigr{)}^{\frac{l_{1}-l_{2}}{2}}(X^{2}+tXY+NY^{2})^{\frac{l_{1}-l_{2}}{2}}.$
Note that ${\bf B}_{\varrho}^{T_{\theta}}(b_{\mathbb{R}}^{\theta})=(\frac{-2}{\sqrt{|D|}})^{-(l_{1}+l_{2})/2}e^{-2\pi\sqrt{|D|}}\,\varrho({}^{t}A_{\theta})^{-1}$, so that ${\bf B}_{\varrho}^{T_{\theta}}(b_{\mathbb{R}}^{\theta})(\mathsf{v}_{\varrho}^{\theta})=e^{-2\pi\sqrt{|D|}}\mathsf{v}_{\varrho}^{0}$. Thus, as a corollary to Lemma 4.4, the vector ${\bf R}(\varphi,E,\Lambda)\in V_{\varrho}$ is non-zero if and only if $B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta})\not=0$, in which case there exists a unique scalar $R(\varphi,E,\Lambda)$ such that ${\bf R}(\varphi,E,\Lambda)=R(\varphi,E,\Lambda)\,\mathsf{v}_{\varrho}^{\theta}$, or equivalently
(4.15) $B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta})=w_{D}^{-1}e^{-2\pi\sqrt{|D|}}\,R(\varphi,E,\Lambda)\mathsf{v}_{\varrho}^{0}.$
Define a $C^{\infty}$-function $B_{\varrho}^{T_{\theta}}:\mathsf{G}({\mathbb{R}})^{0}\longrightarrow V_{\varrho}$ as
(4.16) $B^{T_{\theta}}_{\varrho}(g_{\infty}):={\bf B}^{T_{\theta}}_{\varrho}(g_{\infty})(\mathsf{v}_{\varrho}^{\theta}),\quad g_{\infty}\in\mathsf{G}({\mathbb{R}})^{0}.$
4.1. Sign condition
The Galois group ${\rm Gal}(E/{\mathbb{Q}})$ acts on ${\rm Cl}(E)$ naturally, hence on the orbit space ${{\mathcal{Q}}}_{\rm prim}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ through the bijection in Lemma 4.3(iii). The following lemma describes the conjugate action on ${{\mathcal{Q}}}_{\rm prim}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ explicitly. Recall the element $\sigma\in{\bf GL}_{2}({\mathbb{Z}})$ defined in Lemma 4.4 (ii).
Lemma 4.5.
For $u\in{\rm Cl}_{E}$, the element $T_{\theta}(\bar{u})$ is ${\bf SL}_{2}({\mathbb{Z}})$-equivalent to $-T_{\theta}(u)\,{\mathsf{s}}\left(\delta\right)$. In other words, the conjugate action on ${\mathcal{Q}}_{\rm prim}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ is induced by the map
$T=\left[\begin{smallmatrix}a&2^{-1}b\\ 2^{-1}b&c\end{smallmatrix}\right]\longmapsto-T\,{\mathsf{s}}(\delta)=\left[\begin{smallmatrix}a&-2^{-1}b\\ -2^{-1}b&c\end{smallmatrix}\right].$
Proof.
The defining formula of $I_{\theta}$ in (2.5) yields
$I_{\theta}(a+\theta b)=\left[\begin{smallmatrix}1&\theta\\ 1&\bar{\theta}\end{smallmatrix}\right]^{-1}\left[\begin{smallmatrix}a+\theta b&0\\ 0&a+\bar{\theta}b\end{smallmatrix}\right]\left[\begin{smallmatrix}1&\theta\\ 1&\bar{\theta}\end{smallmatrix}\right],\quad a,b\in{\mathbb{Q}}.$
Set $t:={\operatorname{tr}}_{E/{\mathbb{Q}}}(\theta)$. Then, by $\bar{\theta}=t-\theta$, we have $a+b\bar{\theta}=a^{\prime}+b^{\prime}\theta$ with $a^{\prime}=a+tb\in{\mathbb{Q}}$ and $b^{\prime}=-b\in{\mathbb{Q}}$. Then , a computation reveals
(4.17) $I_{\theta}(a+b\bar{\theta})$ $=\sigma I_{\theta}(a+b\theta)\sigma^{-1}.$
By the decomposition of $I_{\theta}(u)$ for $u\in{\mathbb{A}}_{E}^{\times}$ in (4.11), we may take $\gamma_{\bar{u}}=\sigma\gamma_{u}\sigma^{-1}$. Thus,
$T_{\theta}(\bar{u})=T_{\theta}{\mathsf{s}}(\gamma_{\bar{u}})=T_{\theta}{\mathsf{s}}\left(\sigma\gamma_{u}\sigma^{-1}\right).$
A computation shows $T_{\theta}{\mathsf{s}}\left(\sigma\right)=-T_{\theta}$. Hence,
$T_{\theta}(\bar{u})=-T_{\theta}{\mathsf{s}}\left(\gamma_{u}\sigma^{-1}\right)=-T_{\theta}(u){\mathsf{s}}\left(\sigma^{-1}\right)=-T_{\theta}(u){\mathsf{s}}(\delta){\mathsf{s}}\left(\left[\begin{smallmatrix}1&t\\ 0&1\end{smallmatrix}\right]\right).$
The last matrix is ${\bf SL}_{2}({\mathbb{Z}})$-equivalent to $-T_{\theta}(u){\mathsf{s}}(\delta)$. This completes the proof. ∎
For $\Lambda\in\widehat{{\rm Cl}(E)}$, recall its conjugate $\Lambda^{\dagger}\in\widehat{{\rm Cl}(E)}$ (see (2.25)).
Lemma 4.6.
Let $\varphi\in S_{\varrho}({\bf K}_{0}(N))$ and $\Lambda\in\widehat{{\rm Cl}(E)}$ and suppose $B^{T_{\theta},\Lambda}(\varphi;b_{\mathbb{R}}^{\theta})\not=0$. Then,
$R(\varphi,E,\Lambda^{\dagger})=(-1)^{l_{2}}\,R(\varphi,E,\Lambda).$
Proof.
This is proved by Lemma 4.5 and (4.15) and by $\varrho(\delta)\mathsf{v}_{\varrho}^{0}=(-1)^{l_{2}}\mathsf{v}_{\varrho}^{0}$. Note that $\delta^{-1}\sigma=\left[\begin{smallmatrix}1&t\\ 0&1\end{smallmatrix}\right]\in{\bf SL}_{2}({\mathbb{Z}})$, so that $a_{\varphi}(-T_{\theta}{\mathsf{s}}(\sigma),1)=a_{\varphi}(-T_{\theta}{\mathsf{s}}(\delta),1)$. ∎
4.2. Automorphic representations
Let $\Pi_{\rm cusp}({\mathsf{Z}}\backslash\mathsf{G})$ denote the set of all those irreducible cuspidal representation of $\mathsf{G}({\mathbb{A}})$ with trivial central characters, i.e., those irreducible $({\mathfrak{g}},{\bf K}_{\infty})\times\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}})$-submodules $(\pi,V_{\pi})$ of the space of cusp forms on ${\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})$. We endow $V_{\pi}$ with the restriction of the $L^{2}$-inner product
$(\varphi|\varphi_{1})_{L^{2}}:=\int_{{\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})}\varphi(g)\overline{\varphi_{1}(g)}\,{\rm{d}}g,\qquad\varphi,\,\varphi_{1}\in V_{\pi}).$
For $\pi\in\Pi_{\rm cusp}({\mathsf{Z}}\backslash\mathsf{G})$, we fix its restricted tensor decomposition $\pi\cong\otimes_{p\leqslant\infty}\pi_{p}$ with $(\pi_{p},V_{\pi_{p}})$ being an irreducible admissible unitarizable representation of $\mathsf{G}({\mathbb{Q}}_{p})$ with trivial central character such that a ${\bf K}_{p}$-invariant vector $\xi_{p}^{0}\in V_{\pi_{p}}-(0)$ is preassigned for almost all $p<\infty$. An element $\lambda\in({\mathbb{Z}}^{2})_{\rm dom}$ is the highest weight of the minimal ${\bf K}_{\infty}$-type of a holomorphic discrete series representation (HDS for short) of ${\bf Sp}_{2}({\mathbb{R}})$ if and only if $l_{2}>2$, in which case the Harish-Chandra parameter of the HDS is $(l_{1}-1,l_{2}-2)$. For such $\lambda$ and $N\in{\mathbb{Z}}_{>0}$, let $\Pi_{\rm cusp}(\lambda,N)$ be a subset of $\pi\in\Pi_{\rm cusp}({\mathsf{Z}}\backslash\mathsf{G})$ having the following properties:
  • (i)
    As a $({\mathfrak{g}},{\bf K}_{\infty})$-module $\pi_{\infty}\cong D_{l_{1}-1,l_{2}-2}\oplus D_{-l_{2}+2,-l_{1}+1}$, where $D_{m_{1},m_{2}}$ is the discrete series representation of $\mathsf{G}({\mathbb{R}})^{0}$ of Harish-Chandra parameter $(m_{1},m_{2})$ with central character ${\mathsf{Z}}({\mathbb{R}})\cong{\mathbb{R}}^{\times}\ni z\longmapsto{\rm sgn}(z)^{m_{1}+m_{2}+1}$.
  • (ii)
    $V_{\pi}^{{\bf K}_{0}(N)}\not=(0)$.
  • (iii)
    Let $(\varrho,V_{\varrho})$ be the irreducible rational representation of highest weight $\lambda$ as before. The space $S_{\varrho}({\bf K}_{0}(N))$ is an orthogonal direct sum of spaces $V_{\pi}^{{\bf K}_{0}(N)}[\varrho]:=\{\varphi\in S_{\varrho}({\bf K}_{0}(N))\mid\varphi_{v}\in V_{\pi}\,(\forall v\in V_{\varrho})\}$ for $\pi\in\Pi_{\rm cusp}(\lambda,N)$, where $\varphi_{v}(g):=(\varphi(g)\mid v)_{\varrho}$ for $\varphi\in S_{\varrho}({\bf K}_{0}(N))$ and $v\in V_{\varrho}$.
Although there may be many choices of $\Pi_{\rm cusp}(\lambda,N)$, we fix one of them once and for all. For $\pi\in\Pi_{\rm cusp}(\lambda,N)$, we fix a ${\bf K}_{\infty}$-intertwining map $V_{\varrho}\oplus\overline{V_{\varrho}}\hookrightarrow D_{l_{1}-1,l_{2}-2}\oplus D_{-l_{2}+2,-l_{1}+1}\,(\cong\pi_{\infty})$ once and for all, where $\overline{V_{\varrho}}:=\overline{{\mathbb{C}}}\otimes_{{\mathbb{C}}}V_{\varrho}$ is the complex conjugate of $V_{\varrho}$.
4.3. Basic assumptions
From now on, we fix a pair $(\lambda=(l_{1},l_{2}),N)\in({\mathbb{Z}}^{2})_{\rm dom}\times{\mathbb{Z}}_{>0}$ and a character $\mu=\otimes_{p\leqslant\infty}\mu_{p}:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}{\mathbb{R}}^{\times}\to{\mathbb{C}}^{1}$ with $M:={\rm cond}(\mu)$ satisfying the following conditions:
  • (A-i)
    $N$ is square free.
  • (A-ii)
    $l_{1}\equiv l_{2}\pmod{2}$, and $l_{2}\in{\mathbb{Z}}_{\geqslant 3}$ so that $(l_{1}-1,l_{2}-2)$ is the Harish-Chandra parameter of a HDS.
  • (A-iii)
    $M$ and $N$ are coprime.
  • (A-iv)
    $M$ is odd.
  • (A-v)
    All prime divisors of $NM$ are inert in $E/{\mathbb{Q}}$.
As usual, the character $\mu$ induces an even primitive Dirichlet character $\tilde{\mu}:({\mathbb{Z}}/M{\mathbb{Z}})^{\times}\rightarrow{\mathbb{C}}^{1}$ in such a way that $\mu(u)=\tilde{\mu}(a)^{-1}$ for all $u\in\widehat{\mathbb{Z}}^{\times}$ and $a\in{\mathbb{Z}}$ with $u-a\in M\widehat{\mathbb{Z}}$ ; thus $\mu_{p}(p)=\tilde{\mu}(p)$ for primes $p\nmid M$, and
$\prod_{p\mid M}W_{{\mathbb{Q}}_{p}}(\mu_{p},\psi_{p})=M^{-1/2}G(\tilde{\mu}),$
with $W_{{\mathbb{Q}}_{p}}(\mu_{p},\psi_{p})$ being as in Definition 2 and $G(\tilde{\mu}):=\sum_{a\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}}\tilde{\mu}(a)e^{2\pi ia/M}$, the Gauss sum of $\tilde{\mu}$. Let $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ denote the set of all those $\pi\in\Pi_{\rm cusp}(\lambda,N)$ that satisfies the condition
  • (A-vi)
    $\pi$ admits a global $(T_{\theta},\Lambda)$-Bessel model, i.e., there exists $\varphi\in V_{\pi}$ such that $B^{T_{\theta},\Lambda}(\varphi;g)\not=0$ for some $g\in\mathsf{G}({\mathbb{A}})$.
We remark that the conditions listed above (when $\mu={\bf 1}$ and $l_{1}=l_{2}$) are also imposed in the most part of [7]. Let $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm SK}$ denote the set of $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ that is the Saito-Kurokawa lift ([29]) of an cuspidal automorphic representation of ${\bf PGL}_{2}({\mathbb{A}})$, which is locally described by [38] (see also [39]). From [38, §4], the set $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm SK}=\varnothing$ unless $l_{1}=l_{2}$, i.e., $V_{\varrho}$ is one dimensional. We note that all the representations $\pi\cong\otimes_{p}\pi_{p}$ in $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)\setminus\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm SK}$ are non-CAP (cf. [42], [18, §3.5], [34, Corollary 4.5]); then, by [48], $\pi_{p}$ is tempered for all $p\nmid N$. Moreover, by invoking [3] (see also [40]) , any $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)\setminus\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm SK}$ is either a Yoshida lift, i.e., there exists a pair of irreducible cuspidal automorphic representations $(\sigma_{1},\sigma_{2})$ of ${\bf GL}_{2}({\mathbb{A}})$ such that $L(s,\pi)=L(s,\sigma_{1})L(s,\sigma_{2})$, or a “general type”, that is, there exists an irreducible cuspidal automorphic representation $\Pi$ of ${\bf GL}_{4}({\mathbb{A}})$ such that $L(s,\Pi,\wedge^{2})$ has a pole at $s=1$ and $L(s,\pi)=L(s,\Pi)$. Let $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm Y}$ (resp. $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{\rm G}$) be the set of $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)$ that is a Yoshida lift (resp. a general type). Then,
[…]
On the other hand, the set $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ can be separated into the subsets consisting of all “newforms” (i.e., $N_{\pi}=N)$ and all “oldforms” (i.e., $N_{\pi}\not=N$) denoted by $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm new}$ and $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm old}$ respectively. For $\bullet\in\{{\rm G},{\rm Y},{\rm SK}\}$ and $\ast\in\{{\rm new},{\rm old}\}$, we set
$\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\bullet,\ast}=\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\bullet}\cap\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\ast}.$
Let $N_{\pi}$ be the product of primes $p<\infty$ such that $\pi_{p}$ is not spherical. By condition (ii) in §4.2, we have $N_{\pi}\mid N$. Due to (A-ii), the representations $\pi_{p}$ for $p\mid N_{\pi}$ have non-zero ${\bf K}_{0}(p{\mathbb{Z}}_{p})$-fixed vectors, which, in turn, implies that $\pi_{p}$ is Iwahori spherical so that their isomorphism classes are listed in [37, Table 3] (For extended tables, we refer to [35, Appendix].) Note that in [37] a different symplectic form is used to define the symplectic group; up to the adjustment for this difference, the group $P_{1}$ in [37] is our ${\bf K}_{0}(p{\mathbb{Z}}_{p})$. In [35], all irreducible admissible representations of $\mathsf{G}({\mathbb{Q}}_{p})$ that admit local Bessel models are classified, and the result is conveniently summarized in [32, Table 2]. Here is a summary of what is available for our $\pi$ (as in §4.3):
  • $\pi_{\infty}$ as a representation of $\mathsf{G}({\mathbb{R}})^{0}$ is a direct sum of a HDS and its complex conjugate; as such, $\pi$ is CAP if and only if it is a Saito-Kurokawa lift from a cuspidal representation of ${\bf PGL}_{2}({\mathbb{A}})$; which happens only if $l_{1}=l_{2}$.
  • Suppose $p\nmid N_{\pi}$. Then, the local representation $\pi_{p}$ is of type ${\rm I}$ and tempered if $\pi$ is non-CAP, and is of type ${\rm II}b$ when $\pi$ is CAP.
  • Suppose $p\mid N_{\pi}$. Then, the local representation $\pi_{p}$ is either of type ${\rm III}a$, in which case $\dim_{{\mathbb{C}}}V_{\pi_{p}}^{{\bf K}_{0}(p{\mathbb{Z}}_{p})}=2$, or of type ${\rm VI}b$, in which case $\dim_{{\mathbb{C}}}V_{\pi_{p}}^{{\bf K}_{0}(p{\mathbb{Z}}_{p})}=1$; when $\pi$ is CAP, $\pi_{p}$ has to be of type ${\rm VI}b$.
4.4. Bessel models
Let $p<\infty$; for any irreducible admissible representation $(\pi_{p},V_{\pi_{p}})$ of $\mathsf{G}({\mathbb{Q}}_{p})$, let $(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda}$ denote the space of all ${\mathbb{C}}$-linear forms $\ell:V_{\pi_{p}}\longrightarrow{\mathbb{C}}$ that satisfies
$\ell(\pi_{p}(\mathsf{m}(I_{\theta}(\tau),{\operatorname{N}}_{E/{\mathbb{Q}}}(\tau))n)\xi)=\Lambda(\tau)\psi_{T_{\theta}}(n)\,\ell(\xi),\quad\xi\in V_{\pi_{p}},\,\tau\in E_{p}^{\times},\,n\in\mathsf{N}({\mathbb{Q}}_{p}).$
It is known that $\dim_{{\mathbb{C}}}(V_{\pi}^{*})^{T_{\theta},\Lambda}\leqslant 1$ ([27], [28]). We say that $\pi_{p}$ has a local $(T_{\theta},\Lambda_{p})$-Bessel model if $(V_{\pi}^{*})^{T_{\theta},\Lambda}\not=(0)$; when this is the case, the space of functions of the form $g\mapsto\ell(\pi_{p}(g)\xi)$ with $\xi\in V_{\pi_{p}}$ is independent of $\ell\in(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda_{p}}-(0)$; this space is denoted by ${\mathscr{B}}(T_{\theta},\Lambda_{p})[\pi_{p}]$ and is called the local $(T_{\theta},\Lambda_{p})$-Bessel model of $\pi_{p}$. When $\pi_{p}$ is spherical, it is known that $\pi_{p}$ has a local $(T_{\theta},\Lambda_{p})$-Bessel model, and the space ${\mathscr{B}}(T_{\theta},\Lambda_{p})[\pi_{p}]$ contains a unique ${\bf K}_{p}$-invariant function $B_{\pi_{p}}$ such that $B_{\pi_{p}}^{0}(1_{4})=1$ ([43, Theorem 2-I], [6]). Let $\xi_{p}^{0}$ be the non-zero ${\bf K}_{p}$-fixed vector in $V_{\pi_{p}}$; then there exists a unique element $\ell_{\pi_{p}}^{0}\in(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda_{p}}-(0)$ such that
$\ell_{\pi_{p}}^{0}(\pi_{p}(g)\xi_{\pi_{p}}^{0})=B_{\pi_{p}}^{0}(g),\quad g\in\mathsf{G}({\mathbb{Q}}_{p}).$
The pair $(\ell_{\pi_{p}}^{0},\xi_{\pi_{p}}^{0})$ is referred to as the unramified $(T_{\theta},\Lambda_{p})$-Bessel datum for $\pi_{p}$.
Let $\pi\cong\otimes_{p\leqslant\infty}\pi_{p}\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ (see §4.3). Let $N_{\pi}$ be the product of $p\mid N$ such that $\pi_{p}$ is not spherical.
Definition 1.
A system $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ with $\ell_{p}\in(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda_{p}}-(0)$, $\xi_{p}\in V_{\pi_{p}}^{{\bf K}_{0}(N_{\pi}{\mathbb{Z}}_{p})}$ is called a $(T_{\theta},\Lambda)$-Bessel data for $\pi$ if $(\ell_{p},\xi_{p})=(\ell_{\pi_{p}}^{0},\xi_{\pi_{p}}^{0})$ for $(p,N_{\pi})=1$ and $\ell_{p}(\xi_{p})=1$ for all $p<\infty$.
By [32, Theorem 2 8.2 and 9.3], a $(T_{\theta},\Lambda)$-Bessel data exists for our $\pi$. Once a $(T_{\theta},\Lambda)$-Bessel data $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ is fixed, one can define $\varphi_{\pi,\varrho}^{0}\in V_{\pi}^{{\bf K}_{0}(N_{\pi})}[\varrho]$ to be the $V_{\varrho}$-valued cusp form such that for any $v\in V_{\varrho}$, the function $(\varphi_{\pi,\varrho}^{0}(g)\mid v)_{\varrho}$ in $V_{\pi}\cong\otimes_{p\leqslant\infty}V_{\pi_{p}}$ corresponds to the pure tensor $\bar{v}\otimes(\otimes_{p<\infty}\xi_{p})$, where $\bar{v}=1\otimes v\in\overline{V_{\varrho}}\hookrightarrow V_{\pi_{\infty}}$. A particular choice of $\{(\ell_{p},\xi_{p})\}$ will be made in §6.8 so that $\varphi_{\pi,\varrho}^{0}$ corresponds to a newform on the arithmetic quotient $\Gamma_{0}(N)\backslash{\mathfrak{h}}_{2}$ in the sense of [7, §3.2].
Lemma 4.7.
Let $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ be a $(T_{\theta},\Lambda)$-Bessel data for $\pi\in\Pi_{\rm cusp}(\lambda,N)$. Let $\phi=\otimes_{p}\phi_{p}\in\mathcal{S}({\mathbb{A}}_{E}^{2})$ be a decomposable element. Then, for any $\varphi\in V_{\pi}^{{\bf K}_{0}(N)}[\varrho]$ such that, for any $v\in V_{\varrho}$, $\varphi_{v}$ corresponds to the pure tensor $v\otimes(\otimes_{p<\infty}v_{p})\in\bigotimes_{p\leqslant\infty}V_{\pi_{p}}$, we have
(4.19) $B^{T_{\theta},\Lambda}(\varphi;g)=B^{T_{\theta},\Lambda}(\varphi_{\pi,\varrho}^{0};g_{\infty})\prod_{p<\infty}\ell_{p}(\pi_{p}(g_{p})v_{p}),\quad g=(g_{p})_{p}\in\mathsf{G}({\mathbb{A}}).$
For ${\operatorname{Re}}(s)>1$, $v\in V_{\varrho}$ and $b_{\rm{\bf{f}}}=(b_{p})_{p<\infty}\in\mathsf{G}({\mathbb{A}}_{\rm{\bf{f}}})$,
(4.20) $\langle E(\phi,s,\Lambda,\mu),R(b_{\rm{\bf{f}}}b_{\mathbb{R}}^{\theta})\overline{\varphi_{v}}\rangle=\tfrac{\sqrt{|D|}}{2}Z_{v}^{(\infty)}(\phi_{\infty},\varphi^{0}_{\pi};s,\Lambda)\times\prod_{p\leqslant\infty}Z_{p}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p};b_{p}),$
where
$Z_{v}^{(\infty)}(\phi_{\infty},\varphi^{0}_{\pi};s,\Lambda)$ $:=\int_{{\bf K}_{\infty}^{\#}}\int_{{\mathbb{R}}^{\times}}|a|_{\mathbb{R}}^{s-1}\,f_{\phi_{\infty}}^{(s,1,1)}(k_{\infty}^{\#})(v\mid B^{T_{\theta},\Lambda}(\varphi^{0}_{\pi,\varrho};\mathsf{m}(a1_{2},a)\iota_{\theta}(k^{\#})b_{\mathbb{R}}^{\theta}))_{\varrho}\,{\rm{d}}^{\times}a\,{\rm{d}}k^{\#}_{\infty},$
$Z_{p}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p};b_{p})$ $:=\int_{{\mathbb{Q}}_{p}^{\times}}\int_{{\bf K}_{p}^{\#}}f_{\phi_{p}}^{(s,\Lambda_{p},\mu_{p})}(k_{p}^{\#})\mu_{p}(a)|a|_{p}^{s-1}\,\overline{B_{v_{p}}}(\mathsf{m}(a_{p}1_{2},a_{p})\iota_{\theta}(k_{p}^{\#})b_{p})\,{\rm{d}}^{\times}a_{p}\,{\rm{d}}k^{\#}_{p}$
with
$B_{v_{p}}(g_{p}):=\ell_{p}(\pi_{p}(g_{p})v_{p}),\quad g_{p}\in\mathsf{G}({\mathbb{Q}}_{p}).$
Proof.
Fix $g_{\infty}\in\mathsf{G}({\mathbb{R}})^{0}$ and $v\in V_{\varrho}$ and regard $\varphi\mapsto B^{T_{\theta},\Lambda}(\varphi_{v};g_{\infty})$ as a linear functional on $\otimes_{p<\infty}V_{\pi_{p}}$ by the natural inclusion $\bigotimes_{p<\infty}V_{\pi_{p}}\hookrightarrow v\otimes(\bigotimes_{p<\infty}V_{\pi_{p}})\hookrightarrow V_{\pi}$. Then, by the local multiplicity one theorem for Bessel functional on $\mathsf{G}({\mathbb{Q}}_{p})$ recalled above, there exists $C_{v}(g_{\infty})\in{\mathbb{C}}$ such that
(4.21) $(v\mid B^{T_{\theta},\Lambda}(\varphi;g_{\infty}))_{\varrho}=C_{v}(g_{\infty})\,\prod_{p<\infty}\overline{\ell_{p}(v_{p})}$
for $\varphi$ corresponding to $\otimes_{p<\infty}v_{p}$. To determine $C_{v}(g_{\infty})$, set $v_{p}=\xi_{p}$ for all $p<\infty$; then by $\ell_{p}(\xi_{p})=1$, we get $(v\mid B^{T_{\theta},\Lambda}(\overline{\varphi_{\pi,\varrho}^{0}};g_{\infty}))_{\varrho}=C_{v}(g_{\infty})$. Now we apply (3.7) with $T_{\theta},\Lambda$ being replaced with $-T_{\theta},\Lambda^{-1}$ and with $\varphi$ being replaced with $R(b_{\rm{\bf{f}}})\varphi$. ∎
4.5. Gamma factor and global Bessel period
Recall the point $b_{\mathbb{R}}^{\theta}$ in (2.14) and the vector $v_{\varrho}^{\theta}\in V_{\varrho}$ in (4.16). Since $\nu(b_{{\mathbb{R}}}^{\theta})>0$, the point $\mathsf{m}(a1_{2},a)b_{{\mathbb{R}}}^{\theta}$ with $a\in{\mathbb{R}}^{\times}$ belongs to $\mathsf{G}({\mathbb{R}})^{0}$ if and only if $a>0$. Thus, by (3.6) and (4.6), we have $B^{T_{\theta},\Lambda}(\varphi_{\pi}^{0};\mathsf{m}(a1_{2},a)b_{\mathbb{R}}^{\theta})=0$ for $a<0$. By Lemma 4.4 and (4.16),
$B^{T_{\theta},\Lambda}(\varphi_{\pi}^{0};g_{\infty})=w_{D}^{-1}R(\varphi_{\pi}^{0},E,\Lambda)\,B^{T_{\theta}}_{\varrho}(g_{\infty}),\quad g_{\infty}\in\mathsf{G}({\mathbb{R}})^{0}.$
Substituting this, we have that $Z^{(\infty)}_{v}(\phi_{\infty},\varphi^{0}_{\pi};s,\Lambda)$ equals
(4.22) $w_{D}^{-1}\overline{R(\varphi^{0}_{\pi},E,\Lambda)}\int_{{\bf K}^{\#}_{\infty}}\int_{0}^{\infty}|a|_{\mathbb{R}}^{s-1}\,f_{\phi_{\infty}}^{(s,1,1)}(k_{\infty}^{\#})(v\mid B^{T_{\theta}}_{\varrho}(\mathsf{m}(a1_{2},a)\iota_{\theta}(k_{\infty}^{\#})b_{\mathbb{R}}^{\theta}))_{\varrho}\,{\rm{d}}^{\times}a\,{\rm{d}}k^{\#}_{\infty}.$
Now we specify $\phi_{\infty}$. Recall that the highest weight of $\varrho$ is $\lambda=(l_{1},l_{2})$ and $l_{1}-l_{2}\in 2{\mathbb{Z}}_{>0}$. Set $d:=l_{1}-l_{2}$ and define
$\mathsf{f}_{\varrho}(u):=(d+1)\frac{(\varrho(C\bar{u}C^{-1})\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}}{(\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}},\qquad u\in{\bf K}^{\#}_{\infty}.$
with $C:=\frac{1}{\sqrt{2}}\left[\begin{smallmatrix}1&1\\ i&-i\end{smallmatrix}\right]$. If $u=\left[\begin{smallmatrix}a&-\bar{b}\\ b&\bar{a}\end{smallmatrix}\right]\in{\bf SU}(2)$ with $a=a^{\prime}+ia^{\prime\prime},\,b=b^{\prime}+ib^{\prime\prime}$ and $A,B\in{\bf Mat}_{2}({\mathbb{R}})$ are defined as in (2.16), so that $(b_{\mathbb{R}}^{\theta})^{-1}\iota_{\theta}(u)(b_{\mathbb{R}}^{\theta}=k_{\infty}(A+iB)$ (Lemma 2.2), then a computation reveals $C\bar{u}C^{-1}=A-iB$. Thus, the automorphism $u\mapsto C\bar{u}C^{-1}$ of ${\bf SU}(2)$ brings the subgroup ${\mathsf{B}}^{\#}({\mathbb{R}})\cap({\bf K}^{\#}_{\infty})^{0}=\{\left[\begin{smallmatrix}a&0\\ 0&a^{-1}\end{smallmatrix}\right]\mid a\in{\mathbb{C}}^{1}\}$ to the subgroup ${\bf SO}(2)$, by which $\mathsf{v}_{\varrho}^{0}$ is fixed. Hence, $\mathsf{f}_{\varrho}$ is left ${\mathsf{B}}^{\#}({\mathbb{R}})\cap({\bf K}^{\#}_{\infty})^{0}$-invariant. Since $\mathsf{v}_{\varrho}^{0}$ is also fixed by the element $C\delta C^{-1}=\delta$, the function $\mathsf{f}_{\varrho}$ is left ${\mathsf{B}}^{\#}({\mathbb{R}})\cap{\bf K}^{\#}_{\infty}$-invariant. Thus, there exists a unique element $\mathsf{f}^{(s)}_{\varrho}$ of ${\mathcal{V}}^{\#}(s,1,1)$ such that $\mathsf{f}_{\varrho}^{(s)}|{\bf K}_{\infty}^{\#}=\mathsf{f}_{\varrho}$. Define $\phi_{\infty}\in{\mathcal{S}}({\mathbb{C}}^{2})$ by
(4.23) $\phi_{\infty}(x,y):=\sum_{j=0}^{d/2}\tfrac{(x\bar{x})^{j}}{(j!)^{2}}\tfrac{(-y\bar{y})^{d/2-j}}{((d/2-j)!)^{2}}\,\tfrac{2}{\pi}\exp\left(-\tfrac{2\pi}{\sqrt{|D|}}(x\bar{x}+y\bar{y})\right),\quad(x,y)\in{\mathbb{C}}^{2}.$
By (3.1),
(4.24) $f_{\phi_{\infty}}^{(s,1,1)}=|D|^{\frac{s+1}{2}}(-1)^{\frac{l_{1}-l_{2}}{2}}{\Gamma_{{\mathbb{C}}}\left(s+\tfrac{l_{1}-l_{2}}{2}+1\right)}\,\mathsf{f}_{\varrho}^{(s)}.$
For $u\in({\bf K}_{\infty}^{\#})^{0}$ and $A,B$ as above, by By (4.1) and (4.16) and by Lemma 2.2, we have
$(v\mid B^{T_{\theta}}_{\varrho}(\mathsf{m}(a1_{2},a)\iota_{\theta}(u)b_{\mathbb{R}}^{\theta})_{\varrho}=(-1)^{\frac{l_{1}+l_{2}}{2}}a^{\frac{l_{1}+l_{2}}{2}}\,\exp\left(-2{\sqrt{|D|}\pi}\,a\right)\,(v\mid\varrho(C\bar{u}C^{-1})^{-1}\mathsf{v}^{0}_{\varrho})_{\varrho}.$
By using the orthogonal relation of matrix coefficients on ${\bf SU}(2)$, the integral in (4.22) is computed as
$|D|^{\frac{s}{2}}(-1)^{l_{1}}{\Gamma_{{\mathbb{C}}}\left(s+\tfrac{l_{1}-l_{2}}{2}+1\right)}(v\mid v_{\varrho}^{0})_{\varrho}\int_{0}^{\infty}a^{s+\frac{l_{1}+l_{2}}{2}-1}\,\exp\left(-2\sqrt{|D|}\pi\,a\right)\,{\rm{d}}^{\times}a$
$=|D|^{\frac{1}{2}(1-\frac{l_{1}+l_{2}}{2})}(-1)^{l_{1}}(v\mid v_{\varrho}^{0})_{\varrho}{\Gamma_{{\mathbb{C}}}\left(s+\tfrac{l_{1}-l_{2}}{2}+1\right)}\Gamma_{\mathbb{C}}\left(s+\tfrac{l_{1}+l_{2}}{2}-1\right)$
for ${\operatorname{Re}}(s)+\tfrac{l_{1}+l_{2}}{2}>1$. Recall $L(s,\pi_{\infty})=\Gamma_{{\mathbb{C}}}(s+\tfrac{l_{1}-l_{2}}{2}+\tfrac{1}{2})\Gamma_{{\mathbb{C}}}(s+\tfrac{l_{1}+l_{2}}{2}-\tfrac{3}{2})$. Thus,
(4.25) $Z^{(\infty)}_{v}(\phi_{\infty},\varphi^{0}_{\pi};s,\Lambda)$ $=w_{D}^{-1}\overline{R(\varphi^{0}_{\pi},E,\Lambda)}(v\mid v_{\varrho}^{0})_{\varrho}(-1)^{l_{1}}|D|^{\frac{1}{2}(1-\frac{l_{1}+l_{2}}{2})}{L\left(s+\tfrac{1}{2},\pi_{\infty}\right)}.$
The formula in [47, 7.23 Lemma] yields
(4.26) $M(s)\mathsf{f}_{\varrho}^{(s)}=\tfrac{\pi}{s-d/2}\prod_{j=1}^{d/2}\tfrac{s-d/2+j-1}{s+d/2-j}\,\mathsf{f}_{\varrho}^{(-s)}.$
By this combined with (3.3) and (4.24), we easily deduce
(4.27) $f_{\widehat{\phi_{\infty}}}^{(-s,1,1)}=\tfrac{{|D|}}{4}(-1)^{\frac{l_{1}-l_{2}}{2}}f_{\phi_{\infty}}^{(-s,1,1)}.$
4.6. The spinor $L$-function and its functional equation
Let $\pi\cong\otimes_{v}\pi_{v}$ be a cuspidal automorphic representation of $\mathsf{G}({\mathbb{A}})$ with the trivial central character: then, $\bar{\pi}_{p}\cong\pi_{p}^{\vee}\cong\pi_{p}$ for all $p<\infty$ by [44, Proposition 2.3]. The twist $\pi_{\mu}$ of $\pi$ by an idele class character $\mu:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}\longrightarrow{\mathbb{C}}^{1}$ is defined on the space $V_{\pi}$ of $\pi$ as $\pi_{\mu}(g):=\pi(g)\cdot(\mu\circ\nu)(g)$ for $g\in\mathsf{G}({\mathbb{A}})$, so that the central character of $\pi_{\mu}$ is $\mu^{2}$. Let $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ and $\mu$ be as in §4.3, so that $\lambda=(l_{1},l_{2})\in{\mathbb{Z}}^{2}$, $l_{1}\equiv l_{2}\pmod{2}$, $l_{1}\geqslant l_{2}>2$ and the ramification loci of $\pi$ and $\mu$ are disjoint. We define the spinor $L$-function of $\pi$ twisted by $\mu$ as the Euler product
$L(s,\pi,\mu):=\prod_{p<\infty}L(s,(\pi_{\mu})_{p}).\quad{\operatorname{Re}}(s)>5/2$
with $L(s,(\pi_{\mu})_{p})$ the local $L$-factor listed in table (6.15). The completed $L$-function is defined as
$\widehat{L}(s,\pi,\mu):=\Gamma_{\mathbb{C}}\left(s+\tfrac{l_{1}-l_{2}}{2}+\tfrac{1}{2}\right)\Gamma_{\mathbb{C}}\left(s+\tfrac{l_{1}+l_{2}}{2}-\tfrac{3}{2}\right)\,L(s,\pi,\mu).$
Using Lemmas 3.1, 4.7 and Lemma 4.6, combining (4.25), (4.27) and the computations of the local zeta integrals for cases 1, 4, 5, and 6 in table (6.23), we obtain a meromorphic continuation of $L(s,\pi,\mu)$ to ${\mathbb{C}}$ as well as the functional equation
$\widehat{L}(s,\pi,\mu)=\varepsilon(s,\pi,\mu)\,\widehat{L}(1-s,\pi,\bar{\mu})\quad\text{with}\quad\varepsilon(s,\pi,\mu):=(-1)^{l_{2}}\tilde{\mu}(N_{\pi}^{2})\left(\tfrac{G(\tilde{\mu})}{\sqrt{M}}\right)^{4}(M^{4}N_{\pi}^{2})^{1/2-s}$
as expected. Moreover $\widehat{L}(s,\pi,\mu)$ is holomorphic except for possible simple poles at $s=3/2,-1/2$, which does not occur when $\mu$ is non-trivial. By (4.26), $M(s)\mathsf{f}_{\varrho}^{(s)}$ has a simple zero at $s=1$ if $l_{1}>l_{2}$, which in turn implies the holomorphy at $s=1$ of the global intertwinign operator applied to the section $f_{\phi}^{(s)}$ as well as the Eisenstein series $E(\phi,s,\Lambda,\mu)$ for $\phi_{\infty}$ as above. Hence, by (4.20) with an appropriate $\phi$, $\widehat{L}(s,\pi,\mu)$ is holomorphic at $s=3/2$ when $l_{1}>l_{2}$. If $\mu$ is real valued, then $G(\mu)=\sqrt{M}$ and $\varepsilon(1/2,\pi,\mu)=(-1)^{l_{2}}$, so that $L(1/2,\pi,\mu)=0$ unless $l_{2}$ is even.
5. Spectral average of Rankin-Selberg integrals
The space $S_{\varrho}({\bf K}_{0}(N))$ is endowed with the Hermitian inner product associated to the norm $\int_{{\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})}(\varphi(g)\mid\varphi(g))_{\varrho}\,{\rm{d}}g$ $(\varphi\in S_{\varrho}({\bf K}_{0}(N))$. Let ${\mathscr{H}}(\mathsf{G}({\mathbb{Q}}_{p})\mathbin{/\mkern-6.0mu/}{\bf K}_{p})$ be the Hecke algebra for $(\mathsf{G}({\mathbb{Q}}_{p}),{\bf K}_{p})$ for $p<\infty$. For any finite set $S$ of primes $p$ that is prime to $N$, define ${\mathscr{H}}_{S}:=\otimes_{p\in S}{\mathscr{H}}(\mathsf{G}({\mathbb{Q}}_{p})\mathbin{/\mkern-6.0mu/}{\bf K}_{p})$. The ${\mathbb{C}}$-algebra ${\mathscr{H}}_{S}$ acts on the finite dimensiona Hilbert space $S_{\varrho}({\bf K}_{0}(N))$ normally by
$[R(f_{S})\varphi](g)=\int_{\mathsf{G}({\mathbb{Q}}_{S})}\varphi(gx_{S})f_{S}(x_{S})\,{\rm{d}}x_{S},\quad g\in\mathsf{G}({\mathbb{A}}),\quad f_{S}\in{\mathscr{H}}_{S},\,\varphi\in S_{l}({\bf K}_{0}(N)).$
where $\mathsf{G}({\mathbb{Q}}_{S}):=\prod_{p\in S}\mathsf{G}({\mathbb{Q}}_{p})$. Let $\mu$ and $M$ be as in §4.3. We define the Schwarts-Bruhat function $\phi\in\mathcal{S}({\mathbb{A}}_{E}^{2})$ associated $\mu$ as
$\phi=\prod_{p{\leqslant}\infty}\phi_{p},\quad\phi_{p}(x,y)=\begin{cases}\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{E_{p}}}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{E_{p}}}(y)\quad&(p<\infty,p\nmid M)\\ \operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{p^{e}\mathfrak{o}_{E_{p}}}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{1+p^{e}\mathfrak{o}_{E_{p}}}(y)&(p<\infty,p^{e}\parallel M,e{\geqslant}1),\end{cases}$
and $\phi_{\infty}$ being as in (4.23). In this section, we investigate the averages
(5.1) ${\mathbb{I}}^{(s)}(l,N,f_{S})$ $:=\frac{1}{[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]}\sum_{\varphi\in{\mathcal{B}}(\lambda,N)}\langle E(\phi,s,\Lambda,\mu),R(b)R(f_{S})\overline{\varphi_{\mathsf{v}_{\varrho}^{0}}}\rangle\,{B^{T_{\theta},\Lambda}(\varphi_{\mathsf{v}_{\varrho}^{0}};b_{\mathbb{R}}^{\theta})},$
where $\mathsf{v}_{\varrho}^{0}\in V_{\varrho}^{{\bf SO}(2)}$ is the vector from §4 and $b=(b_{p})_{p\leqslant\infty}\in\mathsf{G}({\mathbb{A}})$ is define by
(5.2) $b_{p}=\begin{cases}1_{4}\quad&(p<\infty,p\nmid NM),\\ \eta_{p}&(p<\infty,p\mid N),\\ b_{p}^{M}:=\left[\begin{smallmatrix}p^{e}1_{2}&T_{\theta}^{\dagger}\\ 0&1_{2}\end{smallmatrix}\right]&(p<\infty,p^{e}\parallel M,e{\geqslant}1),\\ b_{\mathbb{R}}^{\theta}&(p=\infty)\end{cases}$
with $\eta_{p}:=\left[\begin{smallmatrix}0&0&0&-1\\ 0&0&1&0\\ 0&p&0&0\\ -p&0&0&0\end{smallmatrix}\right]\in\mathsf{G}({\mathbb{Q}}_{p})$ being the Atkin-Lehner element, ${\bf K}_{\rm{\bf{f}}}:={\bf K}_{0}(1)$, and ${\mathcal{B}}(\lambda,N)=\bigcup_{\pi\in\Pi_{\rm cusp}(\lambda,N)}{\mathcal{B}}_{\pi}(\lambda,N)$ is an orthonormal basis of $S_{\varrho}({\bf K}_{0}(N))$ with ${\mathcal{B}}_{\pi}(\lambda,N)$ being an orthonormal basis of $V_{\pi}^{{\bf K}_{0}(N)}[\varrho]$. The sum is independent of the choice of an orthonormal basis and can be written as ${\mathbb{I}}^{(s)}(\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N),f_{S})$, where for any subset $X\subset\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ we define
(5.3) ${\mathbb{I}}^{(s)}(X,f_{S})=\frac{1}{[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]}\sum_{\pi\in X}\widehat{f_{S}}(\pi_{S})\sum_{\varphi\in{\mathcal{B}}_{\pi}(\lambda,N)}\langle E(\phi,s,\Lambda,\mu),R(b)\overline{\varphi_{\mathsf{v}_{\varrho}^{0}}}\rangle\,{B^{T_{\theta},\Lambda}(\varphi_{\mathsf{v}_{\varrho}^{0}};b_{\mathbb{R}}^{\theta})},$
where $\widehat{f_{S}}(\pi_{S})$ is the spherical Fourier transforms of $f_{S}$ on $\pi_{S}:=\otimes_{p\in S}\pi_{p}$, which is defined as the eigenvalue of the operator $\pi_{S}(f_{S})$ on the ${\bf K}_{S}=\prod_{p\in S}{\bf K}_{p}$-fixed vectors $\pi_{S}^{{\bf K}_{S}}\cong{\mathbb{C}}$. For $\bullet\in\{{\rm T},{\rm G},{\rm Y},{\rm SK}\}$ and $\ast\in\{{\rm new},{\rm old}\}$, we define ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\bullet}$, ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\ast}$, and ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\bullet,\ast}$ to be ${\mathbb{I}}^{(s)}(X,f_{S})$ with $X=\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\bullet}$, $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\ast}$ and $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\bullet,\ast}$, respectively. Then, due to (4.18), the average (5.1) is rearranged as
(5.4) ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})$ $={\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm G},{\rm new}}+{\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm Y},{\rm new}}$
$+{\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm SK},{\rm new}}+{\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm old}}.$
5.1. A construction of orthonormal basis
Let $\pi\cong\otimes_{p\leqslant\infty}\pi_{p}$ be an element of the set $\Pi_{\rm csup}^{(T_{\theta},\Lambda)}(\lambda,N)$ (see §4.3). We fix a $(T_{\theta},\Lambda)$-Bessel data $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ of $\pi$ (see §4.4) once and for all. Set $\varphi_{\pi}^{0}(g):=(\varphi_{\pi,\varrho}^{0}(g)\mid\mathsf{v}_{\varrho^{0}})_{\varrho}$, which is an element of $V_{\pi}^{{\bf K}_{0}(N)}$. Since $\pi$ is a unitary representation, all of its factors $\pi_{p}$ are unitarizable. For each $p<\infty$, we can uniquely fix a $\mathsf{G}({\mathbb{Q}}_{p})$-invariant inner product $(\cdot|\cdot)_{p}$ on $V_{\pi_{p}}$ by demanding $(\xi_{p}|\xi_{p})_{p}=1$; for $p=\infty$, we fix a $\mathsf{G}({\mathbb{R}})$-invarinat inner product so that its pull back to $V_{\varrho}\hookrightarrow\pi_{\infty}$ (cf. §4.2(i)) coincides with the inner product $(\cdot|\cdot)_{\varrho}$ of $V_{\varrho}$. Let $\varphi_{\pi}^{0}$ be the global new form attached to $\{\xi_{p}\}_{p<\infty}$ (see §4.4). Then,
(5.5) $\frac{(\varphi|\varphi)_{L^{2}}}{(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}}=(v|v)_{\varrho}\prod_{p<\infty}(v_{p}|v_{p})_{p}$
for any $\varphi\in V_{\pi}$ that corresponds to $v\otimes(\otimes_{p<\infty}v_{p})$ with $v\in V_{\varrho}$. As below, for $p<\infty$, we fix an orthonormal basis ${\mathcal{B}}(\pi_{p},{\bf K}_{0}(N{\mathbb{Z}}_{p}))$ of $V_{\pi_{p}}^{{\bf K}_{0}(N{\mathbb{Z}}_{p})}$ in such a way that
${\mathcal{B}}(\pi_{p},{\bf K}_{0}(N{\mathbb{Z}}_{p}))=\{\xi_{\pi_{p}}^{0}\}\quad\text{if $p\nmid N$}.$
Then, a pure tensor of the form
(5.6) $(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}^{-1/2}\times\otimes_{p<\infty}v_{p},\quad v_{p}\in{\mathcal{B}}(\pi_{p},{\bf K}_{0}(N{\mathbb{Z}}_{p})),$
yield an element $\varphi\in V_{\pi}^{{\bf K}_{0}(N)}[l]$ such that for $v\in V_{\varrho}$ the element $\varphi_{v}\in V_{\pi}$ corresponds to $v$ tensored with (5.6). The set of functions $\varphi$ obtained in this way from (5.6) will be denoted by ${\mathcal{B}}_{\pi}(\lambda,N)$. If $\pi\in\Pi_{\rm cusp}(\lambda,N)$ does not satisfy condition (A-iv) in §4.3, then we fix arbitrary orthonormal basis ${\mathcal{B}}_{\pi}(\lambda,N)$ of the space $V_{\pi}^{{\bf K}_{0}(N)}[\lambda]$. Let $\mathcal{B}(\lambda,N)$ be the union of the sets ${\mathcal{B}}_{\pi}(\lambda,N)$ for $\pi\in\Pi_{\rm cusp}(\lambda,N)$; then, by (ii) in §4.2, the set ${\mathcal{B}}(\lambda,N)$ is an orthonormal basis of $S_{\varrho}({\bf K}_{0}(N))$.
5.2. Computation of the average
Let $\varphi$ correspond to a pure tensor as in (5.6). Then, by (4.20) and by the computations recalled in §4.5 and §6.2,
(5.7) $\langle E(\phi,s,\Lambda,\mu),R(b)\overline{\varphi_{\mathsf{v}_{\varrho}^{0}}}\rangle$ $=(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}^{-1/2}\,w_{D}^{-1}\,\overline{R(\varphi^{0}_{\pi},E,\Lambda)}\,\widehat{L}\left(s+\tfrac{1}{2},\pi,\mu\right)$
×$\times(\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}(-1)^{l_{2}}2^{-2}|D|^{\frac{1}{2}(4-\frac{l_{1}+l_{2}}{2})}\,\prod_{p\mid N}Z_{p}^{*}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p};b_{p}),$
where
$Z_{p}^{*}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p};b_{p}):={L\left(s+\tfrac{1}{2},\pi_{p},\mu_{p}\right)}^{-1}\,Z_{p}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p};b_{p})$
is the normalized local zeta-integral. Moreover, by (4.21) and Lemma 4.4,
(5.8) $B^{T_{\theta},\Lambda}(\varphi_{\mathsf{v}_{\varrho}^{0}};b_{\mathbb{R}}^{\theta})$ $=\langle\varphi_{\pi}^{0}\mid\varphi_{\pi}^{0}\rangle_{L^{2}}^{-1/2}\,w_{D}^{-1}R(\varphi_{\pi}^{0},E,\Lambda)\,(B^{T_{\theta}}_{\varrho}(b_{\mathbb{R}}^{\theta})\mid\mathsf{v}_{\varrho}^{0})_{\varrho}\,\prod_{p\mid N}\ell_{p}(v_{p}).$
Note that $B_{\varrho}^{T_{\theta}}(b_{\mathbb{R}}^{\theta})=(-1)^{\frac{l_{1}+l_{2}}{2}}\exp(-2\pi\sqrt{|D|})\,\mathsf{v}_{\varrho}^{0}$ by (4.1) and (2.14). From (5.3), (5.7) and (5.8), we get
Proposition 5.1.
Let $f_{S}\in{\mathscr{H}}_{S}$. Then, ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})$ equals to
$\frac{2^{-2}|D|^{\frac{1}{2}(3-\frac{l_{1}+l_{2}}{2})}e^{-2\pi\sqrt{|D|}}}{w_{D}^{2}[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]}\times M^{s-6}\zeta_{M}(1)\zeta_{M}(4)\tilde{\mu}(2D)G(\tilde{\mu})$
(5.9) ×$\quad\times(\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}^{2}\sum_{\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)}\widehat{f_{S}}(\pi_{S})\frac{|R(\varphi^{0}_{\pi},E,\Lambda)|^{2}}{(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}}\widehat{L}\left(s+\tfrac{1}{2},\pi,\mu\right)$
×$\quad\times\prod_{p<\infty}{\mathbb{I}}_{\pi_{p},{\bf K}_{0}(N{\mathbb{Z}}_{p})}^{(s)}(\xi_{p},\phi_{p},\Lambda_{p},\mu_{p};b_{p}),$
where $\xi_{p}$ for $p<\infty$ is from the fixed $(T_{\theta},\Lambda)$-Bessel data $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ of $\pi$, and
(5.10) ${\mathbb{I}}^{(s)}_{\pi_{p},{\bf K}_{0}(N{\mathbb{Z}}_{p})}(\xi_{p},\phi_{p},\Lambda_{p},\mu_{p};b_{p}):=\sum_{v\in{\mathcal{B}}(\pi_{p},{\bf K}_{0}(N{\mathbb{Z}}_{p})}Z^{*}_{p}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p},b_{p})\,{\ell_{p}(v)}.$
We also use the notation $\zeta_{M}(s)$ to denote $\prod_{p\mid M}(1-p^{-s})^{-1}$.
Note that if $\pi_{p}$ with $p\nmid N$ is unramified and $(\ell_{\pi_{p}}^{0},\xi_{\pi_{p}}^{0})$ is the unramified Bessel datum of $\pi_{p}$, then (5.10) is $1$. For other cases, we compute (5.10) in §6.8 completely. Substituting them, we obtain
Theorem 5.2.
Let $(\lambda=(l_{1},l_{2}),N)$, $\mu$, $\tilde{\mu}$ and $M$ be as in §4.3. Let $S$ be a finite set of prime numbers relatively prime to $DMN$. Then, for any $f_{S}=\otimes_{p\in S}f_{p}\in{\mathscr{H}}_{S}$,
(5.11) ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})$ $=\frac{2^{-2}|D|^{\frac{1}{2}(3-\frac{l_{1}+l_{2}}{2})}e^{-2\pi\sqrt{|D|}}}{w_{D}^{2}[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]}M^{s-6}\zeta_{M}(1)\zeta_{M}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\mid N}(1+p^{-2})^{-1}$
×$\times(\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}^{2}\sum_{\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)}\widehat{f_{S}}(\pi_{S})\frac{|R(\varphi^{0}_{\pi},E,\Lambda)|^{2}}{(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}}\widehat{L}\left(s+\tfrac{1}{2},\pi,\mu\right)\,{\mathfrak{t}}(\pi,\mu)$
with ${\mathfrak{t}}(\pi,\mu)=\prod_{p\mid N}{\mathfrak{t}}(\pi_{p},\mu_{p})$ with
(5.12) ${\mathfrak{t}}(\pi_{p},\mu_{p}):=\begin{cases}1\quad(\text{$p\mid N_{\pi}$ and $\pi_{p}$ is of type VIb}),\\ 2\quad(\text{$p\mid N_{\pi}$ and $\pi_{p}$ is of type IIIa}),\\ 2(p-1)p^{-5}L(1,\pi_{p},{\rm Std})\,\{1+\mu_{p}^{2}(p)-\tfrac{\mu_{p}(p)}{p+1}{\operatorname{tr}}(p^{-1}T_{p}+\eta_{p}|\pi_{p}^{{\bf K}_{0}(p{\mathbb{Z}}_{p})})\}\quad(\text{$p\mid\frac{N}{N_{\pi}}$}).\end{cases}$
To refine (5.11) by considering ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm T},{\rm new}}$ and ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm SK},{\rm new}}$, we refer to the explicit formula of the quantity $\frac{|R(\varphi^{0}_{\pi},E,\Lambda)|^{2}}{(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}}$ first proved by Dickson, Pitale, Saha, and Schmidt ([7, Theorems 1.13 and 3.13]) when $l_{1}=l_{2}$ under the refined GGP conjecture for Bessel periods posed by Liu([23]). Owing to a theorem by Furusawa and Morimoto [11, Theorem 8.1], the formula in [7, Theorems 1.13 and 3.13] is extended to vector-valued forms unconditionally so that it can applied to (5.11) to yield the following theorem. Note that the polynomial $Q_{S,\varrho}$ (with $S=T_{\theta}$) in [10, (8.2.16)] equals $(-1)^{\frac{l_{1}-l_{2}}{2}}(\frac{2}{\sqrt{|D|}})^{\frac{l_{1}+l_{2}}{2}}\mathsf{v}_{\varrho}^{\theta}$, our $R(\varphi,E,\Lambda)$ equals the quantity $w_{D}(\frac{\sqrt{|D|}}{2})^{-\frac{l_{1}+l_{2}}{2}}\,\mathcal{B}_{\Lambda}(\varphi;E)/(Q_{S,\varrho},Q_{S,\varrho})_{l_{1}-l_{2}}$ defined by [10, (8.2.17)], and $(Q_{S,\varrho},Q_{S,\varrho})_{l_{1}-l_{2}}=(\frac{|D|}{4})^{-\frac{l_{1}+l_{2}}{2}}(\mathsf{v}_{\varrho}^{0},\mathsf{v}_{\varrho}^{0})_{l_{1}-l_{2}}=(\frac{|D|}{4})^{-\frac{l_{1}+l_{2}}{2}}(\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}$.
Theorem 5.3.
Let the notations and the assumptions be as in Theorem 5.2. We have
  • (i)
    ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm G},{\rm new}}$ equals
    $\frac{2^{\#S(N)+2l_{1}-6}|D|^{\frac{l_{1}+l_{2}+2}{4}}e^{-2\pi\sqrt{|D|}}}{[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]}$
    ×$\times M^{s-6}\zeta_{M}(1)\zeta_{M}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\in S(N)}(1+p^{-1})$
    ×$\times\sum_{\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm G},{\rm new}}\widehat{f_{S}}(\pi_{S})\frac{\widehat{L}(\frac{1}{2},\pi\times{\mathcal{A}I}(\Lambda^{-1}))}{\widehat{L}(1,\pi;{\rm Ad})}\widehat{L}\left(s+\tfrac{1}{2},\pi,\mu\right),$
    where $S(N)$ is the set of all the prime divisors of $N$, and ${\mathcal{A}I}(\Lambda^{-1})$ is the automorphic representation of ${\bf GL}_{2}({\mathbb{A}})$ defined by the automorphic induction of $\Lambda^{-1}$ from ${\mathbb{A}}_{E}^{\times}$.
  • (ii)
    If $N$ has even number of prime divisors, then ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm Y},{\rm new}}=0$. If $N$ has odd number of prime divisors, then ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm Y},{\rm new}}$ equals
    $\frac{2^{\#S(N)+2l_{1}-7}|D|^{\frac{l_{1}+l_{2}+2}{4}}e^{-2\pi\sqrt{|D|}}}{[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]}$
    ×$\times M^{s-6}\zeta_{M}(1)\zeta_{M}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\in S(N)}(1+p^{-1})$
    ×$\times\sum_{\begin{subarray}{c}\pi_{1}\in\Pi_{{\bf PGL}_{2},{\rm cusp}}(l_{1}+l_{2}-2,N)^{\rm new}\\ \pi_{2}\in\Pi_{{\bf PGL}_{2},{\rm cusp}}(l_{1}-l_{2}+2,N)^{\rm new}\end{subarray}}\widehat{f_{S}}({\rm Y}(\pi_{1},\pi_{2})_{S})\frac{\widehat{L}(\frac{1}{2},\pi_{1}\times{\mathcal{A}I}(\Lambda^{-1}))\widehat{L}(\frac{1}{2},\pi_{2}\times{\mathcal{A}I}(\Lambda^{-1}))}{\widehat{L}(1,\pi_{1};{\rm Ad})\widehat{L}(1,\pi_{2};{\rm Ad})\widehat{L}(1,\pi_{1}\times\pi_{2})}$
    ×$\hskip 227.62204pt\times\widehat{L}\left(s+\tfrac{1}{2},\pi_{1}\times\mu\right)\widehat{L}\left(s+\tfrac{1}{2},\pi_{2}\times\mu\right),$
    where $\Pi_{{\bf PGL}_{2},{\rm cusp}}(k,N)^{\rm new}$ is the set of all irreducible cuspidal representations of ${\bf PGL}_{2}({\mathbb{A}})$ associated to holomorphic newforms of weight $k$ and level $N$ and ${\rm Y}(\pi_{1},\pi_{2})\in\Pi_{\rm cusp}(\lambda,N)^{\rm new}$ denotes the Yoshida lift of $\pi_{1}$ and $\pi_{2}$.
  • (iii)
[…]
    $\frac{3\cdot 2^{2l-3}|D|^{(l+2)/2}e^{-2\pi\sqrt{|D|}}}{[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]}\cdot\frac{s}{4\pi}$
    ×$\times M^{s-6}\zeta_{M}(1)\zeta_{M}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\in S(N)}(1+p^{-1})\cdot\widehat{L}(s+1,\mu)\widehat{L}(s,\mu)$
    ×$\times\sum_{\pi_{0}\in\Pi_{{\bf PGL}_{2},{\rm cusp}}^{(T_{\theta},1)}(2l-2,N)^{\rm new}}\widehat{f_{S}}({\rm SK}(\pi_{0})_{S})\frac{\widehat{L}(\frac{1}{2},\pi_{0}\times\chi_{D})\widehat{L}(1,\chi_{D})^{2}}{\widehat{L}(\frac{3}{2},\pi_{0})\widehat{L}(1,\pi_{0};{\rm Ad})}\widehat{L}\left(s+\tfrac{1}{2},\pi_{0}\times\mu\right),$
[…]
Proof.
The equation (i) and (ii) is a direct corollary to Theorem 5.3 and [10, Theorem 8.1](for the scalar case we refer to [7, Theorems 1.13 and 3.14]; we also note the formula
$\widehat{L}(s,\pi,\mu)$ $=\widehat{L}(s,\pi_{1}\times\mu)\widehat{L}(s,\pi_{2}\times\mu),$
where $\pi=Y(\pi_{1},\pi_{2})$. This obtained immediately by comparing the local factors. It is noted in [7, p.296] that each local representation $\pi_{p}$ of $\pi\cong\otimes_{p}\pi_{p}\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm Y,new}$ for $p\mid N$ is of type VIb. Finally, we check the statement (iii). As quoted in [7, Theorem 3.11], the $(T_{\theta},\Lambda)$-Bessel periods of the Saito-Kurokawa lifts are zero unless $\Lambda$ is trivial due to Qiu. By comparing the local $L$-functions ([39, Theorem 5.2(ii)], cf. [29, Theorem 3.1]),
(5.13) $\widehat{L}(s,\pi,\mu)=\frac{1}{4\pi}\left(s-1/2\right)\widehat{L}(s,\pi_{0}\times\mu)\widehat{L}(s+1/2,\mu)\widehat{L}(s-1/2,\mu),$
where $\pi={\rm SK}(\pi_{0})$. By [39] and the definition, the set $\Pi_{{\rm cusp}}^{(T_{\theta},1)}((l,l),N)^{{\rm SK}}$ corresponds bijectively to the set $\Pi_{{\bf PGL}_{2},{\rm cusp}}^{(T_{\theta},1)}(2l-2,N)$ via the Saito-Kurokawa lifting if $l{\geqslant}3$. Note that the condition $l{\geqslant}3$ is ensured by assumption (A-ii) in §4.3. ∎
By (5.4), we have that ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})-{\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm G,new}$ is the sum of ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm Y,new}$, ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm SK,new}$ and ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm old}$. Our main focus is their values at $s=0$. For those, we have the following upper bounds in the level aspect.
Theorem 5.4.
Let $l$, $\mu$, $\widetilde{\mu}$, $M$, S be as in Theorem 5.2.
  • (i)
    Let $\lambda$ be fixed as before and $N$ an inert prime in $E/{\mathbb{Q}}$. Then, there exist constants $C_{D,\lambda,\mu,f_{S}},C_{D,\lambda,\mu,f_{S}}^{\prime}>0$ depending only on $D$, $l$, $\mu$, and $f_{S}$ such that
    (5.14) $|{\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{{\rm old}}|<C_{D,\lambda,\mu,f_{S}}N^{-8}.$
    (5.15) $|{\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{{\rm Y,new}}|<C_{D,\lambda,\mu,f_{S}}^{\prime}N^{-\frac{3}{2}}.$
  • (ii)
    Let $N$ be $1$ or an inert prime in $E/{\mathbb{Q}}$. When $\mu\neq{\bf 1}$, ${\mathbb{I}}^{(0)}(l,N,f_{S})^{{\rm SK,new}}=0.$ When $\mu={\bf 1}$, there exists a constant $C_{f_{S}}>0$ depending only on $f_{S}$ such that
    (5.16) $|{\mathbb{I}}^{(0)}((l,l),N,f_{S})^{{\rm SK,new}}|<C_{f_{S}}|D|^{\frac{l}{2}}l^{\frac{1}{2}}N^{-\frac{3}{2}}.$
Proof.
Because $N$ is a prime, one can check that the summation range of ${\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{{\rm old}}$ is $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{{\rm old}}=\Pi_{\rm cusp}^{(T_{\theta},\chi)}(\lambda,1)$, which is independent of $N$. From the formula of ${\mathfrak{t}}(\pi_{p},\mu_{p})$ in (5.12) for $p=N$, combined with the temperedness of $\pi_{N}$ due to [48] and the matrix of $T_{1,0}$ in table (6.49), we get ${\mathfrak{t}}(\pi_{N},\mu_{N})=O(N^{-4})$. This and the equation $[{\bf K}_{\bf f}:{\bf K}_{0}(N)]=N^{3}(1+N^{-2})(1+N^{-1})$ for prime $N$ yield the bound (5.14).
Next we treat the average for Yoshida lifts. For $\pi_{1}\in\Pi_{{\bf PGL}_{2},{\rm cusp}}(l_{1}+l_{2}-2,N)^{\rm new}$ and $\pi_{2}\in\Pi_{{\bf PGL}_{2},{\rm cusp}}(l_{1}-l_{2}+2,N)^{\rm new}$, we need the lower bound
$L(1,\pi_{1}\times\pi_{2})\gg\exp(-C\sqrt{\log N}),$
uniform in $N$. This is a special case of Lemma 5.5, because $\pi_{1}$ and $\pi_{2}$ are everywhere tempered by Deligne’s estimate and by the fact that the local $p$-components of $\pi_{i}$ for $p\mid N$ are the (twisted) Steinberg representation, which is tempered. Note that $\pi_{1}\not\cong\pi_{2}$ due to the weight condition. Now, We deduce the inequality (5.15) bounding the sum from above by a product of the average considered in [8, Theorem 1.1]; to do this, we invoke the subconvexity bound
(5.17) $L(\tfrac{1}{2},\pi)=O(C(\pi)^{\frac{1}{4}-\delta}),\quad(\exists\,\delta>0)$
for automorphic representations $\pi$ of ${\bf GL}_{2}({\mathbb{A}})$ ([24]) and the non-negativity of $L(\frac{1}{2},\pi_{i}\times\mathcal{AI}(\Lambda^{-1})=L(\frac{1}{2},{\rm BC}_{E/{\mathbb{Q}}}(\pi_{i})\otimes\Lambda^{-1})$ for $i=1,2$ ([16])).
Let us prove (5.16); by Theorem 5.3 (ii), we may assume $\lambda=(l,l)$. Suppose $\mu\not={\bf 1}$; then, by Theorem 5.3 (ii), we have that ${\mathbb{I}}^{(s)}((l,l),N,f_{S})^{{\rm SK,new}}\times s^{-1}$ is entire, hence $\mathbb{I}^{(0)}((l,l),N,f_{S})^{\rm SK,new}=0$ because $\widehat{L}(s,\mu)$, as well as $\widehat{L}(s,\pi_{0}\times\chi_{D})$, is entire. In the rest of the proof, we assume $\mu={\bf 1}$. Then, $s\widehat{L}(s,\mu)\widehat{L}(s+1,\mu)$ has a simple pole at $s=0$. By [38, Theorem 3.1 and Table 2], ${\rm SK}(\pi_{0})$ with $\pi_{0}\in\Pi_{{\rm cusp},{\bf PGL_{2}}}(2l-2,N)$ has the global $(T_{\theta},\Lambda)$-Bessel model only if the sign of the functional equation of $\pi_{0}$ is $-1$ so that $\widehat{L}(s+1/2,\pi_{0})$ has a zero at $s=0$. Hence, theorem 5.3 (ii) gives us the following majorant of ${\mathbb{I}}^{(0)}((l,l),N,f_{S})^{{\rm SK,new}}$.
$2^{3}\pi^{\frac{5}{2}}(l-1)^{-2}\frac{\Gamma(l)}{\Gamma(l-\frac{1}{2})}|D|^{\frac{l}{2}}[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]^{-1}\sum_{\pi_{0}\in\Pi_{{\bf PGL}_{2},{\rm cusp}}^{(T_{\theta},{\bf 1})}(2l-2,N)}\frac{L(\frac{1}{2},\pi_{0}\times\chi_{D})\,|L^{\prime}(\frac{1}{2},\pi_{0})|}{L(1,\pi_{0};{\rm Ad})}.$
To estimate this, we invoke a subconvexity bound $L({1}/{2},\pi_{0}\times\chi_{D})=O((l^{2}N)^{\frac{1}{4}-\delta})$ for some $\delta>0$ from (5.17) and a lower bound $L(1,\pi_{0};{\rm Ad})>c_{0}\exp(-c_{1}\sqrt{\log(1+l^{2}N)})$ for some constant $c_{0},c_{1}>0$ which is known by [13, Theorem 0.1] (see also the remark after [22, Corollary 7]). The lower bound implies $L(1,\pi_{0};{\rm Ad})^{-1}=O((l^{2}N)^{2\delta})$. Moreover, by a common argument, a subconvexity bound for the central derivative of $L^{\prime}(1/2,\pi_{0})=O((l^{2}N)^{\frac{1}{4}-\delta})$ is derived from a subconvexity bound for $L(1/2,\pi_{0}|\cdot|_{\mathbb{A}}^{it})$ available from (5.17). By the uniform bound #$\#\Pi_{{\bf PGL}_{2},\rm cusp}(2l-2,N)=O(lN)$ and the asymptotic $\frac{\Gamma(l)}{\Gamma(l-\frac{1}{2})}\sim l^{\frac{1}{2}}\ (l\to\infty)$, the bound (5.16) follows. ∎
Lemma 5.5.
Let $\pi_{i}$ ($i=1,2)$ be irreducible cuspidal automorphic representations of ${\bf PGL}_{n}$ such that $C(\pi_{1}),C(\pi_{2})\leqslant Q$ with $Q>2$. We assume that $\pi_{1}\not\cong\pi_{2}$ and both of them are self-dual and tempered everywhere. Then,
$L(1,\pi_{1}\times\pi_{2})\gg\exp(-C\sqrt{\log Q}).$
with an absolute constant $C>0$.
Proof.
We recall the argument indicated in [36] (attributed originally to [25]), which eliminates a possibility of Siegel zeros of the $L$-function $L(s,\pi_{1}\times\pi_{2})$. Fix $t\in{\mathbb{R}}$. Then, $L(s,\Pi\times\widetilde{\Pi})$ with $\Pi$ being the isobaric sum
$\Pi:=\pi_{1}\boxplus(\pi_{1}\times|\cdot|_{\mathbb{A}}^{it})\boxplus(\pi_{1}\times|\cdot|_{\mathbb{A}}^{-it})\boxplus\pi_{2}\boxplus(\pi_{2}\times|\cdot|_{\mathbb{A}}^{it})\boxplus(\pi_{2}\times|\cdot|_{\mathbb{A}}^{-it})$
is a Dirichlet series with non-negative coefficients ([14, Lemma a]). Moreover, by ([31, Proposition 9.4]) and due to $\pi_{1}$ and $\pi_{2}$ being self-dual, $L(s,\Pi\times\widetilde{\Pi})$ is expressed as
$L(s,\Pi\times\widetilde{\Pi})=\prod_{1{\leqslant}i,j{\leqslant}2}\left[{\begin{subarray}{c}L(s,\pi_{i}\times\pi_{j})^{3}L(s-it,\pi_{i}\times\pi_{j})^{2}L(s+it,\pi_{i}\times\pi_{j})^{2}\\ \times L(s-2it,\pi_{i}\times\pi_{j})L(s+2it,\pi_{i}\times\pi_{j})\end{subarray}}\right]^{2},$
which shows that $L(s,\Pi\times\widetilde{\Pi})$ has a pole of order $6$ at $s=1$ (due to $\pi_{1}\ncong\pi_{2}$), and has a zero of order $8$ at $s=\sigma$ if $L(\sigma+it,\pi_{i}\times\pi_{j})=0$ $(i,j=1,2)$. Hence, by [12, Lemma, p.178], one can show that $L(s,\pi_{1}\times\pi_{2})$ has no zeros in the interval $(1-\frac{C_{0}}{\log M},1)$ for some constant $C_{0}>0$ with $M=(1+|t|)^{24}C(\pi_{2}\times\pi_{2})^{18}C(\pi_{1}\times\pi_{1})^{9}C(\pi_{2}\times\pi_{2})^{2}$, where $C(\pi_{i}\times\pi_{j})$ is the analytic conductor of $L(s,\pi_{i}\times\pi_{j})$. By [14, lemma b], we have $C(\pi_{i}\times\pi_{j})\ll Q^{2n}$. Thus, for an absolute constant $C_{0}^{\prime}>0$, $L(s,\pi_{1}\times\pi_{2})$ has no zeros in the interval $(1-\frac{C_{0}^{\prime}}{\log Q},1)$. Now, we apply the argument of [22, Corollary 7]) to the $L$-function $L(s,\pi_{1}\times\pi_{2})$. Since $\pi_{1},\pi_{2}$ are assumed to be tempered everywhere, the optimal bound of the lambda function $\Lambda_{\mathcal{A}}(n)$ for $\mathcal{=}\pi_{1}\times\pi_{2}$ is available so that the automorphy of $\pi_{1}\times\pi_{2}$ is not necessary, which simplifies the proof to get the lower bound of $L(1,\pi_{1}\times\pi_{2})$. ∎
As a consequence of Theorem 5.4, we get
(5.18) ${\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{\rm{G,new}}={\mathbb{I}}^{(0)}(\lambda,N,f_{S})+O_{\Lambda,\mu,\lambda,f_{S}}(N^{-\frac{3}{2}}).$
6. Computation of local zeta-integrals for $p$-adic fields
In this section, let $F$ be a non-archimedean local field of characteristic 0, $\mathfrak{o}$ the integer ring of $F$, ${\mathfrak{p}}$ the maximal ideal of $\mathfrak{o}$, $\varpi$ a generator of ${\mathfrak{p}}$ and $q=\#(\mathfrak{o}/{\mathfrak{p}})$. Let ||$\cdot$|| ⋅ | denote the normalized absolute value of $F$, i.e., $|\varpi|=q^{-1}$. Fix a non-trivial additive character $\psi:F\longrightarrow{\mathbb{C}}^{1}$ with ${\rm cond}(\psi):=\min\{n\in{\mathbb{Z}}\ ;\ \psi|_{\varpi^{n}\mathfrak{o}}=1\}=0$.
We compute the local zeta-integral a la Piatetski-Shapiro ([30]) for several representations, taking particular test functions; as a result, we determine the local $L$-factors and the local $\varepsilon$-factors in [30] to confirm that they coincides with the expected ones listed in [35, Tables A8,A9]. As explained in §4.3, for a particular global application in mind, we only deal with representations of types I, IIb, IIIa and VIb (but allowing the central characters to be non-trivial.)
6.1. Local zeta integral for Bessel models
We first review some generalities on local zeta integrals and then recall results from [32] on explicit formulas of Bessel functions for Iwahori spherical representations of $\mathsf{G}$, which are possible local components of $\pi\cong\otimes_{p}\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\varrho,N)$.
Let $K=\mathsf{G}(\mathfrak{o})$ be a standard maximal compact subgroup of $\mathsf{G}(F)$ and $K_{0}({\mathfrak{p}}^{e}):=\{\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in K\ |\ C\equiv 0\ ({\rm mod}\ {\mathfrak{p}}^{e})\}$ be the congruence subgroup of level ${\mathfrak{p}}^{e}$.
For a symmetric matrix $S=[\begin{smallmatrix}{\mathsf{a}}&\mathsf{b}/2\\ \mathsf{b}/2&\mathsf{c}\end{smallmatrix}]\in{\rm Sym}_{2}(F)$ such that ${\mathsf{d}}:=\mathsf{b}^{2}-4{\mathsf{a}}\mathsf{c}\neq 0$, set $\xi=[\begin{smallmatrix}\mathsf{b}/2&\mathsf{c}\\ -{\mathsf{a}}&-\mathsf{b}/2\end{smallmatrix}]$. Let $L:=F\oplus F\xi$ be the two-dimensional $F$-algebra, $\mathfrak{o}_{L}$ be its integer ring, and $\psi_{L}$ be the additive character on $L$ given by $\psi_{L}:=\psi\circ{\operatorname{tr}}_{L/F}$. The maps
$L\owns x+y\xi\longmapsto x+\tfrac{\sqrt{{\mathsf{d}}}}{2}y\in F(\sqrt{{\mathsf{d}}})\quad({\mathsf{d}}\notin(F^{\times})^{2}),$
$L\owns x+y\xi\longmapsto\left(x+\tfrac{\sqrt{{\mathsf{d}}}}{2}y,x-\tfrac{\sqrt{{\mathsf{d}}}}{2}y\right)\in F\oplus F\quad({\mathsf{d}}\in(F^{\times})^{2})$
are isomorphisms of $F$-algebras. We assume that
  • (i)
    ${\mathsf{a}}\in\mathfrak{o}^{\times}$ and $\mathsf{b},\mathsf{c}\in\mathfrak{o}$.
  • (ii)
    If ${\mathsf{d}}\notin(F^{\times})^{2}$, then $\sqrt{{\mathsf{d}}}\mathfrak{o}_{L}$ is the different ideal of $L/F$.
  • (iii)
    If ${\mathsf{d}}\in(F^{\times})^{2}$, then ${\mathsf{d}}\in\mathfrak{o}^{\times}$.
We can check that
$L^{\times}=\left\{\tau\in{\bf GL}_{2}(F)\ \middle|\ {}^{t}\tau S\tau=\det\tau\cdot S\right\}.$
[…]
$\langle x,y\rangle:={\operatorname{tr}}_{L/F}(-(2\xi)^{-1}(x_{1}y_{2}-x_{2}y_{1})),\quad x=\left[\begin{smallmatrix}x_{1}\\ x_{2}\end{smallmatrix}\right],y=\left[\begin{smallmatrix}y_{1}\\ y_{2}\end{smallmatrix}\right]\in L^{2}.$
We set
$G^{\#}(F)$ :=$:=\left\{g\in{\bf GL}_{2}(L)\ \middle|\ \det g\in F^{\times}\right\},$
$B^{\#}(F)$ :=$:=\left\{\left[\begin{smallmatrix}\tau&\beta\\ 0&a\tau^{\dagger}\end{smallmatrix}\right]\ \middle|\ \tau\in L^{\times},\beta\in L,a\in F^{\times}\right\},$
$K^{\#}$ $:=G^{\#}(F)\cap{\bf GL}_{2}(\mathfrak{o}_{L}).$
$K_{0}^{\#}({\mathfrak{p}}^{e})$ $:=G^{\#}(F)\cap\left[\begin{smallmatrix}\mathfrak{o}_{L}&\mathfrak{o}_{L}\\ {\mathfrak{p}}^{e}\mathfrak{o}_{L}&\mathfrak{o}_{L}\end{smallmatrix}\right].$
Since $\langle g^{\#}x,g^{\#}y\rangle=\det g^{\#}\langle x,y\rangle$ for any $x,y\in L^{2}$ and $g^{\#}\in\mathsf{G}(F)$, we get a natural embedding $G^{\#}(F)\owns g\mapsto\iota(g^{\#})\in\mathsf{G}(F)$. More precisely, $\iota(g)$ is the representation matrix of with respect to an $F$-basis $\{[\begin{smallmatrix}{\mathsf{a}}\\ 0\end{smallmatrix}],[\begin{smallmatrix}\theta_{0}\\ 0\end{smallmatrix}],[\begin{smallmatrix}0\\ -{\mathsf{a}}^{-1}\theta_{0}^{\dagger}\end{smallmatrix}],[\begin{smallmatrix}0\\ 1\end{smallmatrix}]\}$. A computation yields
$\iota\left(\left[\begin{smallmatrix}\tau&0\\ 0&a\tau^{\dagger}\end{smallmatrix}\right]\right)=\left[\begin{smallmatrix}\tau&0\\ 0&a\cdot{}^{t}\tau^{\dagger}\end{smallmatrix}\right],\quad\tau\in L^{\times},a\in F^{\times},$
(6.1) $\iota\left(\left[\begin{smallmatrix}1&\beta\\ 0&1\end{smallmatrix}\right]\right)=\left[\begin{smallmatrix}1_{2}&X_{\beta}\\ 0&1_{2}\end{smallmatrix}\right],\quad\beta=\beta_{2}{\mathsf{a}}+\beta_{3}\theta_{0}\in L,\ X_{\beta}:=\left[\begin{smallmatrix}-{\mathsf{a}}^{-1}\mathsf{b}\beta_{2}-{\mathsf{a}}^{-1}\mathsf{c}\beta_{3}&\beta_{2}\\ \beta_{2}&\beta_{3}\end{smallmatrix}\right],$
$\iota\left(\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]\right)=\left[\begin{smallmatrix}0&0&-{\mathsf{a}}^{-2}\mathsf{b}&{\mathsf{a}}^{-1}\\ 0&0&{\mathsf{a}}^{-1}&0\\ 0&{\mathsf{a}}&0&0\\ {\mathsf{a}}&\mathsf{b}&0&0\end{smallmatrix}\right].$
We note that $K^{\#}=\iota^{-1}(K)$ and $K_{0}^{\#}({\mathfrak{p}}^{e})=\iota^{-1}(K_{0}({\mathfrak{p}}^{e}))$. We call
$R:=\left\{\iota\left(\left[\begin{smallmatrix}\tau&0\\ 0&\tau^{\dagger}\end{smallmatrix}\right]\right)n\ \middle|\tau\in L^{\times},n\in\mathsf{N}(F)\right\}$
the Bessel subgroup of $\mathsf{G}(F)$ (with respect to $S$). For a character $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$, we can check that the map
$R\owns\iota\left(\left[\begin{smallmatrix}\tau&0\\ 0&\tau^{\dagger}\end{smallmatrix}\right]\right)\mathsf{n}(X)\mapsto\Lambda(\tau)\psi({\operatorname{tr}}{(SX)})\in{\mathbb{C}}^{1}$
defines a character on $R$. We denote this character by $\Lambda\otimes\psi_{S}$.
Let $(\pi,V_{\pi})$ be an irreducible admissible representation of $\mathsf{G}(F)$. Let $(V_{\pi}^{*})^{S,\Lambda}$ denote the space of all ${\mathbb{C}}$-linear forms $\ell:V_{\pi}\longrightarrow{\mathbb{C}}$ that satisfies
$\ell(\pi(r)\xi)=\Lambda\otimes\psi_{S}(r)\,\ell(\xi),\quad\xi\in V_{\pi},\,r\in R.$
Then, it holds that $\dim_{{\mathbb{C}}}(V_{\pi}^{*})^{S,\Lambda}\leqslant 1$ from ([27],[28]). We say that $\pi$ has a local $(S,\Lambda)$-Bessel model if $(V_{\pi}^{*})^{S,\Lambda}\not=(0)$. In this case, we can define the local $(S,\Lambda)$-Bessel model of $\pi$ as in §4.4. Moreover, when $\pi$ is spherical, we define the unramified $(S,\Lambda)$- Bessel datum $(\ell_{\pi}^{0},\xi_{\pi}^{0})\in(V_{\pi}^{*})^{S,\Lambda}\times V_{\pi}^{K}$ for $\pi$ as in §4.4, so that
$B_{\pi}^{0}(g)=\ell_{\pi}^{0}(\pi(g)\xi_{\pi}^{0}),\quad g\in\mathsf{G}({\mathbb{Q}}_{p}).$
is the unramified Bessel function such that $B_{\pi}^{0}(1_{4})=1$. Recall that all irreducible admissible representations of $\mathsf{G}(F)$ that admit local Bessel models are classified in [35], and the result is conveniently summarized in [32, Table 2]. For our cases, as we mentioned in §4.4, we may assume that $\pi$ is of type I ,IIb, IIIa, or VIb. We refer some of its properties in ([32, Table 2]);
(6.7) $\begin{array}[]{|c|c|c|c|c|}\hline\cr{{\rm type}}&{\pi}&{\dim V_{\pi}^{K}}&{\dim V_{\pi}^{K_{0}({\mathfrak{p}})}}&{{\rm Cent.\ Char.}}\\ \hline\cr{{\rm I}}&{\chi\times\chi^{\prime}\rtimes\sigma}&{1}&{4}&{\chi\chi^{\prime}\sigma^{2}}\\ \hline\cr{{\rm IIb}}&{\chi 1_{{\bf GL}_{2}}\rtimes\sigma}&{1}&{3}&{\chi^{2}\sigma}\\ \hline\cr{{\rm IIIa}}&{\chi\rtimes\sigma_{{\rm St}_{{\bf GL}_{2}}}}&{0}&{2}&{\chi\sigma^{2}}\\ \hline\cr{{\rm VIb}}&{\tau(T,|\cdot|^{-\frac{1}{2}}\sigma)}&{0}&{1}&{\sigma^{2}}\\ \hline\cr\end{array}$
Here, $\chi$, $\chi^{\prime}$, and $\sigma$ are unramified quasicharacters on $F^{\times}$ such that representations are irreducible and admit $(S,\Lambda)$-Bessel models. For $a\in F^{\times}$. For later use we set $\alpha=\chi(\varpi)$, $\beta=\chi^{\prime}(\varpi)$, and $\gamma=\sigma(\varpi)$. According to [35, Table A2], for type IIIa and type VIb, the corresponding representations are unitarizable if and only if the inducing quasi-characters are unitary, whereas, for type I, the unitarity of the inducing quasi-characters is equivalent to the corresponding representations being unitarizable and tempered. In what follows, having a global application in mind, we suppose the unitarity of all the inducing characters, i.e., $|\alpha|=|\beta|=|\gamma|=1$.
Similarly as (3.1), for a Schwartz-Bruhat function $\phi\in\mathcal{S}(L^{2})$, a character $\mu$ on $F^{\times}$, and $s\in{\mathbb{C}}$ with ${\operatorname{Re}}(s)>1$, we define a function on $\mathsf{G}^{\#}(G)$ by
$f_{\phi}^{(s,\Lambda,\mu)}(g^{\#})=\mu(\det g^{\#})|\det g^{\#}|^{s+1}\int_{L^{\times}}\phi([\begin{smallmatrix}0&1\end{smallmatrix}][\begin{smallmatrix}\tau&0\\ 0&\tau^{\dagger}\end{smallmatrix}]g^{\#})\Lambda\mu_{L}(\tau)|\tau\tau^{\dagger}|^{s+1}{\rm{d}}^{\times}\tau,$
where $\mu_{L}(\tau):=\mu(\tau\tau^{\dagger}),\tau\in L^{\times}$ and ${\rm{d}}^{\times}\tau$ is the normalized Haar measure on $L^{\times}$ such that ${\operatorname{vol}}(\mathfrak{o}_{L})=1$. For an irreducible smooth admissible representation $\pi$ of ${\mathsf{Z}}(F)\backslash\mathsf{G}(F)$, admitting $(S,\Lambda)$-Bessel model and $B\in{\mathscr{B}}(S,\Lambda)[\pi]$, define the zeta integral by
(6.8) $Z(\phi,B,s,\mu;g)=\int_{F^{\times}}\int_{K^{\#}}B(\mathsf{m}(a1_{2},a)\iota(k^{\#})g)\mu(a)|a|^{s-1}f_{\phi}^{(s,\Lambda,\mu)}(k^{\#}){\rm{d}}k^{\#}{\rm{d}}^{\times}a,\quad s\in{\mathbb{C}},{\operatorname{Re}}(s)>1.$
It is shown that $Z(\phi,B,\mu;g)$ is a rational function in $q^{-s}$ ([41, Lemma 5.3.1]). Then, the local $L$-function $L(s,\pi,\mu)$ is defined to be the unique function such that $L(s,\pi,\mu)^{-1}$ is a polynomial in $q^{-s}$ with constant term $1$ and such that $Z(\phi,B,s,\mu,1)/L(s,\pi,\mu)\in{\mathbb{C}}[q^{s},q^{-s}]$ for all $\phi$ and $B$ ([30], see also [41, Proposition 5.3.3]). By [30, Proposition 3.2], there exists an entire function $\varepsilon(\pi,s,\mu,\psi)$ such that the local functional equation
(6.9) $\varepsilon(\pi,s+1/2,\mu,\psi)\frac{Z(\phi,B,s,\mu;g)}{L(s+1/2,\pi,\mu)}=\frac{Z(\widehat{\phi},B^{\dagger},-s,\mu^{-1};g)}{L(-s+1/2,\hat{\pi},\mu^{-1})}$
holds. Here, we set $B^{\dagger}\in{\mathscr{B}}(S,\Lambda^{\dagger})[\pi]$ as $B^{\dagger}(g):=B\left(\left[\begin{smallmatrix}1&{\mathsf{a}}^{-1}\mathsf{b}&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&-{\mathsf{a}}^{-1}\mathsf{b}&1\end{smallmatrix}\right]g\right)$, $g\in\mathsf{G}(F)$, and $\Lambda^{\dagger}$ denote the Galois conjugate of $\Lambda$ defined by $\Lambda^{\dagger}(\tau):=\Lambda(\tau^{\dagger})$ for $\tau\in L^{\times}$, and $\widehat{\phi}$ is the Fourier transform of $\phi\in{\mathcal{S}}(L^{2})$ defined as
$\widehat{\phi}(x,y)=\int_{L^{2}}\phi(u,v)\psi_{L}\left(\langle\left[\begin{smallmatrix}x\\ y\end{smallmatrix}\right],\left[\begin{smallmatrix}u\\ v\end{smallmatrix}\right]\rangle\right)\,{\rm{d}}u{\rm{d}}v$
with ${\rm{d}}u$ and ${\rm{d}}v$ being the Haar measures on $L$ such that ${\operatorname{vol}}(\mathfrak{o}_{L})=1$. The $L$-functions $L(s,\pi,\mu)$ are explicitly determined for several cases as shown in [41, Table 5], which, for $\mu=1$, coincide with the local $L$-functions listed in [35, Table A.8]:
(6.15) $\begin{array}[]{|c|c|c|}\hline\cr{{\rm type}}&{\pi}&{L(s,\pi)}\\ \hline\cr{{\rm I}}&{\chi\times\chi^{\prime}\rtimes\sigma}&{(1-\alpha\beta\gamma q^{-s})^{-1}(1-\alpha\gamma q^{-s})^{-1}(1-\beta\gamma q^{-s})^{-1}(1-\gamma q^{-s})^{-1}}\\ \hline\cr{{\rm IIb}}&{\chi 1_{{\bf GL}_{2}}\rtimes\sigma}&{(1-\alpha^{2}\gamma q^{-s})^{-1}(1-\gamma q^{-s})^{-1}(1-\alpha\gamma q^{-s-\frac{1}{2}})^{-1}(1-\alpha\gamma q^{-s+\frac{1}{2}})^{-1}}\\ \hline\cr{{\rm IIIa}}&{\chi\rtimes\sigma_{{\rm St}_{{\bf GL}_{2}}}}&{(1-\alpha\gamma q^{-s-\frac{1}{2}})^{-1}(1-\gamma q^{-s-\frac{1}{2}})^{-1}}\\ \hline\cr{{\rm VIb}}&{\tau(T,|\cdot|^{-\frac{1}{2}}\sigma)}&{(1-\gamma q^{-s-\frac{1}{2}})^{-2}}\\ \hline\cr\end{array}$
We shall get explicit evaluations of the zeta integrals $Z(\phi,B,s,\mu;g)$ for particular choices of $\phi$, $B$ and $g$ as shown in the next table, which allow us to determine $\varepsilon(s,\pi,\mu,\psi)$:
(6.23) $\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr&{{\rm type}}&{\phi}&{B}&{\mu}&L/F&{g}\\ \hline\cr{\rm case\,1}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K{\rm-fixed}}&{{\rm unramified}}&-&{1_{4}}\\ \hline\cr{\rm case\,2}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}\oplus 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}}&{K{\rm-fixed}}&{{\rm ramified}}&{\rm inert}&{b}\\ \hline\cr{\rm case\,3}&{\text{I or IIb}}&{\widehat{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}\oplus 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}}}&{K{\rm-fixed}}&{{\rm ramified}}&{\rm inert}&{b}\\ \hline\cr{\rm case\,4}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K_{0}({\mathfrak{p}}){\rm-fixed}}&{{\rm unramified}}&{\rm inert}&{\eta}\\ \hline\cr{\rm case\,5}&{\text{IIIa}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K_{0}({\mathfrak{p}}){\rm-fixed}}&{{\rm unramified}}&{\rm inert}&{\eta}\\ \hline\cr{\rm case\,6}&{\text{VIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K_{0}({\mathfrak{p}}){\rm-fixed}}&{{\rm unramified}}&{\rm inert}&{\eta}\\ \hline\cr\end{array}$
Here,
(6.24) $b:=\left[\begin{smallmatrix}\varpi^{e}1_{2}&S^{\dagger}\\ 0&1_{2}\end{smallmatrix}\right]\in\mathsf{G}(F),$
and $L/F$ is said to be inert if it is an unramified field extension. For $l,m\in{\mathbb{Z}}$, set
$h(\ell,m):=\left[\begin{smallmatrix}\varpi^{\ell+2m}&&&\\ &\varpi^{\ell+m}&&\\ &&1&\\ &&&\varpi^{m}\end{smallmatrix}\right]\in\mathsf{G}(F).$
As in [32], for smooth representation $(\pi,V_{\pi})$ of $\mathsf{G}(F)$, we define an operator $T_{m,n}\in{\operatorname{End}}(V_{\pi})$ by
$T_{m,n}v:={\operatorname{vol}}(K_{0}({\mathfrak{p}}))\int_{K_{0}({\mathfrak{p}})h(l,m)K_{0}({\mathfrak{p}})}\pi(g)v\,{\rm{d}}g,\quad v\in V_{\pi}$
with ${\rm{d}}g$ the Haar measure on $\mathsf{G}(F)$ such that ${\operatorname{vol}}(K)=1$, and the Atkin-Lehner involution by
$\eta v:=\pi(\eta)v,\quad\eta:=\left[\begin{smallmatrix}&&&-1\\ &&1&\\ &\varpi&&\\ -\varpi&&&\end{smallmatrix}\right],\quad v\in V_{\pi}.$
If $\pi$ is irreducible and has the $(S,\Lambda)$-Bessel models, then for any $B\in{\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$, by [32, Lemma 4.4 (i)], we have the support conditions:
(6.25) $B(h(\ell,m))=B(h(\ell,m)s_{2})=0\quad{\rm if\ }\ell<0,$
(6.26) $B(h(\ell,m)s_{1}s_{2})=B(h(\ell,m)s_{2}s_{1}s_{2})=0\quad{\rm if\ }\ell<-1.$
6.2. Unramified computation (case1)
When $\pi$ is spherical, $\pi$ is of type I or IIb in our cases. Recall that there exists a unique element $B_{\pi}^{0}\in{\mathscr{B}}(S,\Lambda)[\pi]^{K}$ such that $B_{\pi}^{0}(1_{4})=1$. Then, the following formula for an unramified $\mu$ in its greatest generality is due to Sugano ([43, Tehorem 2.1]):
$Z(\phi,B_{\pi}^{0},s,\mu;1_{4})=L(s+1/2,\pi,\mu),\quad{\operatorname{Re}}(s)\gg 0.$
In [20, Proposition 5.9], a proof of this formula is given when $L=F\oplus L$ and $\pi$ is of type I based on the explicit formula of $B_{\pi}^{0}$ due to [6].
6.3. Unramified computation (case2 & case3)
Now we assume that $e:={\rm cond}(\mu)>0$ and that $S\in{\rm{Sym^{2}}}(F)$ satisfies the following conditions;
  • $L/F$ is an unramified field extension.
  • $S\in{\bf GL}_{2}(\mathfrak{o}),\quad-\frac{{\mathsf{d}}}{2}={\operatorname{tr}}(SS^{\dagger})\in\mathfrak{o}^{\times}.$
Then $\varpi$ remains a prime element in ${\mathfrak{p}}\mathfrak{o}_{L}$. Let $b$ as in (6.24), and set $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}\oplus 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}\in\mathcal{S}(L^{2})$. Our goal of this subsection is to calculate the zeta integrals $Z(\phi,B_{\pi}^{0},s,\mu;b)$ and $Z(\widehat{\phi},B_{\pi}^{0},s,\mu;b)$ explicitly. We remark that $Z(\phi,B_{\pi}^{0},s,\mu;1_{4})$ must be zero because the $a$-integral in (6.8) vanishes.
Definition 2 (Root number).
Let $\psi$ and $\mu$ be as above, we define the Gauss sum $W_{F}(\mu,\psi)$ by
$W_{F}(\mu,\psi):=q^{-\frac{e}{2}}\mu(\varpi)^{-e}\sum_{\alpha\in(\mathfrak{o}/{\mathfrak{p}}^{e})^{\times}}\psi(\varpi^{-e}\alpha)\mu(\alpha).$
Note that this sum has the absolute value $1$ and is independent of the choice of $\varpi$. The following lemma is more or less well-known ([45, (2.18)]).
Lemma 6.1.
Let $\psi$ and $\mu$ be as above. For $n\in{\mathbb{Z}}$,
$\int_{\mathfrak{o}^{\times}}\psi(\varpi^{n}a)\mu(a){\rm{d}}^{\times}a=\begin{cases}q^{-\frac{e}{2}+1}(q-1)^{-1}\mu(\varpi)^{e}W_{F}(\mu,\psi)\quad&(n=-e),\\ 0&(n\neq-e).\end{cases}$
Lemma 6.2.
Let $L/F$ be an unramified quadratic extension. Then, for any character $\mu$ of $F$ with $e:={\rm cond}(\mu)>0$,
$W_{L}(\mu_{L},\psi_{L})=(-1)^{e}W_{F}(\mu,\psi)^{2}.$
Proof.
For a virtual representation $V$ of the Weil group $\mathfrak{W}_{F}$ of $F$, let $\varepsilon(s,V,\psi)$ denote the local epsilon factor a la Deligne-Langlands ([46]). By ([46, (3.4.8)]), we have
(6.27) $\varepsilon(s,{\rm Ind}_{{\mathfrak{W}}_{L}}^{{\mathfrak{W}}_{F}}\rho,\psi)=\varepsilon(s,\rho,\psi_{L})$
for any $\rho$ of degree 00. Any character $\mu$ of $F^{\times}$ is viewed as a one-dimensional representation of ${\mathfrak{W}}_{F}$ by $F^{\times}\cong{\mathfrak{W}}_{F}^{\rm ab}$. If $\mu_{L}:=\mu\circ{\operatorname{N}}_{L/F}$, then ${\rm Ind}_{{\mathfrak{W}}_{L}}^{{\mathfrak{W}}_{F}}\mu_{L}\cong\mu\oplus\mu\eta_{L/F}$. If $\mu$ is unramified, then $\varepsilon(s,\mu,\psi)=1$; since $L/F$ is unramified, so is the character $\eta_{L/F}$. By these observations, the equality (6.27) holds true when $\rho={\bf 1}$. Thus, (6.27) is true when $\rho=\mu$ with $e>0$. Since $\varepsilon(s,\mu\eta_{L/F}^{j},\psi)=q^{-e(s-1/2)}W_{F}(\bar{\mu}\eta_{L/F}^{j},\psi)$ for $j=0,1$ and $\varepsilon(s,\mu_{L},\psi_{L})=q^{-2e(s-1/2)}W_{L}(\bar{\mu}_{L},\psi_{L})$, we obtain $W_{L}(\mu_{L},\psi_{L})=W_{F}(\mu,\psi)W_{F}(\mu\eta_{L/F},\psi)$. By $e>0$ and $\eta_{L/F}(\varpi)=-1$, we have $W_{F}(\mu\eta_{L/F},\psi)=\eta_{L/F}(\varpi)^{e}W_{F}(\mu,\psi)=(-1)^{e}W_{F}(\mu,\psi)$. ∎
Lemma 6.3.
For $k^{\#}=\left[\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right]\in K^{\#}$, we have
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})$ =$=\begin{cases}q^{-2e+2}(q^{2}-1)^{-1}\mu_{L}(d)^{-1}\mu(\det k^{\#})\quad&(c\in{\mathfrak{p}}^{e}\mathfrak{o}_{L},d\in 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}),\\ 0&({\rm otherwise}),\end{cases}$
$f_{\widehat{\phi}}^{(s,\Lambda,\mu)}(k^{\#})$ =$=\begin{cases}q^{e(2s-3)+2}(q^{2}-1)^{-1}\Lambda(\varpi)^{-e}\mu_{L}(c)^{-1}W_{L}(\mu_{L},\psi_{L})\quad&(c\in\mathfrak{o}_{L}^{\times}),\\ 0&(c\notin\mathfrak{o}_{L}^{\times}).\end{cases}$
Proof.
First, we prove the first equation. By a direct calculation, we have
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})$ $=\mu(\det k^{\#})\int_{L^{\times}}\phi([\begin{smallmatrix}c\tau^{\dagger}&d\tau^{\dagger}\end{smallmatrix}])\Lambda\mu_{L}(\tau)|\tau\tau^{\dagger}|^{s+1}{\rm{d}}^{\times}\tau$
$=\mu(\det k^{\#})\int_{L^{\times}}\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}}(c\tau^{\dagger})\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}(d\tau^{\dagger})\Lambda\mu_{L}(\tau)|\tau\tau^{\dagger}|^{s+1}{\rm{d}}^{\times}\tau.$
Since $c\in\mathfrak{o}_{L}^{\times}$ or $d\in\mathfrak{o}_{L}^{\times}$, it is easy to check that $(c\tau^{\dagger},d\tau^{\dagger})\in{\mathfrak{p}}^{e}\mathfrak{o}_{L}\times(1+{\mathfrak{p}}^{e}\mathfrak{o}_{L})$ if and only if
$c\in{\mathfrak{p}}^{e}\mathfrak{o}_{L},\quad d\in\mathfrak{o}_{L}^{\times},\quad\tau\in(d^{\dagger})^{-1}(1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}).$
Hence, if $c\in{\mathfrak{p}}^{e}\mathfrak{o}_{L}$, $d\in\mathfrak{o}_{L}^{\times}$, we get
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})$ $=\mu(\det k^{\#})\cdot{\operatorname{vol}}(1+{\mathfrak{p}}^{e}\mathfrak{o}_{L})\cdot\Lambda\mu_{L}((d^{\dagger})^{-1})$
$=q^{-2e+2}(q^{2}-1)^{-1}\mu_{L}(d)^{-1}\mu(\det k^{\#}).$
Next, we prove the second equation. It is easy to check that
$\widehat{\phi}(x,y)=q^{-4e}\psi_{L}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{{}^{e}}\mathfrak{o}}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{{}^{e}}\mathfrak{o}}(y),\quad x,y\in L.$
Hence we have
$f_{\widehat{\phi}}^{(s,\Lambda,\mu)}(k^{\#})$ $=\mu(\det k^{\#})\int_{L^{\times}}\widehat{\phi}([\begin{smallmatrix}c\tau^{\dagger}&d\tau^{\dagger}\end{smallmatrix}])\Lambda\mu_{L}(\tau)|\tau\tau^{\dagger}|_{p}^{s+1}{\rm{d}}^{\times}\tau$
$=q^{-4e}\mu(\det k^{\#})\int_{L^{\times}}\psi_{L}(c\tau^{\dagger})\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{-e}\mathfrak{o}_{L}}(c\tau^{\dagger})\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{-e}\mathfrak{o}_{L}}(d\tau^{\dagger})\Lambda\mu_{L}(\tau)|\tau\tau^{\dagger}|_{p}^{s+1}{\rm{d}}^{\times}\tau$
(6.28) $=q^{-4e}\mu(\det k^{\#})\sum_{n{\geqslant}-e}\Lambda\mu_{L}(\varpi)^{n}q^{-2n(s+1)}\int_{\mathfrak{o}_{L}^{\times}}\psi_{L}(c\varpi^{n}\tau)\mu_{L}(\tau){\rm{d}}^{\times}\tau.$
We put $c=c_{0}\varpi^{m}$, $c_{0}\in\mathfrak{o}_{L}^{\times}$. Then by Lemma 6.1, with $(F,\mu,\psi)$ replaced by $(L,\Lambda\mu_{L},\psi_{L})$, we have
$\int_{\mathfrak{o}_{L}^{\times}}\psi_{L}(c\varpi^{n}\tau)\mu_{L}(\tau){\rm{d}}^{\times}\tau$ $=\mu_{L}(c_{0})^{-1}\int_{\mathfrak{o}_{L}^{\times}}\psi_{L}(\varpi^{n+m}\tau)\mu_{L}(\tau){\rm{d}}^{\times}\tau$
$=\delta_{n,-e-m}q^{-e+2}(q^{2}-1)^{-1}\mu_{L}(c_{0}^{-1}\varpi)^{e}W_{L}(\mu_{L},\psi_{L}).$
[…]
$f_{\widehat{\phi}}^{(s,\Lambda,\mu)}(k^{\#})$ $=q^{-5e+2}(q^{2}-1)^{-1}\Lambda(\varpi)^{-e}\mu_{L}(c_{0})^{-1}W_{L}(\mu_{L},\psi_{L})\sum_{n{\geqslant}-e}\delta_{n,-e-m}\Lambda\mu_{L}(\varpi)^{n+e}q^{-2n(s+1)}$
=$=\begin{cases}q^{e(2s-3)+2}(q^{2}-1)^{-1}\Lambda(\varpi)^{-e}\mu_{L}(c)^{-1}W_{L}(\mu_{L},\psi_{L})\quad&(m=0),\\ 0&(m>0).\end{cases}\qed$
Lemma 6.4.
For $u\in\mathfrak{o}^{\times}$,
$\sum_{\begin{subarray}{c}\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}\\ u+{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}\end{subarray}}\mu(u+{\operatorname{N}}_{L/F}(\eta))=(-1)^{e}q^{e}\mu(u).$
Proof.
Because $L/F$ is an unramified extension, the norm map $\mathfrak{o}_{L}^{\times}\owns\eta\mapsto{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}$ is surjective. Thus there exists $v\in\mathfrak{o}_{L}^{\times}$ such that $u={\operatorname{N}}_{L/F}(v)$. By replacing $\eta$ by $v\eta$, without loss of generality, we can assume that $u=1$.
For an integer $i$ with $10{\leqslant}i{\leqslant}e+$ 1, we set
$\overline{\mathfrak{o}_{L,i}}=\begin{cases}\left\{\eta\in\left(\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}\right)^{\times}\ \middle|\ 1+{\operatorname{N}}_{L/F}(\eta)\not\equiv 0\ (\text{mod }{\mathfrak{p}})\right\}\quad&(i=0),\\ {\mathfrak{p}}^{i}\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}&(0<i{\leqslant}e),\\ \varnothing&(i=e+1)\end{cases}$
and
$\overline{{\mathcal{U}}_{i}}=\begin{cases}\mathfrak{o}^{\times}/(1+{\mathfrak{p}}^{e})\quad&(i=0),\\ (1+{\mathfrak{p}}^{i})/(1+{\mathfrak{p}}^{e})\quad&(0<i{\leqslant}e),\\ \varnothing&(i=e+1).\end{cases}$
When $20{\leqslant}i<e/$ 2, we can check that
$\overline{\mathfrak{o}_{L,i}}\setminus\overline{\mathfrak{o}_{L,i+1}}\owns\eta\ (\text{mod }{\mathfrak{p}}^{e}\mathfrak{o}_{L})\mapsto 1+N_{L/F}(\eta)\ (\text{mod }{\mathfrak{p}}^{e})\in\overline{{\mathcal{U}}_{2i}}\setminus\overline{{\mathcal{U}}_{2i+1}}$
is a surjective map having the fiber of cardinality $q^{e-1}(q+1)$ at each point.
When $e/2{\leqslant}i{\leqslant}e$, it is immediate that
$\mu(1+{\operatorname{N}}_{L/F}(\eta))=1,\quad\eta\in\overline{\mathfrak{o}_{L,i}}.$
Hence we have
$\sum_{\begin{subarray}{c}\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}\\ 1+{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}\end{subarray}}\mu(1+{\operatorname{N}}_{L/F}(\eta))$ $=\sum_{i=0}^{e}\sum_{\eta\in\overline{\mathfrak{o}_{L,i}}\setminus\overline{\mathfrak{o}_{L,i+1}}}\mu(1+{\operatorname{N}}_{L/F}(\eta))$
$=q^{e-1}(q+1)\sum_{0{\leqslant}i<e/2}\sum_{\varepsilon\in\overline{{\mathcal{U}}_{2i}}\setminus\overline{{\mathcal{U}}_{2i+1}}}\mu(\varepsilon)+\sum_{e/2{\leqslant}i{\leqslant}e}\#(\overline{\mathfrak{o}_{L,i}}\setminus\overline{\mathfrak{o}_{L,i+1}}).$
Then the Lemma follows from
$\sum_{\varepsilon\in\overline{{\mathcal{U}}_{2i}}\setminus\overline{{\mathcal{U}}_{2i+1}}}\mu(\varepsilon)=\begin{cases}0&(0{\leqslant}i<\frac{e-1}{2}),\\ -1&(i=\frac{e-1}{2})\\ \end{cases}$
and
$\sum_{e/2{\leqslant}i{\leqslant}e}\#(\overline{\mathfrak{o}_{L,i}}\setminus\overline{\mathfrak{o}_{L,i+1}})=q^{2e-2\lceil e/2\rceil}.$
where $\lceil\cdot\rceil:{\mathbb{R}}\to{\mathbb{Z}}$ is the ceiling function. ∎
Lemma 6.5.
Let $B\in{\mathscr{B}}(S,\Lambda)[\pi]^{K}$ and $\eta\in\mathfrak{o}$ with $\frac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\in\varpi^{j}\mathfrak{o}^{\times}$ $(0{\leqslant}j{\leqslant}e)$.
  • (i)
    If we set $Y_{\eta}=-{\mathsf{a}}^{2}S^{\dagger}+X_{\eta}$, then we have
    $\det Y_{\eta}=-\tfrac{{\mathsf{a}}^{4}{\mathsf{d}}}{4}-{\mathsf{a}}^{-2}{\operatorname{N}}_{L/F}(\eta).$
  • (ii)
    There exist $\tau\in\mathfrak{o}_{L}^{\times}$ and $A\in{\bf GL}_{2}(\mathfrak{o})$ such that
    $Y_{\eta}=\tau\left[\begin{smallmatrix}\varpi^{j}&0\\ 0&1\end{smallmatrix}\right]A.$
  • (iii)
    For $n\in{\mathbb{Z}}$ and $a\in\mathfrak{o}^{\times}$,
    $B\left(\mathsf{m}(\varpi^{n}a1_{2},\varpi^{n}a)\iota(\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right])b\right)$
    $=\psi\left(-\varpi^{n}a\cdot\tfrac{{\mathsf{a}}^{4}{\mathsf{d}}}{2}\left(\tfrac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\right)^{-1}\right)\cdot B\left(h(e+n-2j,j)\right).$
Proof.
(i) By noting that $\det S=-{\mathsf{d}}/4$, $\det X_{\eta}=-{\mathsf{a}}^{-2}{\operatorname{N}}_{L/F}(\eta)$, and ${\operatorname{tr}}(SX_{\eta})={\operatorname{tr}}(X_{\eta}S^{\dagger})=0$, we obtain
$\det Y_{\eta}=\tfrac{1}{2}{\operatorname{tr}}(Y_{\eta}Y_{\eta}^{\dagger})=\tfrac{1}{2}{\operatorname{tr}}({\mathsf{a}}^{4}SS^{\dagger}+X_{\eta}^{\dagger}X_{\eta})={\mathsf{a}}^{4}\det S+\det(X_{\eta})=-\tfrac{{\mathsf{a}}^{4}{\mathsf{d}}}{4}-{\mathsf{a}}^{-2}{\operatorname{N}}_{L/F}(\eta).$
(ii) By (i) and the assumption that $\frac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\in\varpi^{j}\mathfrak{o}^{\times}$, we have $Y_{\eta}\in{\bf GL}_{2}(F)$. The disjoint union
${\bf GL}_{2}(F)=\sqcup_{m{\geqslant}0}L^{\times}\left[\begin{smallmatrix}\varpi^{m}&0\\ 0&1\end{smallmatrix}\right]{\bf GL}_{2}(\mathfrak{o}_{L}).$
from [43, Lemma 2–4] implies that there exist $m\in{\mathbb{Z}}_{{\geqslant}0}$, $\ell\in{\mathbb{Z}}$, $\tau\in\mathfrak{o}_{L}^{\times}$, $A\in{\bf GL}_{2}(\mathfrak{o})$, such that
$Y_{\eta}=(\varpi^{\ell}\tau)\left[\begin{smallmatrix}\varpi^{m}&0\\ 0&1\end{smallmatrix}\right]A=\tau\left[\begin{smallmatrix}\varpi^{m+\ell}&0\\ 0&\varpi^{\ell}\end{smallmatrix}\right]A.$
We now prove that $m=j$ and $\ell=0$. Because $\tau\in{\bf GL}_{2}(\mathfrak{o})$, the Smith normal form of $Y_{\eta}$ is $\left[\begin{smallmatrix}\varpi^{m+\ell}&0\\ 0&\varpi^{\ell}\end{smallmatrix}\right]$. Hence by the theory of the Smith normal form, we only have to show that the elements of $Y_{\eta}$ are coprime. This follows from the identity ${\operatorname{tr}}(Y_{\eta}S)=\frac{{\mathsf{a}}^{2}{\mathsf{d}}}{2}$ and the assumption $\frac{{\mathsf{d}}}{2}\in\mathfrak{o}^{\times}$.
(iii) By using (6.1), we can check that
$\iota(\left[\begin{smallmatrix}1&0\\ -\eta^{\dagger}&1\end{smallmatrix}\right])=\left[\begin{smallmatrix}1_{2}&0\\ X_{\eta}^{\dagger}&1_{2}\end{smallmatrix}\right],\quad\iota(\left[\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right])b\iota(\left[\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right])=\left[\begin{smallmatrix}1_{2}&0\\ -{\mathsf{a}}^{2}S&\varpi^{e}1_{2}\end{smallmatrix}\right].$
Hence we have
$\mathsf{m}(\varpi^{n}a1_{2},\varpi^{n}a)\iota(\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right])b=\left[\begin{smallmatrix}\varpi^{n}a1_{2}&0\\ Y_{\eta}^{\dagger}&\varpi^{e}1_{2}\end{smallmatrix}\right]\iota(\left[\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right]).$
By (ii), we may choose elements $\tau\in\mathfrak{o}_{L}^{\times}$ and $A\in{\bf GL}_{2}(\mathfrak{o})$ satisfying $Y_{\eta}=\tau\left[\begin{smallmatrix}\varpi^{j}&0\\ 0&1\end{smallmatrix}\right]A$. Then, a computation shows that
$\left[\begin{smallmatrix}\varpi^{n}a1_{2}&0\\ Y_{\eta}^{\dagger}&\varpi^{e}1_{2}\end{smallmatrix}\right]=\mathsf{n}(\varpi^{n}a(\det Y_{\eta})^{-1}Y_{\eta})\left[\begin{smallmatrix}\tau&0\\ 0&{}^{t}\tau^{\dagger}\end{smallmatrix}\right]h(e+n-2j,j)k$
where $k=\left[\begin{smallmatrix}A&0\\ 0&{}^{t}A^{\dagger}\end{smallmatrix}\right]\left[\begin{smallmatrix}a\varpi^{j}(\det Y_{\eta})^{-1}1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\left[\begin{smallmatrix}0&-1_{2}\\ 1_{2}&\varpi^{e}(\det Y_{\eta})^{-1}Y_{\eta}\end{smallmatrix}\right]\in K$. Since $\Lambda$ is unramified and $B$ is $K$-invariant, we have
$B\left(\mathsf{m}(\varpi^{n}a1_{2},\varpi^{n}a)\iota(\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right])b\right)$
$=\psi(\varpi^{n}a(\det Y_{\eta})^{-1}\cdot\tfrac{{\mathsf{a}}^{2}{\mathsf{d}}}{2})\cdot B\left(h(e+n-2j,j)\right)$
$=\psi\left(-\varpi^{n}a\cdot\tfrac{{\mathsf{a}}^{4}{\mathsf{d}}}{2}\left(\tfrac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\right)^{-1}\right)\cdot B\left(h(e+n-2j,j)\right).\qed$
Proposition 6.6.
Suppose that $\pi$ is of type I or IIb. Let $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ be characters such that $\text{cond}(\mu)=e>0$ and $\Lambda$ is unramified. Suppose that $S=\left[\begin{smallmatrix}{\mathsf{a}}&\mathsf{b}/2\\ \mathsf{b}/2&\mathsf{c}\end{smallmatrix}\right]\in{\rm{Sym^{2}}}(F)\cap{\bf GL}_{2}(\mathfrak{o})$ satisfies that $L/F$ is an unramified field extension and $2d\in\mathfrak{o}^{\times}$. When $\phi=1_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}\oplus 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}$ and ${\operatorname{Re}}(s)>1$, we have
$Z(\phi,B_{\pi}^{0},s,\mu;b)=q^{e(s-\frac{11}{2})+5}(q^{4}-1)^{-1}(q-1)^{-1}\mu(-2^{-1}{\mathsf{d}})^{-1}W_{F}(\mu,\psi),$
$Z(\widehat{\phi},B_{\pi}^{0},s,\mu;b)=(-1)^{e}q^{e(3s-\frac{11}{2})+5}(q^{4}-1)^{-1}(q-1)^{-1}\Lambda(\varpi)^{-e}\mu(-2^{-1}{\mathsf{a}}^{2})W_{L}(\mu_{L},\psi_{L})W_{F}(\mu,\psi).$
Proof.
Let $K^{\#}({\mathfrak{p}}^{e})=\{k^{\#}\in K^{\#}\ |\ k^{\#}\equiv 1_{2}\ ({\rm mod}\ {\mathfrak{p}}^{e}\mathfrak{o}_{L})\}$ be the principal congruence subgroup of level ${\mathfrak{p}}^{e}\mathfrak{o}_{L}$. If $\phi\in\mathcal{S}(L^{2})$ is right $K^{\#}({\mathfrak{p}}^{e})$-invariant, the function
$K^{\#}\owns k^{\#}\mapsto B(\mathsf{m}(a1_{2},1)\iota(k^{\#})b)f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})\in{\mathbb{C}}$
is left $B^{\#}(\mathfrak{o})$-invariant and right $K^{\#}({\mathfrak{p}}^{e})$-invariant because $b^{-1}\iota(K^{\#}({\mathfrak{p}}^{e}))b\subset K$. Since $K^{\#}({\mathfrak{p}}^{e})$ is a normal subgroup of $K^{\#}$ and $B^{\#}(\mathfrak{o})K^{\#}({\mathfrak{p}}^{e})=K_{0}^{\#}({\mathfrak{p}}^{e})$, we have
$Z(\phi,B_{\pi}^{0},s,\mu;b)$
$=\sum_{[\gamma]\in B^{\#}(\mathfrak{o})\backslash K^{\#}/K^{\#}({\mathfrak{p}}^{e})}{\operatorname{vol}}(B^{\#}(\mathfrak{o})\gamma K^{\#}({\mathfrak{p}}^{e}))f_{\phi}^{(s,\Lambda,\mu)}(\gamma)\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\gamma)b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
$=[K^{\#}:K_{0}^{\#}({\mathfrak{p}}^{e})]^{-1}\sum_{[\gamma]\in K_{0}^{\#}({\mathfrak{p}}^{e})\backslash K^{\#}}f_{\phi}^{(s,\Lambda,\mu)}(\gamma)\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\gamma)b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
(6.29) $=q^{-2e+2}(q^{2}+1)^{-1}\sum_{[\gamma]\in K_{0}^{\#}({\mathfrak{p}}^{e})\backslash K^{\#}}f_{\phi}^{(s,\Lambda,\mu)}(\gamma)\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\gamma)b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a.$
A complete set of representative for $K^{\#}({\mathfrak{p}}^{e})\backslash K^{\#}$ is given by
(6.30) $\left[\begin{smallmatrix}1&0\\ \xi&1\end{smallmatrix}\right]\ (\xi\in{\mathfrak{p}}\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}),\quad\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right]\ (\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}).$
Now we prove the first equation. By Lemma 6.3, when $\gamma$ runs through the above elements, then we have $f_{\phi}^{(s,\Lambda,\mu)}(1_{2})=q^{-2e+2}(q^{2}-1)^{-1}$ and $f_{\phi}^{(s,\Lambda,\mu)}(\gamma)=0$ if $\gamma\neq 1$. By substituting this in (6.29) and Lemma 6.1, we have
$Z(\phi,B_{\pi}^{0},s,\mu;b)$
$=q^{-4e+4}(q^{4}-1)^{-1}\int_{F^{\times}}B_{\pi}^{0}\left(\left[\begin{smallmatrix}a\varpi^{e}1_{2}&aS^{\dagger}\\ 0&1_{2}\end{smallmatrix}\right]\right)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
$=q^{-4e+4}(q^{4}-1)^{-1}\int_{F^{\times}}B_{\pi}^{0}\left(\left[\begin{smallmatrix}a\varpi^{e}1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\right)\psi(a{\operatorname{tr}}(SS^{\dagger}))\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
$=q^{-4e+4}(q^{4}-1)^{-1}\mu(-2^{-1}{\mathsf{d}})^{-1}\sum_{n\in{\mathbb{Z}}}B_{\pi}^{0}\left(\left[\begin{smallmatrix}\varpi^{e+n}1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\right)\mu(\varpi)^{n}q^{-n(s-1)}\int_{\mathfrak{o}^{\times}}\psi(\varpi^{n}a)\mu(a){\rm{d}}^{\times}a$
$=q^{e(s-\frac{11}{2})+5}(q^{4}-1)^{-1}(q-1)^{-1}\mu(-2^{-1}{\mathsf{d}})^{-1}W_{F}(\mu,\psi).$
Next, we prove the second equation. When we replace $\phi$ by $\widehat{\phi}$ in (6.29) and consider the representative (6.30), we can ignore the contribution of $\xi$-terms in the summation from Lemma 6.3. Hence, by noting the $K$-invariance of $B$, we have
$Z(\widehat{\phi},B_{\pi}^{0},s,\mu;b)$ $=q^{e(2s-5)+4}(q^{4}-1)^{-1}\Lambda(\varpi)^{-e}W_{L}(\mu_{L},\psi_{L})$
(6.31) ×$\times\sum_{\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}}\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right])b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a.$
By replacing $\eta$ by $\eta+\varpi^{e}u$ for some $u\in\mathfrak{o}_{L}$ if we need, we may assume that ${\mathsf{a}}^{6}{\mathsf{d}}+{\operatorname{N}}_{L/F}(\eta)\in\varpi^{j}\mathfrak{o}_{L}^{\times}$ with $0{\leqslant}j{\leqslant}e$. Then, Lemmas 6.1 and 6.5 (iii) imply
$\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right])b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
$=\sum_{n\in{\mathbb{Z}}}B_{\pi}^{0}(h(e+n-2j,j))\mu(\varpi)^{n}q^{-n(s-1)}\int_{\mathfrak{o}^{\times}}\psi\left(-\varpi^{n}a\cdot\tfrac{{\mathsf{a}}^{4}{\mathsf{d}}}{2}\left(\tfrac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\right)^{-1}\right)\mu(a){\rm{d}}^{\times}a$
$=\mu\left(-2{\mathsf{a}}^{-4}{\mathsf{d}}^{-1}\varpi^{-j}\left(\tfrac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\right)\right)\sum_{n\in{\mathbb{Z}}}B_{\pi}^{0}(h(e+n-2j,j))\mu(\varpi)^{n}q^{-n(s-1)}\int_{\mathfrak{o}^{\times}}\psi(\varpi^{n-j}a)\mu(a){\rm{d}}^{\times}a$
$=q^{-\frac{e}{2}+1+(e-j)(s-1)}(q-1)^{-1}\mu(-2{\mathsf{a}}^{-4}{\mathsf{d}}^{-1})\mu\left(\tfrac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\right)W_{F}(\mu,\psi)\cdot B_{\pi}^{0}(h(-j,j)).$
The last equation is equal to 00 unless $j=0$ from (6.25). By substituting this into (6.31) and Lemma 6.4, we have that $Z(\phi,B_{\pi}^{0},s,\mu;b)$ equals
$q^{e(3s-\frac{13}{2})+5}(q^{4}-1)^{-1}(q-1)^{-1}\Lambda(\varpi)^{-e}\mu(-2{\mathsf{a}}^{-4}{\mathsf{d}}^{-1})$
×$\times W_{L}(\mu_{L},\psi_{L})W_{F}(\mu,\psi)\sum_{\begin{subarray}{c}\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}\\ {\mathsf{a}}^{6}{\mathsf{d}}+{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}\end{subarray}}\mu\left(\tfrac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\right)$
$=(-1)^{e}q^{e(3s-\frac{11}{2})+5}(q^{4}-1)^{-1}(q-1)^{-1}\Lambda(\varpi)^{-e}\mu(-2^{-1}{\mathsf{a}}^{2})W_{L}(\mu_{L},\psi_{L})W_{F}(\mu,\psi).\qed$
Corollary 6.7.
Let $\pi$, $S$, $\mu$, and $\Lambda$ be as in Proposition 6.6, then
$\varepsilon(\pi,s,\mu,\psi)$ $=(-1)^{e}q^{4e(\frac{1}{2}-s)}\Lambda(\varpi)^{-e}\mu(-{\mathsf{a}}^{-2}{\mathsf{d}})\overline{W_{L}(\mu_{L},\psi_{L})W_{F}(\mu,\psi)^{2}}$
$=q^{4e(\frac{1}{2}-s)}\Lambda(\varpi)^{-e}\mu(-{\mathsf{a}}^{-2}{\mathsf{d}})\overline{W_{F}(\mu,\psi)^{4}}.$
Proof.
The first equality follows from Proposition 6.6 and (6.9); the second equality is due to Lemma 6.2. ∎
6.4. Computation for old forms of type ${\rm I}$ and ${\rm IIb}$ (case4)
Now we assume that $\pi$ has the trivial character. Then $\pi$ has the $(S,\Lambda)$-Bessel model if and only if $\Lambda=1$. Recall that the dimension of $V_{\pi}^{K_{0}({\mathfrak{p}})}$ is equal to $4$ or $3$ according as $\pi$ is of type I or type IIb.
Proposition 6.8.
Let $\pi$ be a smooth admissible irreducible representation of type I or type IIb with trivial central character. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ is an unramified character. When $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$ and ${\operatorname{Re}}(s)>1$, we have
$Z(\phi,B,s,\mu;\eta)=\frac{1}{q^{2}+1}L(s+1/2,\pi,\mu)$
×$\times\left[\eta B+q^{-1}T_{1,0}B+\left\{\mu(\varpi)^{-1}q^{s+1}+\mu(\varpi)q^{-s+1}-{\operatorname{tr}}((q^{-1}T_{1,0}+\eta)|_{V_{\pi}^{K_{0}({\mathfrak{p}})}})\right\}B\right](1_{4})$
for $B\in{\mathscr{B}}(S,1)[\pi]^{K_{0}({\mathfrak{p}})}$.
Proof.
We will give the proof of the case that $\pi$ is of type I. In the case of type IIb, the proof is similar.
Let $\{B_{j}^{*}\}_{1{\leqslant}j{\leqslant}4}\subset{\mathscr{B}}(S,1)[\pi]^{K_{0}({\mathfrak{p}})}$ be an eigenbasis of $T_{1,0}$ with eigenvalues $\{\mu_{j}\}_{1{\leqslant}i{\leqslant}4}$ and $(\eta_{ij})_{1{\leqslant}i,j{\leqslant}4}$ be the representation matrix of $\eta|_{K_{0}({\mathfrak{p}})}$ with respect to $\{B_{j}^{*}\}_{1{\leqslant}i{\leqslant}4}$. Then, the first relation in (6.40) shows that
$B_{j}^{*}(h(l+1,0))=\mu_{j}q^{-3}B_{j}^{*}(h(l,0)),\quad l\in{\mathbb{Z}}_{{\geqslant}0}.$
It is easy to check that
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})=L(s+1,\Lambda\mu_{L}),\quad k^{\#}\in K^{\#}.$
By the left $K_{0}^{\#}({\mathfrak{p}})$-coset decomposition of $K^{\#}$;
(6.32) $K^{\#}=K_{0}^{\#}({\mathfrak{p}})\sqcup\left(\sqcup_{\xi\in\mathfrak{o}_{L}/{\mathfrak{p}}\mathfrak{o}_{L}}\left[\begin{smallmatrix}1&\xi\\ 0&1\end{smallmatrix}\right]\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]K_{0}^{\#}({\mathfrak{p}}^{e})\right),$
the zeta integral (6.8) becomes
(6.33) $Z(\phi,B,s,\mu;\eta)$ $=Z(\phi,\eta B,s,\mu;1_{4})$
$=[K^{\#}:K_{0}^{\#}({\mathfrak{p}})]^{-1}L(s+1,\Lambda\mu_{L})$
×$\times\int_{F^{\times}}\left\{\eta B(\mathsf{m}(a1_{2},a))+\sum_{\xi\in\mathfrak{o}_{L}/{\mathfrak{p}}\mathfrak{o}_{L}}\eta B\left(\mathsf{m}(a1_{2},a)\iota\left(\left[\begin{smallmatrix}1&\xi\\ 0&1\end{smallmatrix}\right]\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]\right)\right)\right\}\mu^{-1}(a)|a|^{s-1}{\rm{d}}^{\times}a$
$=\frac{1}{q^{2}+1}L(s+1,\Lambda\mu_{L})$
×$\times\left\{\sum_{l\in{\mathbb{Z}}}\eta B(h(l,0))+q^{2}\Lambda(\varpi)^{-1}B(h(l+1,0))\right\}\mu(\varpi)^{l}q^{-l(s-1)}.$
Since $B(h(l,0))=0$ if $l<0$ by (6.25), the right hand side equals
(6.34) $\frac{L(s+1,\Lambda\mu_{L})}{q^{2}+1}\left\{\sum_{l=0}^{\infty}\eta B(h(l,0))\mu(\varpi)^{l}q^{-l(s-1)}+q^{s+1}\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}\sum_{l=0}^{\infty}B(h(l,0))\mu(\varpi)^{l}q^{-l(s-1)}\right\}.$
By substituting $B=B_{j}^{*}$ in (6.34) and $\Lambda=1$,
$Z(\phi,B_{j}^{*},s,\mu;\eta)=\frac{L(s+1,\Lambda\mu_{L})}{q^{2}+1}\left\{\sum_{i=1}^{4}\frac{\eta_{ij}B_{i}^{*}(1_{4})}{1-\mu_{j}\mu(\varpi)q^{-(s+2)}}+\frac{\mu(\varpi)^{-1}q^{s+1}B_{j}^{*}(1_{4})}{1-\mu_{i}\mu(\varpi)q^{-(s+2)}}\right\}.$
Thus, we may put
$Z(\phi,B_{j}^{*},s,\mu;\eta)$ $=\frac{1}{q^{2}+1}\prod_{j=1}^{4}(1-\mu_{j}q^{-2}X)^{-1}$
×$\times\frac{A_{0}+A_{1}X+A_{2}X^{2}+A_{3}X^{3}+A_{4}X^{4}}{X(1-q^{-2}X^{2})},\quad(X=\mu(\varpi)q^{-s},A_{i}\in{\mathbb{C}},1{\leqslant}i{\leqslant}4).$
By noting that $\{\mu_{1},\mu_{2},\mu_{3},\mu_{4}\}=\{\alpha\beta\gamma q^{3/2},\beta\gamma q^{3/2},\alpha\gamma q^{3/2},\gamma q^{3/2}\}$ from the table (6.15), we have
(6.35) $(q^{2}+1)\cdot\frac{Z(\phi,B_{j}^{*},s,\mu;1_{4})}{L(s+1/2,\pi,\mu)}=\frac{A_{0}+A_{1}X+A_{2}X^{2}+A_{3}X^{3}+A_{4}X^{4}}{X(1-q^{-2}X^{2})}$
A direct computation shows that
(6.36) $A_{0}=qB_{j}^{*}(1_{4}),$
$A_{1}=\sum_{i=1}^{4}\eta_{ij}B_{i}^{*}(1_{4})+qB_{j}^{*}(1_{4})\left\{\left(\sum_{k=1}^{4}q^{-2}\mu_{k}\right)-q^{-2}\mu_{i}\right\}$
(6.37) $=\eta B_{j}^{*}(1_{4})+q^{-1}{\operatorname{tr}}(T_{1,0}|_{V_{\pi}^{K_{0}({\mathfrak{p}})}})\cdot B_{j}^{*}(1_{4})-q^{-1}\cdot T_{1,0}B_{j}^{*}(1_{4}).$
The functional equation (6.9) and the fact $\varepsilon(\pi,s,\mu,\psi)=1$ from ([30]) show that the right hand side of (6.35) is invariant under the transformation $X\to X^{-1}$. By comparing the coefficients, we have
(6.38) $A_{2}=q^{-2}(q^{2}-1)A_{0},\quad A_{3}=-q^{-2}A_{1},\quad A_{4}=-q^{-2}A_{0}$
Substituting (6.36), (6.37), and (6.38) into (6.35), we get
$(q^{2}+1)\cdot\frac{Z(\phi,B_{j}^{*},s,\mu;\eta)}{L(s+1/2,\pi,\mu)}$
$=\frac{A_{0}+A_{1}X+q^{-2}(q^{2}-1)A_{0}X^{2}-q^{-2}A_{1}X^{3}-q^{-2}A_{0}X^{4}}{X(1-q^{-2}X^{2})}$
$=A_{0}(X+X^{-1})+A_{1}$
$=\left[\eta B+q^{-1}T_{1,0}B+\left\{\mu(\varpi)^{-1}q^{s+1}+\mu(\varpi)q^{-s+1}-{\operatorname{tr}}(q^{-1}T_{1,0}|_{V_{\pi}^{K_{0}({\mathfrak{p}})}})\right\}B\right](1_{4}).\qed$
6.5. Computation for newforms of type ${\rm III}a$ (case5)
By [32, §9.1], there exists a basis $\{B_{1},B_{2}\}$ of ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ such that
$T_{1,0}B_{1}=\alpha\gamma qB_{1},\quad T_{0,1}B_{2}=\gamma qB_{2},$
(6.39) $T_{0,1}B_{1}=\alpha\gamma^{2}(\alpha q+1)qB_{1},\quad T_{0,1}B_{2}=\alpha\gamma^{2}(\alpha^{-1}q+1)qB_{2},$
$\eta B_{1}=\alpha\gamma B_{2},\quad\eta B_{2}=\gamma B_{1}.$
We now consider the values of $B_{1}$ and $B_{2}$ at the identity element. We remark that [32, Lemma 9.2] ensures its non-vanishing and
$\Lambda(\varpi)=\alpha\gamma^{2}$
by the condition of central character of $\pi$.
Lemma 6.9.
Let $S\in{\rm{Sym^{2}}}(F)$ such that $L/F$ is an unramified field extension. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ are unramified characters. Then we have
$B_{2}(1_{4})=\alpha^{-1}B_{1}(1_{4})\not=0.$
Proof.
For $B\in{\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ and any non-negative integer $l$, by Lemmas 5.1 and 5.3 in [32], we get the following relations;
$T_{1,0}B(h(l,0))=q^{3}B(h(l+1,0)),$
$T_{1,0}B(h(l,0)s_{2})=q^{2}(q-1)B(h(l+1,0))+q^{2}B(h(l-1,1)s_{1}s_{2}),$
$T_{1,0}B(h(l,0)s_{2}s_{1}s_{2})=q^{2}(q-1)B(h(l+1,0))+\mu(\varpi)B(h(l-1,0)s_{2}s_{1}s_{2})$
(6.40) $+(q^{2}-1)B(h(l-1,1)s_{1}s_{2}),$
$T_{0,1}B(h(l,0))=q^{3}(q+1)B(h(l,1)),$
$T_{0,1}B(h(l,0)s_{2})=q^{4}B(h(l,1)s_{1}s_{2})+q\mu(\varpi)B(h(l-2,1)s_{1}s_{2})$
$+q^{2}(q-1)B(h(l,1))-q\mu(\varpi)B(1_{4}).$
Set $B=B_{1}$. By putting $l=0$ in the above relations, applying the relation (6.39), and using the equation that $B_{2}(h(l,0))=\gamma B_{1}(h(l-1,0)s_{2}s_{1}s_{2})$ for $l{\geqslant}0$, we have
(6.41) $\alpha\gamma qB_{1}(s_{2})=\alpha\gamma(q-1)B_{1}(1_{4})+q^{2}B_{1}(h(-1,1)s_{1}s_{2}),$
(6.42) $\mu(\varpi)\gamma^{-1}(q^{-1}-1)B_{2}(1_{4})=\alpha\gamma(q-1)B_{1}(1_{4})+(q^{2}-1)B_{1}(h(-1,1)s_{1}s_{2}),$
(6.43) $\mu(\varpi)(\alpha q+1)B_{1}(s_{2})=q^{4}B_{1}(h(0,1)s_{1}s_{2})+\frac{\mu(\varpi)(\alpha q^{2}-\alpha q-q^{2}-1)}{q+1}B_{1}(1_{4}).$
Similarly, by putting $l=1$ in the third relation in (6.40), we have
(6.44) $\mu(\varpi)q^{-3}(1-q)B_{2}(1_{4})=\mu(\varpi)\alpha q^{-2}(q-1)B_{1}(1_{4})+(q^{2}-1)B_{1}(h(0,1)s_{1}s_{2}).$
From (6.41), (6.42), (6.43), and (6.44), we obtain $B_{2}(1_{4})=\alpha^{-1}B_{1}(1_{4})$. The non-vanishing of $B_{1}$ at $1_{4}$ is a consequence of [32, Theorem 9.3]. ∎
Proposition 6.10.
Suppose that $\pi$ is of type IIIa. Let $S\in{\rm{Sym^{2}}}(F)$ be such that $L/F$ is an unramified field extension. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ are unramified characters. When $\phi=1_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$ and ${\operatorname{Re}}(s)>1$, we have
$Z(\phi,B,s,\mu;\eta)=\frac{\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}L(s+1/2,\pi,\mu)\cdot B(1_{4}),\quad B\in{\mathscr{B}}(s,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}.$
Proof.
By substituting $B=B_{i}$ in (6.34) for $i=1,2$, the statement is immediate for $B=B_{i}$ by (6.39), the first relation in (6.40), and Lemma 6.9. Since $B_{1}$ and $B_{2}$ span ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$, we complete the proof. ∎
Corollary 6.11.
Let $\pi$, $S$, $\mu$, and $\Lambda$ be as in Proposition 6.10, then
$\varepsilon(\pi,s,\mu,\psi)=\mu(\varpi)^{2}q^{2(\frac{1}{2}-s)}.$
6.6. Computation for newforms of type ${\rm VI}b$ (case6)
Suppose that $\pi$ is of type VIb. By [32, Table 3], any element $B\in{\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ satisfies
(6.45) $T_{1,0}B=\gamma qB,\quad\eta B=\gamma B.$
By [32, Theorem 9.3], the value $B(1_{4})\not=0$ if and only if $B=0$.
Proposition 6.12.
Suppose that $\pi$ is of type VIb. Let $S\in{\rm{Sym^{2}}}(F)$ be such that $L/F$ is an unramified field extension. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ are unramified characters. When $\phi=1_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$ and ${\operatorname{Re}}(s)>1$, we have
$Z(\phi,B,s,\mu;\eta)=\frac{\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}L(s+1/2,\pi,\mu)\cdot B(1_{4}),\quad B\in{\mathscr{B}}(s,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}.$
Proof.
In our case, the relations (6.40) are also true. The proof is similar to that of Proposition 6.10. ∎
Corollary 6.13.
Let $\pi$, $S$, $\mu$, and $\Lambda$ be as in Proposition 6.12, then
$\varepsilon(\pi,s,\mu,\psi)=\mu(\varpi)^{2}q^{2(\frac{1}{2}-s)}.$
6.7. The values of Bessel models at $1_{4}$ for old forms
For quasi-characters $\chi$, $\chi^{\prime}$, and $\sigma$ on $F^{\times}$ which are trivial on $\mathfrak{o}^{\times}$, the representation $(\chi\times\chi^{\prime}\rtimes\sigma,V)$ can be realized as the right regular representation of the space of all smooth functions $f$ on $\mathsf{G}(F)$ satisfying
$f\left(\left[\begin{smallmatrix}a&\ast&\ast&\ast\\ 0&b&\ast&\ast\\ 0&0&ca^{-1}&0\\ 0&0&\ast&cb^{-1}\end{smallmatrix}\right]g\right)=|a^{2}b||c|^{-\frac{3}{2}}\chi(a)\chi^{\prime}(b)\sigma(c)f(g),\quad a,b,c\in F^{\times},g\in\mathsf{G}(F).$
The space $V$ has a unique $K$-invariant element $f_{K}\in V$ such that $f_{K}(1_{4})=1$. Let $I=K\cap\left[\begin{smallmatrix}\mathfrak{o}&{\mathfrak{p}}&\mathfrak{o}&\mathfrak{o}\\ \mathfrak{o}&\mathfrak{o}&\mathfrak{o}&\mathfrak{o}\\ {\mathfrak{p}}&{\mathfrak{p}}&\mathfrak{o}&\mathfrak{o}\\ {\mathfrak{p}}&{\mathfrak{p}}&{\mathfrak{p}}&\mathfrak{o}\end{smallmatrix}\right]$ be the Iwahori subgroup of $K$. For $w\in W$, let $f_{w}\in V^{I}$ be the unique element such that $f_{w}|_{K}=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{IwI}$.
By [33, Table 3], the space $V^{I}$ is 8-dimensional space spanned by
$f_{1_{4}},\quad f_{s_{1}},\quad f_{s_{2}},\quad f_{s_{2}s_{1}},\quad f_{s_{1}s_{2}s_{1}},\quad f_{s_{1}s_{2}},\quad f_{s_{1}s_{2}s_{1}s_{2}},\quad f_{s_{2}s_{1}s_{2}}.$
A basis of $4$-dimensional space $V^{K_{0}({\mathfrak{p}})}$ is given by
$f_{1}^{{\rm I}}:=f_{1_{4}}+f_{s_{1}},\quad f_{2}^{{\rm I}}:=f_{s_{2}}+f_{s_{2}s_{1}},\quad f_{3}^{{\rm I}}:=f_{s_{1}s_{2}s_{1}}+f_{s_{1}s_{2}},\quad f_{4}^{{\rm I}}:=f_{s_{1}s_{2}s_{1}s_{2}}+f_{s_{2}s_{1}s_{2}}.$
Recall that $\pi$ is isomorphic to $\chi\times\chi^{\prime}\rtimes\sigma$ or $\chi 1_{{\bf GL}_{2}}\rtimes\sigma$ for some characters $\chi$, $\chi^{\prime}$, and $\sigma$ on $F^{\times}$ according as $\pi$ is of type I or IIb. The representation $(\chi 1_{{\bf GL}_{2}}\rtimes\sigma,V)$ can be realized as a subrepresentation of $|\cdot|^{1/2}\chi\times|\cdot|^{-1/2}\chi\rtimes\sigma$, and a basis of $3$-dimensional space $V^{K_{0}({\mathfrak{p}})}$ is given by
$f_{1}^{{\rm II}b}:=f_{1_{4}}+f_{s_{1}},\quad f_{2}^{{\rm IIb}}:=f_{s_{2}}+f_{s_{2}s_{1}}+f_{s_{1}s_{2}s_{1}}+f_{s_{1}s_{2}},\quad f_{3}^{{\rm IIb}}:=f_{s_{1}s_{2}s_{1}s_{2}}+f_{s_{2}s_{1}s_{2}}.$
The representation matrices of operators $T_{1,0}$ and $\eta$ on $V_{\pi}^{K_{0}({\mathfrak{p}})}$ are given in [33, Table 3] as follows;
(6.49) $\begin{array}[]{|c|c|c|}\hline\cr{}\hfil&{{\rm type\ I}}&{{\rm type\ IIb}}\\ \hline\cr{T_{1,0}}&{\left[\begin{smallmatrix}\alpha\beta\gamma q^{3/2}&0&0&0\\ \alpha\beta\gamma(q-1)q^{1/2}&\beta\gamma q^{3/2}&0&0\\ \alpha\beta\gamma(q-1)q^{1/2}&\beta\gamma(q-1)q^{1/2}&\alpha\gamma q^{3/2}&0\\ \alpha\beta\gamma(q-1)q^{1/2}&\beta\gamma(q-1)q^{1/2}&\alpha\gamma(q-1)q^{1/2}&\gamma q^{3/2}\end{smallmatrix}\right]}&{\left[\begin{smallmatrix}\alpha^{2}\gamma q^{3/2}&0&0\\ \alpha^{2}\gamma(q-1)q^{1/2}&\alpha\gamma q^{2}&0\\ \alpha^{2}\gamma(q-1)q^{1/2}&\alpha\gamma(q^{2}-1)&\gamma q^{3/2}\end{smallmatrix}\right]}\\ \hline\cr{\eta}&{\left[\begin{smallmatrix}0&0&0&\gamma q^{3/2}\\ 0&0&\alpha\gamma q^{1/2}&0\\ 0&\beta\gamma q^{-1/2}&0&0\\ \alpha\beta\gamma q^{-3/2}&0&0&0\end{smallmatrix}\right]}&{\left[\begin{smallmatrix}0&0&\gamma q^{3/2}\\ 0&\alpha\gamma&0\\ \alpha^{2}\gamma q^{-3/2}&0&0\end{smallmatrix}\right]}\\ \hline\cr\end{array}$
Now we assume that $\pi$ has the trivial character. Let $(\ell_{\pi}^{0},f_{K})$ be the Bessel data for $\pi$ defined above. For $\bullet\in\{{\rm I},{\rm IIb}\}$, we set a basis $\{B_{i}^{\bullet}\}_{1{\leqslant}i{\leqslant}\dim V_{\pi}^{K_{0}({\mathfrak{p}})}}$ of $V_{\pi}^{K_{0}({\mathfrak{p}})}$ by
$B_{i}^{\bullet}(g):=\ell_{\pi}^{0}(\pi(g)f_{i}^{\bullet}),\quad g\in\mathsf{G}(F).$
Then the following result holds.
Proposition 6.14.
Let $\pi$ be a smooth admissible irreducible $K$-spherical representation of type I or IIb, and $\{B_{i}^{\bullet}\}$ be as above for $\bullet\in\{{\rm I},{\rm IIb}\}$. Then
  • (i)
    If $\pi$ is of type I,
    $B_{1}^{{\rm I}}(1_{4})=\frac{\alpha\beta}{(q-\alpha)(q-\beta)},\quad B_{2}^{{\rm I}}(1_{4})=\frac{-q\beta}{(q-\alpha)(q-\beta)},$
    $B_{3}^{{\rm I}}(1_{4})=\frac{-q\alpha}{(q-\alpha)(q-\beta)},\quad B_{4}^{{\rm I}}(1_{4})=\frac{q^{2}}{(q-\alpha)(q-\beta)}.$
  • (ii)
    If $\pi$ is of type IIb,
    $B_{1}^{{\rm IIb}}(1_{4})=\frac{\alpha^{2}}{(q^{1/2}-\alpha)(q^{3/2}-\alpha)},\quad B_{2}^{{\rm IIb}}(1_{4})=\frac{-q^{1/2}(1+q)\alpha}{(q^{1/2}-\alpha)(q^{3/2}-\alpha)},$
    $B_{3}^{{\rm IIb}}(1_{4})=\frac{q^{2}}{(q^{1/2}-\alpha)(q^{3/2}-\alpha)}.$
Proof.
Set $d:=\dim V_{\pi}^{K_{0}({\mathfrak{p}})}$. By the construction of the elements $\{f_{i}^{\bullet}\}_{1{\leqslant}i{\leqslant}d}$, we have $B_{\pi}^{0}=\sum_{i=1}^{d}B_{i}^{\bullet}$. For any $\ell\in{\mathbb{Z}}_{{\geqslant}0}$, we act by $T_{1,0}^{l}$ on the both sides and evaluate for $g=1_{4}$, then the first relation in (6.40) implies that
$B_{\pi}^{0}(h(\ell,0))=\sum_{i=1}^{d}(T_{1,0}^{\ell}B_{i}^{\bullet})(1_{4}).$
The values of the left hand side is given in [6, Corollary 1.8]. Hence by using the table (6.49) and a direct calculation, we can solve the values of $B_{i}^{\bullet}(1_{4})$. ∎
6.8. Local periods
[…]
$\mathbb{I}_{\pi,K_{0}({\mathfrak{p}})}^{(s)}(\xi^{\pi},\phi,\Lambda,\mu;g)=\sum_{B\in\mathcal{B}((V_{\pi}^{*})^{S,\Lambda},K_{0}({\mathfrak{p}}))}Z^{*}(\phi,B,s,\mu;g)\overline{B(1_{4})}.$
where
$Z^{*}(\phi,B,s,\mu;g):=L(s+1/2,\pi,\mu)^{-1}Z(\phi,B,s,\mu;g)$
is the normalized zeta integral and $\mathcal{B}((V_{\pi}^{*})^{S,\Lambda},K_{0}({\mathfrak{p}}))$ is an orthonormal basis of ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ with respect to the inner product on ${\mathscr{B}}(S,\Lambda)[\pi]$ via the $\mathsf{G}(F)$-isomorphism
(6.50) $V_{\pi}\owns\xi\mapsto(g\mapsto\ell_{\pi}(\pi(g)\xi))\in{\mathscr{B}}(S,\Lambda)[\pi].$
6.8.1. Local periods for type I and IIb
[…]
For $\bullet\in\{{\rm I},{\rm IIb}\}$, the set $\{\langle B_{i}^{\bullet}|B_{i}^{\bullet}\rangle^{-\frac{1}{2}}B_{i}^{\bullet}\}_{1{\leqslant}i{\leqslant}\dim{V_{\pi}^{K_{0}({\mathfrak{p}})}}}$ forms an orthonormal basis of ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$. By [37], the values $\langle B_{i}^{\bullet}|B_{i}^{\bullet}\rangle$ are given as follows:
$\langle B_{i}^{\rm{I}}|B_{i}^{\rm{I}}\rangle=q^{i-1}(q+1)\quad(1{\leqslant}i{\leqslant}4),$
$\langle B_{1}^{\rm{IIb}}|B_{1}^{\rm{IIb}}\rangle=(q+1),\quad\langle B_{2}^{\rm{IIb}}|B_{2}^{\rm{IIb}}\rangle=q(q+1)^{2},\quad\langle B_{3}^{\rm{IIb}}|B_{3}^{\rm{IIb}}\rangle=q^{3}(q+1).$
By Propositions 6.8 and 6.14, and the matrix representation of $T_{1,0}$ and $\eta$ given in (6.49), we can calculate the local period by a direct calculation.  [^2] We have used Mathematica 13.
Proposition 6.15.
Let $\pi$ be a smooth admissible irreducible unitary representation of type I or type IIb with trivial central character. Suppose that $\mu:F^{\times}\to{\mathbb{C}}^{1}$ is an unramified character. When $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$, we have
$\mathbb{I}_{\pi,K_{0}({\mathfrak{p}})}^{(s)}(\xi_{0}^{\pi},\phi,\Lambda,\mu;\eta)$ $=\frac{2(q-1)}{q^{5}(q^{2}+1)}L(1,\pi;{\rm Std})$
×$\times\left\{\mu(\varpi)^{-1}q^{s+1}+\mu(\varpi)q^{-s+1}-\frac{1}{q+1}{\operatorname{tr}}(T_{1,0}+q\eta|_{V_{\pi}^{K_{0}({\mathfrak{p}})}})\right\},$
where $L(s,\pi;{\rm Std})$ is the standard L-function of $\pi$ defined by
$L(s,\pi;{\rm Std})=(1-\alpha q^{-s})^{-1}(1-\beta q^{-s})^{-1}(1-\alpha^{-1}q^{-s})^{-1}(1-\beta^{-1}q^{-s})^{-1}(1-q^{-s})^{-1}.$
or
$L(s,\pi;{\rm Std})=(1-\alpha q^{-s+\frac{1}{2}})^{-1}(1-\alpha q^{-s-\frac{1}{2}})^{-1}(1-\alpha^{-1}q^{-s+\frac{1}{2}})^{-1}(1-\alpha^{-1}q^{-s-\frac{1}{2}})^{-1}(1-q^{-s})^{-1}.$
according as $\pi$ is of type I or IIb
6.8.2. Local period for type IIIa
Suppose that $\pi$ is of type IIIa. Then, as remarked before, $|\alpha|=|\gamma|=1$ and $\alpha\not=1$. By Lemma 6.9, there exists a basis $\{B_{i}\}_{i=1,2}$ of ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{(}{\mathfrak{p}})}$, unique up to constant, satisfying the relations (6.39) and $B_{2}(1_{4})=\alpha^{-1}B_{1}(1_{4})\not=0$. Thus, we can uniquely fix $\{B_{i}^{{\rm IIIa}}\}_{i=1,2}$ by requiring $B_{1}^{{\rm IIIa}}(1_{4})=1$ and the datum $(\ell_{\pi},\xi^{\pi})$ by $\ell_{\pi}(\xi^{\pi})=B^{{\rm IIIa}}_{1}$ to form (6.50) and induce the $\mathsf{G}(F)$-invariant inner product on ${\mathscr{B}}(S,\Lambda)[\pi]$ such that $\langle B_{1}^{{\rm IIIa}}|B_{1}^{{\rm IIIa}}\rangle=1$. Since $\langle B_{1}|B_{2}\rangle=0$ by [7, Lemma 2.11], we then have $\langle B_{-}^{{\rm IIIa}}|B_{-}^{{\rm IIIa}}\rangle=1$ and $\langle B_{+}^{{\rm IIIa}}|B_{-}^{{\rm IIIa}}\rangle=0$ due to $|\alpha|=1$.
Proposition 6.16.
Let $\pi$ be a smooth admissible irreducible unitary representation of type IIIa. Suppose that $:F^{\times}\to{\mathbb{C}}^{1}$ is an unramified character. When $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$, we have
$\mathbb{I}_{\pi,K_{0}({\mathfrak{p}})}^{(s)}(\xi^{\pi},\phi,\Lambda,\mu;\eta)=\frac{2\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}$
Proof.
By Proposition 6.10 and noting that $|\alpha|=1$, we have
$\mathbb{I}_{\pi,K_{0}({\mathfrak{p}})}^{(s)}(\xi^{\pi},\phi,\Lambda,\mu;\eta)$ $=\frac{\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}\left\{|B_{1}^{{\rm IIIa}}(1_{4})|^{2}+|B_{2}^{{\rm IIIa}}(1_{4})|^{2}\right\}$
$=\frac{\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}\left\{1+|\alpha^{-1}|^{2}\right\}$
$=\frac{2\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}.\qed$
6.8.3. Local period for type VIb
Suppose that $\pi$ is of type VIb. We recall that $\dim(V_{\pi}^{K_{0}({\mathfrak{p}})})=\dim({\mathscr{B}}(S,\Lambda)^{K_{0}({\mathfrak{p}})})=1$; thus, by [32, Theorem 9.3], there exists a unique element $B^{{\rm VIb}}\in{\mathscr{B}}(S,\Lambda)^{K_{0}({\mathfrak{p}})}$ satisfying $B^{{\rm VIb}}(1_{4})=1$. We fix the datum $(\ell_{\pi},\xi^{\pi})$ by $\ell_{\pi}(\xi^{\pi})=B^{{\rm VIb}}$. Then the following result is immediate from Proposition 6.12.
Proposition 6.17.
Let $\pi$ be a smooth admissible irreducible unitary representation of type VIb. Suppose that $:F^{\times}\to{\mathbb{C}}^{1}$ is an unramified character. When $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}$, we have
$\mathbb{I}_{\pi,K_{0}({\mathfrak{p}})}^{(s)}(\xi^{\pi},\phi,\Lambda,\mu;\eta)=\frac{\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}q^{s+1}}{q^{2}+1}.$